### v1

# Factor Augmented Tensor-on-Tensor Neural Networks

Abstract
This paper studies the prediction task of tensor-on-tensor regression in which both covariates and responses are multi-dimensional arrays (a.k.a., tensors) across time with arbitrary tensor order and data dimension. Existing methods either focused on linear models without accounting for possibly nonlinear relationships between covariates and responses, or directly employed black-box deep learning algorithms that failed to utilize the inherent tensor structure. In this work, we propose a Factor Augmented Tensor-on-Tensor Neural Network (FATTNN) that integrates tensor factor models into deep neural networks. We begin with summarizing and extracting useful predictive information (represented by the “factor tensor”) from the complex structured tensor covariates, and then proceed with the prediction task using the estimated factor tensor as input of a temporal convolutional neural network. The proposed methods effectively handle nonlinearity between complex data structures, and improve over traditional statistical models and conventional deep learning approaches in both prediction accuracy and computational cost. By leveraging tensor factor models, our proposed methods exploit the underlying latent factor structure to enhance the prediction, and in the meantime, drastically reduce the data dimensionality that speeds up the computation. The empirical performances of our proposed methods are demonstrated via simulation studies and real-world applications to three public datasets. Numerical results show that our proposed algorithms achieve substantial increases in prediction accuracy and significant reductions in computational time compared to benchmark methods.
Keywords: Tensor regression; Tensor-on-tensor regression; Tensor factor model;
Time series forecasting; Temporal convolutional network.
1 Introduction
Motivated by a wide range of industrial and scientific applications, forecasting tasks using one tensor series to predict another has become increasingly prevalent in various fields. In finance, tensor models are used to forecast future stock prices by analyzing multi-dimensional stock attributes over time (Tran et al.,, 2017). In meteorology, tensor models predict future weather conditions using historical data formatted as multi-dimensional tensors (latitude, longitude, time, etc.) (Bilgin et al.,, 2021). In neuroscience, tensor models are employed to process medical imaging like MRI/fMRI scans (Wei et al.,, 2023).
Though predictive modeling of tensors has emerged rapidly over the past decade, most existing studies primarily focus on scenarios when either covariates or responses are tensors, such as scalar-on-tensor regression (Guo et al.,, 2011; Zhou et al.,, 2013; Wimalawarne et al.,, 2016; Li et al.,, 2018; Ahmed et al.,, 2020), matrix-on-tensor regression (Hoff,, 2015; Kossaifi et al.,, 2020), and tensor response regression (Li and Zhang,, 2017; Sun and Li,, 2017). Tensor-on-tensor [^1] Following the literature, we distinguish “tensor-on-tensor regression” from “tensor regression”. “Tensor regression” refers to regression models with tensor covariates and scalar/vector/matrix responses. “Tensor-on-tensor regression” refers to regression models with both tensor covariates and tensor responses. As a matter of fact, tensor-on-tensor regression encompasses tensor regression. prediction, when both covariates and responses are tensors, remains a challenging task due to the inherent complexities of tensor structures.
One commonly used strategy for tensor-on-tensor modeling is to flatten the tensors into matrics (or even vectors) (Kilmer and Martin,, 2011; Choi and Vishwanathan,, 2014) so that matrix-based analytical tools become applicable. However, flattening can lead to a loss of spatial or temporal information, especially for datasets that are inherently dependent on data structures. For instance, in image processing, the relative positions of pixels carry important information about object shapes, which is lost when the image is flattened.
Alternatively, there is a handful of recent works on tensor-on-tensor regression models that do not adopt flattening but deal with the tensor structures directly; see Lock, (2018); Liu et al., (2020); Gahrooei et al., (2021); Luo and Zhang, (2022). However, these methods posit a linear prediction model between covariates and responses, limiting their capacity to capture potential nonlinear relationships.
Another emerging trend in tensor-on-tensor prediction involves neural networks, such as recurrent neural networks, convolutional neural networks, convolutional recurrent neural networks, and temporal convolutional networks, to name a few. Deep neural network is a powerful tool in prediction tasks owing to its ability to learn intricate patterns from complex datasets. Structured with multiple layers of interconnected neurons, neural networks excel in capturing nonlinearity in data. Nevertheless, they often operate in a black-box manner and typically require extensive computational resources.
In this work, we develop a Factor Augmented Tensor-on-Tensor Neural Network (FATTNN) for forecasting a sequence of tensor responses using tensor covariates of arbitrary tensor order and data dimensions. One key innovation of our method is the integration of tensor factor models into deep neural networks for tensor-on-tensor regression. Factor models have been a widely utilized tool in matrix/vector data analysis (Bai and Ng,, 2002; Stock and Watson,, 2002; Pan and Yao,, 2008; Lam and Yao,, 2012) for understanding common dependence among multi-dimensional covariates. In recent years, researchers have advanced the methodology of factor models to the context of tensor time series (Chen et al., 2022a, ; Chen et al., 2022b, ; Han et al.,, 2020, 2024, 2022; Han and Zhang,, 2022). In our models, we borrow the strengths from tensor factor models, offering a more general approach than traditional vector factor analysis for prediction (Sen et al.,, 2019; Fan et al.,, 2017; Yu et al.,, 2022).
Our contributions can be summarized in three folds.
  • We propose a FATTNN model for tensor-on-tensor prediction that integrates tensor factor models into deep neural networks. Via a tensor factor model, FATTNN exploits the latent factor structures among the complex structured tensor covariates, and extracts useful predictive information for subsequent modeling. The utilization of the factor model reduces the data dimension drastically while preserving the tensorial data structures, contributing to substantial increases in prediction accuracy and significant reductions in computational time.
  • In addition to the improved prediction accuracy and reduced computation cost, FATTNN offers several advantages over existing methods. First, it preserves the inherent tensor structure without any flattening or tensor contraction operations, preventing the possible loss of spatial or temporal information. Second, the utilization of neural networks equips FATTNN with exceptional capability in modeling nonlinearity in response-covariate relationships.
  • FATTNN provides a comprehensive framework for the general supervised learning question of predicting tensor responses using tensor covariates. Although initially introduced under the problem settings of tensor-on-tensor time series forecasting, the proposed FATTNN framework is not limited to time series data, but can accommodate a variety of data types. With different types of factor models and choices of neural network architectures tailored to specific characteristics of the data, FATTNN can be naturally adapted to tensor-on-tensor regression for various data types.
This paper is organized as follows. Section 2 introduces some useful tensor terminology and sets up the problem of interest. Section 3 presents the methodological details of the proposed FATTNN algorithms. Section 4 includes applications to three real-world datasets to demonstrate the finite-sample performance of our proposed methods. Section 5 concludes the paper with discussions. Theoretical proofs, simulation studies, and supplemental numerical results are provided in appendices.
2 Preliminaries
Notation. Before proceeding, we first set up some notation. Let $\mathcal{X}\in\mathbb{R}^{d_{1}\times\dots\times d_{K}}$ denote a $K$-th order tensor of dimension $d_{1}\times d_{2}\times\dots\times d_{K}$. Let $\mathcal{X}_{i_{1},\dots,i_{K}}$ denote the $({i_{1},\dots,i_{K}})$-th entry of the tensor $\mathcal{X}$ and $\mathcal{X}[\mathcal{I}_{1},\dots,\mathcal{I}_{K}]$ denotes the sub-tensor of $\mathcal{X}$ for $\mathcal{I}_{1}\subseteq\{1,\dots,d_{1}\},\dots,\mathcal{I}_{K}\subseteq\{1,\dots,d_{K}\}$. The matricization operation $\hbox{\rm mat}_{k}(\cdot)$ unfolds an order-$K$ tensor along mode $k$ to a matrix, say ${\cal X}\in\mathbb{R}^{d_{1}\times\dots\times d_{K}}$ to $\hbox{\rm mat}_{k}({\cal X})\in\mathbb{R}^{d_{k}\times d_{-k}}$ where $d_{k}=\prod_{j\neq k}d_{j}$ and its detailed definition is provided in the appendix. The Frobenius norm of tensor ${\cal X}$ is defined as $\|{\cal X}\|_{\rm F}=(\sum_{i_{1},i_{2},\cdots,i_{K}}{\cal X}_{i_{1},\dots,i_{K}}^{2})^{1/2}$. The Tucker rank of an order-$K$ tensor ${\cal X}$, denoted by Tucrank$({\cal X})$, is defined as a $K$-tuple $(r_{1},...,r_{K})$, where $r_{k}=\text{rank}(\hbox{\rm mat}_{k}({\cal X}))$. Any Tucker rank-$(r_{1},...,r_{K})$ tensor ${\cal X}$ admits the following Tucker decomposition (Tucker,, 1966): ${\cal X}={\cal S}\times_{1}U_{1}\times_{2}\cdots\times_{K}U_{K}$, where ${\cal S}\in\mathbb{R}^{r_{1}\times\cdots\times r_{K}}$ is the core tensor and $U_{k}=\text{SVD}_{r_{k}}(\hbox{\rm mat}_{k}({\cal X}))$ is the mode-$k$ top $r_{k}$ left singular vectors. Here, the mode-$k$ product of ${\cal X}\in\mathbb{R}^{d_{1}\times\cdots\times d_{K}}$ with a matrix $B\in\mathbb{R}^{d_{k}\times r_{k}}$, denoted by ${\cal X}\times_{k}B$, is a $d_{1}\times\cdots\times d_{k-1}\times r_{k}\times d_{k+1}\times\cdots\times d_{K}$-dimensional tensor, and its detailed definition is provided in Section 3 and Appendix A. The following abbreviations are used to denote the tensor-matrix product along multiple modes: ${\cal X}\times_{k=1}^{K}U_{k}={\cal X}\times_{1}U_{1}\times_{2}\cdots\times_{K}U_{K}$ and ${\cal X}\times_{\ell\neq k}U_{\ell}={\cal X}\times_{1}U_{1}\times_{2}\cdots\times_{k-1}U_{k-1}\times_{k+1}U_{k+1}\times_{k+2}\cdots\times_{K}U_{K}$. For any two matrices $A\in\mathbb{R}^{m_{1}\times r_{1}},B\in\mathbb{R}^{m_{2}\times r_{2}}$, denote the Kronecker product $\odot$ as $A\odot B\in\mathbb{R}^{m_{1}m_{2}\times r_{1}r_{2}}$. For any two tensors ${\cal A}\in\mathbb{R}^{m_{1}\times m_{2}\times\cdots\times m_{K}},{\cal B}\in\mathbb{R}^{r_{1}\times r_{2}\times\cdots\times r_{N}}$, denote the tensor product $\otimes$ as ${\cal A}\otimes{\cal B}\in\mathbb{R}^{m_{1}\times\cdots\times m_{K}\times r_{1}\times\cdots\times r_{N}}$, such that $({\cal A}\otimes{\cal B})_{i_{1},...,i_{K},j_{1},...,j_{N}}=({\cal A})_{i_{1},...,i_{K}}({\cal B})_{j_{1},...,j_{N}}.$
Problem Setup. Suppose we observe a time series dataset comprising $n$ temporal samples ${\cal D}=\{({\cal X}_{t},{\cal Y}_{t}),1\leq t\leq n\}$, where $\mathcal{X}_{t}\in\mathbb{R}^{d_{1}\times\dots\times d_{K}}$ is a tensor of covariates and $\mathcal{Y}_{t}\in\mathbb{R}^{p_{1}\times\dots\times p_{q}}$ is the tensor response variable. For some newly observed covariates $\{{\cal X}_{t}\}_{t=n+1}^{n+m}$, the forecasting task of tensor-on-tensor time series regression is to accurately predict the corresponding future tensor responses $\{{\cal Y}_{t}\}_{t=n+1}^{n+m}$, given the observed time series $\{({\cal X}_{t},{\cal Y}_{t})\}_{t=1}^{n}$ and the new covariates $\{{\cal X}_{t}\}_{t=n+1}^{n+m}$. Here, $m$ is the forecast horizon. In subsequent discussions, we denote the series in training time range as $\mathcal{X}^{(tr)}=\{\mathcal{X}_{t}\}_{t=1}^{n}$ and $\mathcal{Y}^{(tr)}=\{\mathcal{Y}_{t}\}_{t=1}^{n}$, and the covariates in forecasting time range as $\mathcal{X}^{(te)}=\{\mathcal{X}_{t}\}_{t=n+1}^{n+m}$.
3 Methodology: Factor Augmented Tensor-on-Tensor time series regression
In this section, we introduce our Factor Augmented Tensor-on-Tensor Neural Network, dubbed as FATTNN. Our work explores a broad category of challenges known as tensor-on-tensor regression, which seeks to elucidate the relationships between covariates and responses that may manifest as scalars, vectors, matrices, or higher-order tensors. As illustrated in Figure 1, our method consists of two components. First, factor tensor features are extracted by low-rank tensor factorization. Second, we develop a Tensor-on-Tensor Neural Network based on Temporal Convolutional Network (TCN) (Bai et al.,, 2018). In Section 3.1, we present a tensor factor model. This model can capture global patterns among the observed time series covariate tensors, by representing each of the original covariate tensor ${\cal X}_{t}$ as a multilinear combination of $r_{1}\times\cdots\times r_{K}$ basis across each tensor mode $k$ plus noise, where $r_{k}\ll d_{k}$. It offers a more general approach than the traditional vector factor analysis for prediction (Sen et al.,, 2019; Fan et al.,, 2017; Luo et al.,, 2022; Yu et al.,, 2022; Fan and Gu,, 2023), and preserves the tensor structures of features. It also provides dimension reduction, which significantly reduces computational complexity. In Section 3.2, we describe how the output from the global tensor factor model can be used as an input covariate for a TCN to predict future response tensors.
Figure 1: A graphical illustration of the proposed FATTNN.
The input is the observed time series $\{({\cal X}_{t},{\cal Y}_{t})\}_{t=1}^{n}$ and the new covariates $\{{\cal X}_{t}\}_{t=n+1}^{n+m}$.
The output is the forecasted future tensor responses, denoted by $\{\widehat{\cal Y}_{t}\}_{t=n+1}^{n+m}$
3.1 Tensor Factor Model (TFM)
Our representation learning module utilizes low rank tensor factorization for covariate tensor. The idea is to factorize the training covariate tensors ${\cal X}_{t}\in\mathbb{R}^{d_{1}\times\cdots\times d_{K}},1\leq t\leq n,$ into low dimensional tensors ${\cal F}_{t}\in\mathbb{R}^{r_{1}\times\cdots\times r_{K}}$ and linear combination matrices (loading matrices) $A_{k}\in\mathbb{R}^{d_{k}\times r_{k}}$, where $r_{k}\ll d_{k}$. Specifically, the model can be expressed as
$\mathcal{X}_{t}=\mathcal{F}_{t}\times_{1}A_{1}\times_{2}\ldots\times_{K}A_{K}+{\cal E}_{t},$ (1)
where ${\cal E}_{t}$ is idiosyncratic noise of the tensor ${\cal X}_{t}$, and $t$ is a time point. Here the $k$-mode product of $\mathcal{X}\in\mathbb{R}^{d_{1}\times d_{2}\times\ldots\times d_{K}}$ with a matrix $U\in\mathbb{R}^{d_{k}^{\prime}\times d_{k}}$, denoted as $\mathcal{X}\times_{k}U$, is an order $K$-tensor of size $d_{1}\times\ldots\times d_{k-1}\times d_{k}^{\prime}\times d_{k+1}\times\ldots\times d_{K}$ such that $(\mathcal{X}\times_{k}U)_{i_{1},\ldots,i_{k-1},j,i_{k+1},\ldots,i_{K}}=\sum_{i_{k}=1}^{d_{k}}\mathcal{X}_{i_{1},i_{2},\ldots,i_{K}}U_{j,i_{k}}.$ If $K=2$, the matrix factor model can be written as ${\cal X}_{t}=A_{1}{\cal F}_{t}A_{2}^{\top}+{\cal E}_{t}$. The latent core tensor ${\cal F}_{t}$, which typically encapsulates the most critical information, generally possesses dimensions significantly smaller than those of ${\cal X}_{t}$. Thus, the covariate tensor ${\cal X}_{t}$ can be thought of to be comprised of a basis tensor features ${\cal F}_{t}$ that capture the global patterns in the whole data-set, and all the original covariates are multilinear combinations of these basis tensor features. In the following subsection, we will discuss how a TCN can be utilized to leverage the temporal structure of the training data.
For predicting future response tensors, given a new covariate tensor ${\cal X}_{t}$, we can also extract low-dimensional feature tensors ${\cal F}_{t}$ using the estimated deterministic loading matrices $A_{k}$ and model (1). This approach efficiently captures the essential characteristics from the high-dimensional input data. The reduced dimensionality allows for more efficient data processing and analysis while retaining essential features.
To capture the underlying tensor features of the covariates, we employ a Time series Inner-Product Unfolding Procedure (TIPUP). It utilizes the linear temporal dependence among the covariates. The TIPUP method performs singular value decomposition (SVD) for each tensor mode $k=1,...,K$:
$\widehat{A}_{k}=\text{LSVD}_{r_{k}}\left(\frac{1}{n}\sum_{t=1}^{n}\hbox{\rm mat}_{k}({\cal X}_{t})\hbox{\rm mat}_{k}^{\top}({\cal X}_{t})\right).$ (2)
The estimation methods in (2) can be further enhanced through an iterative refinement algorithm, detailed in the appendix. The factor tensors in the training data thus can be estimated by
^$\widehat{\cal F}_{t}={\cal X}_{t}\times_{k=1}^{K}\widehat{A}_{k}^{\top},\quad\text{for }1\leq t\leq n.$
In the next proposition, we demonstrate the provable benefits of using a tensor factor model with Tucker low-rank structure (1). Specifically, we establish the convergence rates of the linear space of the loading matrices, which in turn ensure the accuracy of the estimated factor tensors.
Proposition 3.1.
Assume each elements of the idiosyncratic noise ${\cal E}_{t}$ in (1) are i.i.d. $N(0,1)$. The ranks $r_{1},...,r_{K}$ are fixed. The factor process ${\cal F}_{t}$ is weakly stationary and its cross-outer-product process is ergodic in the sense of $\frac{1}{n}\sum_{t=1}^{n}{\cal F}_{t}\otimes{\cal F}_{t}\rightarrow\mathbb{E}({\cal F}_{t}\otimes{\cal F}_{t})$ in probability, where the elements of $\mathbb{E}({\cal F}_{t}\otimes{\cal F}_{t})$ are all finite. In addition, the condition numbers of $A_{k}^{\top}A_{k}$ ($k=1,...,K$) are bounded. Let $\lambda=\prod_{k=1}^{K}\|A_{k}\|_{2}$ and grow as dimensions $d_{k}$ increases. Then, the iterative TIPUP estimator satisfies
$\|\widehat{A}_{k}(\widehat{A}_{k}^{\top}\widehat{A}_{k})^{-1}\widehat{A}_{k}^{\top}-A_{k}(A_{k}^{\top}A_{k})^{-1}A_{k}\|_{2}=O_{\mathbb{P}}\left(\frac{\max_{k}\sqrt{d_{k}}}{\lambda\sqrt{n}}+\frac{\max_{k}\sqrt{d_{k}}}{\lambda^{2}\sqrt{n}}\right),\qquad 1\leq k\leq K.$
3.2 Tensor-on-Tensor Neural Network based TCN (TTNN)
If we are equipped with a TCN that identifies the temporal patterns in the training dataset $({\cal X}^{(tr)},{\cal Y}^{(tr)})$, we can use this model to enhance the temporal structures in $({\cal F}^{(tr)},{\cal Y}^{(tr)})$. Thus, a TCN is applied to the input sequence of estimated factor tensors ^$\widehat{\cal F}_{1},\ldots,\widehat{\cal F}_{t}$ along with the corresponding output sequence of response tensors ${\cal Y}_{1},\ldots,{\cal Y}_{t}$ to encapsulate the temporal dynamics. Let $\eta(\cdot)$ represent this network. The whole temporal network $\eta(\cdot)$ can be trained using the low-rank factor tensors ^$\widehat{\cal F}_{1},\ldots,\widehat{\cal F}_{n}$ derived from the training dataset.
The trained temporal network $\eta(\cdot)$ can be effectively utilized for multi-step look-ahead prediction in a standard approach. The factor tensors for future time steps can be estimated using the following formula:
^$\widehat{\cal F}_{n+h}={\cal X}_{n+h}\times_{k=1}^{K}\widehat{A}_{k}^{\top},$
where $h>0$. Using the historical data points of factor tensors ^$\widehat{\cal F}_{t},1\leq t\leq n+h-1$, the prediction for the subsequent time step, ${\cal Y}_{n+h}$, is generated by
$\widehat{\cal Y}_{n+h}=\eta(\widehat{\cal F}_{1},...,\widehat{\cal F}_{n+h}).$
Our hybrid forecasting model incorporates a TCN that inputs past data points from the original raw response time series and the estimated factor tensors, enhancing prediction accuracy. The pseudo-code for our model is outlined in Algorithm 1.
Algorithm 1 Factor Augmented Tensor-on-Tensor Neural Network (FATTNN)
1:  Input: Observed tensor time series $(\mathcal{X}_{1},\mathcal{Y}_{1}),\ldots,(\mathcal{X}_{n},\mathcal{Y}_{n}),\mathcal{X}_{n+1},\ldots,\mathcal{X}_{n+m}$, rank $r_{k}$ for each mode $k$, forecasting horizon $m$.
2:  Output: Forecasts $\widehat{\mathcal{Y}}_{n+1},\ldots,\widehat{\mathcal{Y}}_{n+m}$.
3:  Compute the factor tensor ^$\widehat{\cal F}_{t}$ for $t=1,\ldots,n$ by TIPUP (2) or its iterative version in Algorithm S.1.
4:  Fit a TCN $\eta(\cdot)$ to the input sequence of factor tensors ^$\widehat{\cal F}_{1},\ldots,\widehat{\cal F}_{n}$ and the output sequence of response tensor ${\cal Y}_{1},...,{\cal Y}_{n}$ to capture the temporal dynamics.
5:  Forecast ${\cal Y}_{n+1},\ldots,{\cal Y}_{n+m}$ using the fitted TCN.
6:  for $h=1$ to $m$ do
7:     ^$\widehat{\cal F}_{n+h}={\cal X}_{n+h}\times_{k=1}^{K}\widehat{A}_{k}^{\top}$.
8:     $\widehat{\mathcal{Y}}_{n+h}=\eta(\widehat{\cal F}_{1},\ldots,\widehat{\mathcal{F}}_{n+h})$.
9:  end for
10:  return  $\widehat{\mathcal{Y}}_{n+1},\ldots,\widehat{\mathcal{Y}}_{n+m}$.
4 Numerical Results
In this section, we investigate the finite-sample performance of our proposed FATTNN methods via simulation studies and real-world applications. Specifically, we implement the proposed FATTNN [^2] Our code for implementing FATTNN is available in the supplement. with TIPUP algorithm to estimate the factor structures. For benchmark methods [^3] Multiway is implemented using the code provided by https://github.com/lockEF/MultiwayRegression, and TCN is implemented using the provided by https://github.com/locuslab/TCN . , we consider one traditional statistical model - the multiway tensor-on-tensor regression (Lock,, 2018) (denoted by Multiway), and one conventional deep learning approach – the temporal convolutional network (Bai et al.,, 2018) of $\mathcal{Y}$ regressed on $\mathcal{X}$ (denoted by TCN). Due to space limitations, simulation results are presented in Appendix 4.1.
Evaluation Metrics. We evaluate the performance of various methods with emphasis on two aspects: prediction accuracy and computational efficiency. To evaluate prediction accuracy, we compute the Mean Squared Error (MSE) over the testing data, i.e., $\text{MSE}=(n_{\text{test}}\cdot p_{1}\cdots p_{q})^{-1}\sum_{i\in\mathcal{D}_{\text{test}}}\|\mathcal{Y}_{i}^{(\text{obs})}-\mathcal{Y}_{i}^{(\text{pred})}\|_{F}^{2}$, where $\mathcal{D}_{\text{test}}$ denotes the testing set, $n_{\text{test}}$ is the number of samples in the testing set $\mathcal{D}_{\text{test}}$, and $(\mathcal{Y}_{i}^{(\text{obs})},\mathcal{Y}_{i}^{(\text{pred})})$ are the observed and predicted values of the $i$-th tensor response. In addition, we record the computational time of different methods to evaluate computational efficiency.
4.1 Simulation Studies
We carry out simulation studies under different scenarios with various configurations of sample size, data dimensions, tensor ranks, and relationships between covariates and responses. Recall that ${\cal X}_{t}\in\mathbb{R}^{d_{1}\times\cdots\times d_{K}}$, ${\cal F}_{t}\in\mathbb{R}^{r_{1}\times\cdots\times r_{K}}$, ${\cal Y}_{t}\in\mathbb{R}^{p_{1}\times\cdots\times p_{q}}$, for $1\leq t\leq n$.
Data Generating Process. We begin with generating the low-rank core tensor ${\cal F}_{t}$ from a tensor autoregressive model ${\rm vec}({\cal F}_{t})=\Phi\cdot{\rm vec}({\cal F}_{t-1})+{\rm vec}({\cal W}_{t})$ for $t=1,\ldots,n$. Here, ${\rm vec}({\cal F}_{t})$ denotes the vectorization [^4] We would like to clarify that our proposed FATTNN approach does not involve tensor vectorization in the model fitting. This vectorization is solely for the convenience of data generation in simulation. of a tensor ${\cal F}_{t}$, ${\cal W}_{t}\in\mathbb{R}^{r_{1}\times\cdots\times r_{K}}$ is an error tensor in which every element is randomly generated from a normal distribution. The coefficient matrix $\Phi\in\mathbb{R}^{(r_{1}r_{2}\cdots r_{K})\times(r_{1}r_{2}\cdots r_{K})}$ is constructed using the Kronecker product of a sequence of matrices $Q_{k}$, i.e., $\Phi=\odot_{k=1}^{K}Q_{k}$, in which each $Q_{k}\in\mathbb{R}^{r_{k}\times r_{k}}$ is a matrix with i.i.d. standard normal entries and orthonormalized through QR decomposition. The tensor time series ${\cal F}_{t}$ is randomly initiated in a distance past, and we discard the first 500 time points to stabilize the process.
Next, we generate the tensor covariates ${\cal X}_{t}$ from ${\cal F}_{t}$, following $\mathcal{X}_{t}=\lambda\mathcal{F}_{t}\times_{1}A_{1}\times_{2}\ldots\times_{K}A_{K}+{\cal E}_{t},$ for $t=1,\ldots n$, where $\lambda$ is a scalar that controls the signal-to-noise ratio in the tensor factor model. In our experiments, we set $\lambda=(\prod_{k=1}^{K}r_{k})^{1/2}$. The loading matrices $A_{k}\in\mathbb{R}^{d_{k}\times r_{k}}$ are generated with independent $N(0,1)$ entries, and then orthonormalized using QR decomposition. The error ${\cal E}_{t}\in\mathbb{R}^{d_{1}\times\cdots\times d_{K}}$ is also generated with each element independently drawn from normal distributions.
[…]
We consider the following three experimental configurations:
  1. (1)
    $(d_{1},d_{2},d_{3})=(25,25,12)$,$(r_{1},r_{2},r_{3})=(6,8,6)$, $(p_{1},p_{2},p_{3})=(3,3,2)$
    $n=500$, $s(z)=cos(z)$, $\sigma_{u}^{2}=1$.
  2. (2)
    $(d_{1},d_{2},d_{3})=(30,6,12)$, $(r_{1},r_{2},r_{3})=(6,3,2)$, $(p_{1},p_{2},p_{3})=(8,6,4)$,
    $n=400$, $s(z)=log(|z|)$, $\sigma_{u}^{2}=1$.
  3. (3)
    $(d_{1},d_{2},d_{3})=(12,3,12)$, $(r_{1},r_{2},r_{3})=(4,3,4)$, $(p_{1},p_{2},p_{3})=(3,3,3)$,
    $n=100$, $s(z)=log(e^{(z)}+1)$, $\sigma_{u}^{2}=0.5$.
Simulation Evaluations. In each setting, we use 70% for training the models and 30% for model evaluation. Comparisons of MSE, and computational time are summarized in Table 1 and Table 2. All reported metrics are averaged over 20 replications. Experiments were run on the research computing cluster provided by the University Center for Research Computing. From Table 1, we can see that FATTNN improves over the two benchmark methods by a large margin. The improvements can be attributed to two facets: (a) the utilization of tensor factors, summarizing useful predictive information effectively and reducing the number of noise covariates, and (b) the utilization of neural networks, possessing extraordinary ability in capturing complex relationships in the data and therefore exhibiting exceptional predictive power. On a separate note, Table 2 reveals the evident computational advantage of FATTNN compared to other neural network models such as TCN, owing to the effective dimension reduction through tensor factor models.
Table 1: Comparisons of prediction accuracy of different methods in simulation studies
(measured by MSE over the testing data)
Prediction Task TCN FATTNN MultiWay
Simulation 1 3.642 (1.91) 2.451 (1.57) 3.987 (2.00)
Simulation 2 670.5 (25.8) 511.9 (22.5) 805.1 (28.4)
Simulation 3 1074 (32.6) 696.7 (26.4) 1670 (40.3)
Note: The results are averaged over 20 replications.
Numbers in the brackets are standard errors.
Table 2: Comparisons of computational time (in the format of hh:mm:ss)
of different methods in simulation studies
Prediction Task TCN FATTNN Muiltway
Simulation 1 01:22:02 00:16:18 00:02:15
Simulation 2 01:26:30 00:17:54 00:02:30
Simulation 3 00:11:16 00:03:41 00:01:03
4.2 Real Data Examples
We evaluate the performance of different methods using five prediction tasks on three real-world datasets. The information about sample sizes in the datasets, data dimensions of covariates $(d_{1},\ldots,d_{K})$ and responses $(p_{1},\ldots,p_{q})$, and the ranks of core tensors $(r_{1},\ldots,r_{K})$ utilized in the analyses are summarized in Table 3. For the prediction tasks using the New York Taxi data and FMRI image datasets, we split the data into training sets (70%) and testing (30%). For the prediction tasks using the FAO dataset, we use 80% data for training and 20% for testing due to its relatively small sample size. We compare the prediction accuracy and computation time of three methods: FATTNN, TCN, and Multiway. The numerical prediction results along with the corresponding computation time are summarized in Table 4.
Table 3: Data information of the real-world applications
Prediction Task Sample Size Dimension of $\mathcal{X}_{t}$ Dimension of $\mathcal{Y}_{t}$ Rank of ${\cal F}_{t}$
FAO crop $\sim$ crop 62 (33,2,11) (13,2,11) (6,2,4)
FAO crop $\sim$ livestock 62 (26,4,5) (26,3,11) (6,4,2)
NYC taxi – midtown 504 (2,12,12,8) (2,12,12,8) (2,4,4,2)
NYC taxi – downtown 504 (2,19,19,8) (2,19,19,8) (2,4,4,2)
FMRI data 1452 (25,64,64) (1,64,64) (8,23,23)
Table 4: Comparisons of prediction accuracy (measured by MSE over the testing data) and computational time (in the format of hh:mm:ss) of different methods in real-world applications
MSE Time
Prediction Task TCN FATTNN Multiway TCN FATTNN Multiway
FAO crop $\sim$ crop 14.11 7.437 12.29 00:55:33 00:15:28 00:03:18
FAO $\sim$ livestock 32.53 21.65 48.04 00:24:34 00:12:57 00:03:21
NYC taxi – midtown 12.87 7.399 22.84 01:57:15 00:34:34 00:04:52
NYC taxi – downtown 14.58 9.493 28.10 03:25:55 00:32:46 00:25:35
FMRI data 0.1236 0.06634 0.1397 03:54:50 00:20:58 10:01:20
(1) The United Nations Food and Agriculture Organization (FAO) Crops and Livestock Products Data. The database provides agricultural statistics (including crop, livestock, and forestry sub-sectors) collected from countries and territories since 1961. It is publicly available at https://www.fao.org/faostat/en/#data/QCL. Our analyses focus on 11 crops and 5 livestock products of 46 countries in East Asia, North America, South America, and Europe from 1961 to 2022. Detailed information about the types of crops, livestock, and countries are presented in Table S.1 of Appendix B. We study two prediction tasks using the FAO dataset: (i) using Yield and Production data of 11 different crops for 33 countries in East Asia, North America, and Europe (i.e., ${\cal X}_{t}\in\mathbb{R}^{33\times 2\times 11}$) to predict the Yield and Production quantities of the same crops for 13 selected countries in South America (i.e., ${\cal Y}_{t}\in\mathbb{R}^{13\times 2\times 11}$); and (ii) using four agricultural statistics (Producing Animals, Animals-slaughtered, Milk Animals, Laying) associated with 5 kinds of livestock of 26 selected countries in Europe (i.e., ${\cal X}_{t}\in\mathbb{R}^{26\times 4\times 5}$) to predict three metrics (Area-harvested, Production, and Yield) of 11 crops in the same countries (i.e., ${\cal Y}_{t}\in\mathbb{R}^{26\times 3\times 11}$). Observing high variability in the raw data, we adopt a log transformation on the raw series and fit our models using log-transformed data.
Table 4 shows that FATTNN brings in substantial reductions in MSE compared to TCN and Multiway. In task (i), FATTNN yields 46.92% and 39.49% reductions in MSE compared to TCN and Multiway, respectively, and in task (ii), the improvements are 33.45% and 54.93% reductions in MSE, respectively. As reflected by the table, FATTNN computes much faster than TCN, about 3.6 times faster in task (i) and 2 times faster in tasks (ii). We also provide graphical comparisons between ground truth and predictions from different methods; see Figures 2 and 3. The plots confirm that FATTNN yields results that most closely align with the ground-truth patterns.
Figure 2: Comparisons between ground-truth values and predicted Production of Treenuts (top row) and Oilcrops (bottom row) in South America using the crop data of 33 countries in East Asia, North America, and Europe. Values are reported on the log-transformed scale. From left to right: Ground truth, FATTNN, and TCN.
Figure 3: Comparisons between ground-truth values and predicted Yield (top row) and Area-harvested (bottom row) of citrus fruit in 26 selected countries in Europe using livestock data of the same 26 countries. Values are plotted on the log-transformed scale. For readability, numerical numbers are reported in Tables S.2S.3 instead of being superimposed on the plots. From left to right: Ground truth, FATTNN, and TCN.
(2) New York City (NYC) Taxi Trip Data. The data contains 24-hour taxi pick-up and drop-off information of 69 areas in New York City for all the business days in 2014 and 2015. It is publicly available at https://www.nyc.gov/site/tlc/about/tlc-trip-record-data.page. Considering the autoregressive nature of this taxi data, we include the lag-1 response as an additional covariate when fitting the models. We focus on two prediction tasks: using the pick-up and drop-off data from 6:00-14:00 to predict the pick-up and drop-off outcomes from 14:00-22:00 in 12 districts in Midtown Manhattan and 19 districts in Downtown Manhattan, respectively. A detailed Manhattan district map is provided in Figure S.1 of Appendix B. The Midtown prediction task yields a tensor covariate and response $({\cal X}_{t},{\cal Y}_{t})\in\mathbb{R}^{2\times 12\times 12\times 8}\times\mathbb{R}^{2\times 12\times 12\times 8}$ for each business day, in which the 1st dimension represents two different taxi companies, the 2nd and 3rd dimensions encode the pick-up and drop-off districts, and the 4th dimension represents the hour of day. Similarly, the Downtown prediction task involves $({\cal X}_{t},{\cal Y}_{t})\in\mathbb{R}^{2\times 19\times 19\times 8}\times\mathbb{R}^{2\times 19\times 19\times 8}$ for each business day.
From Table 4, we observe that FATTNN significantly outperforms the other approaches in terms of prediction accuracy, with a 42.5% reduction in MSE compared to TCN and a 67.6% reduction in MSE compared to Multiway in Midtown traffic prediction. In Downtown traffic prediction, FATTNN shows a 34.9% reduction in MSE compared to TCN and a 66.2% reduction in MSE compared to Multiway. As shown in the table, FATTNN is much more computationally efficient than TCN, about 6.4 times faster than TCN in Downtown prediction and 3.4 times faster in Midtown prediction. Additionally, Figure 4 visualizes the comparisons using a 5-day moving average of ground-truth values and predicted pick-up and drop-off volumes by day (where we sum the number of pick-up and drop-off across 8-hour testing periods). The FATTNN-predicted lines appear to be closer to the ground-truth lines than TCN-predicted lines in most plots.
(3) Functional Magnetic Resonance Imaging (FMRI) Data. We consider the Haxby dataset (Haxby et al.,, 2001), a well-known public dataset for research in brain imaging and cognitive neuroscience, which can be retrieved from the “NiLearn” Python library. The data contains slices of $64\times 64$ FMRI brain images collected from 1452 samples. We use the first 25 slices of each sample to predict the 26th slice of that sample, i.e., $(\mathcal{X}_{t},\mathcal{Y}_{t})\in\mathbb{R}^{25\times 64\times 64}\times\mathbb{R}^{1\times 64\times 64}$. Similar to the previous two datasets, FATTNN significantly outperforms the other methods in both prediction accuracy and computational efficiency. Compared to TCN, FATTNNyields a 46.3% reduction in MSE and computes 11 times faster. To further illustrate the algorithm performances, we visualize the predicted images in Figure 5. It is evident that FATTNN produces images that more accurately resemble the ground truth.
In summary, our proposed methods, especially the FATTNN, achieve substantial increases in prediction accuracy compared to benchmark methods – with $\sim 40\%$ reductions in MSE compared to TCN, and even larger reductions compared to Multiway Regression. At the same time, the computational costs of FATTNN methods are much lower compared to TCN which directly feeds tensor observations into a convolutional network, ranging from 2 to 11 times faster than TCN.
Figure 4: Comparisons between ground-truth values and predicted pick-up and drop-off volumes using various methods. “District A to B” denotes the traffic volume that passengers were picked up in District A and dropped off in District B. The district numbering is assigned according to the Manhattan district map shown in Figure S.1. Left panel: Midtown Manhattan; Right panel: Downtown Manhattan.
Figure 5: Comparison between ground-truth values and predicted outputs for selected samples in FMRI dataset. From left to right: Ground truth, FATTNN, and TCN.
5 Summary & Discussion
In this paper, we propose FATTNN, a hybrid method that integrates a tensor factor model into deep neural networks, for tensor-on-tensor time series forecasting. The key advantage of FATTNN is its ability to exploit and utilize the underlying tensor structure among the covariate tensors. Through a low-rank tensor factor representation, we preserve the tensor structures and, in the meantime, drastically reduce the data dimensionality, leading to enhanced prediction results and reduced computational costs. This approach unifies the strengths of both factor models and neural networks. Numerical studies confirm that our proposed methods successfully achieve substantial increases in prediction accuracy and significant reductions in computational time, compared to traditional linear tensor-on-tensor regression models (which often fail to capture the intricate nonlinearity in covariate-response relationships) and conventional deep learning approaches (that often operate in a black-box manner without intention to making use of the inherent structures among observations).
Limitations. In this work, we focus on tensor-on-tensor time series forecasting. The integration of TIPUP and TCN, both designed specifically for handling temporal data, makes our proposed approach a powerful tool for forecasting tensor series over time. When the data are not time series, the TIPUP-TCN-based FATTNN is still applicable for tensor-on-tensor prediction tasks, but the strengths of TIPUP and TCN may not be fully utilized for modeling non-temporal data. This limitation can be easily rectified by adopting proper types of tensor factor models and neural networks in the implementation of FATTNN, depending on the nature of the prediction tasks and data types involved. Please see the paragraph on generalizations of FATTNN.
Generalizations. Though we presented our models for the purpose of tensor-on-tensor time series forecasting, our proposed FATTNN framework is not confined to temporal data but can accommodate a variety of data types. The main idea of FATTNN is to use a low-rank tensor factor model to capture the intrinsic patterns among the observed covariates, and then proceed with a neural network to model the intricate relationships between covariates and responses. With appropriate choices of factor models and neural network architectures, the proposed FATTNN can be naturally adapted to any simple (e.g., i.i.d observations) or complex tensor-type data (e.g., image, graph, network, text, video). When handling time series tensors, we adopt TIPUP factorization and TCN to fully capitalize on their temporal nature. To extend beyond time series data, different types of factor models and neural network architectures should be deployed, depending on the nature of the data. For example, for graph data, Graph Neural Network is a more suitable alternative to TCN. For text data, Transformer models could substitute for the TCN in our FATTNN framework. Our proposed FATTNN is flexible in accommodating different types of tensor factor models and deep learning architecture for tensor-on-tensor prediction using diverse data types.
Appendices
Appendix A Tensor Factor Models
A.1 Notation
For a vector $x=(x_{1},...,x_{p})^{\top}$, define $\|x\|_{q}=(x_{1}^{q}+...+x_{p}^{q})^{1/q}$, $q\geq 1$. For a matrix $A=(a_{ij})\in\mathbb{R}^{m\times n}$, write the SVD as $A=U\Sigma V^{\top}$, where $\Sigma=\text{diag}(\sigma_{1}(A),\sigma_{2}(A),...,\sigma_{\min\{m,n\}}(A))$, with singular values $\sigma_{\max}(A)=$ $\sigma_{1}(A)\geq\sigma_{2}(A)\geq\cdots\geq\sigma_{\min\{m,n\}}(A)\geq 0$ in descending order. The matrix spectral norm is denoted as $\|A\|_{2}=\sigma_{1}(A).$ Write $\sigma_{\min}(A)$ the smallest nontrivial singular value of $A$. For two sequences of real numbers $\{a_{n}\}$ and $\{b_{n}\}$, write $a_{n}=O(b_{n})$ (resp. $a_{n}\asymp b_{n}$) if there exists a constant $C$ such that $|a_{n}|\leq C|b_{n}|$ (resp. $1/C\leq a_{n}/b_{n}\leq C$) for all sufficiently large $n$, and write $a_{n}=o(b_{n})$ if $\lim_{n\to\infty}a_{n}/b_{n}=0$. Write $a_{n}\lesssim b_{n}$ (resp. $a_{n}\gtrsim b_{n}$) if there exist a constant $C$ such that $a_{n}\leq Cb_{n}$ (resp. $a_{n}\geq Cb_{n}$). Write $a\wedge b=\min\{a,b\}$ and $a\vee b=\max\{a,b\}$. We use $C,C_{1},c,c_{1},...$ to denote generic constants, whose actual values may vary from line to line.
For any two $m\times r$ matrices with orthonormal columns, say, $U$ and $\widehat{U}$, suppose the singular values of $U^{\top}\widehat{U}$ are $\sigma_{1}\geq\sigma_{2}\geq\cdots\geq\sigma_{r}\geq 0$. Define the principal angles between $U$ and $\widehat{U}$ as
$\Theta(U,\widehat{U})=\text{diag}(\arccos(\sigma_{1}),\arccos(\sigma_{2}),...,\arccos(\sigma_{r})).$
A natural measure of distance between the column spaces of $U$ and $\widehat{U}$ is then
$\|\widehat{U}\widehat{U}^{\top}-UU^{\top}\|_{2}=\|\sin\Theta(U,\widehat{U})\|_{2}=\sqrt{1-\sigma_{r}^{2}},$ (S.1)
which equals to the sine of the largest principle angle between the column spaces of $U$ and $\widehat{U}$. For any two matrices $A\in\mathbb{R}^{m_{1}\times r_{1}},B\in\mathbb{R}^{m_{2}\times r_{2}}$, denote the Kronecker product $\odot$ as $A\odot B\in\mathbb{R}^{m_{1}m_{2}\times r_{1}r_{2}}$. For any two tensors ${\cal A}\in\mathbb{R}^{m_{1}\times m_{2}\times\cdots\times m_{K}},{\cal B}\in\mathbb{R}^{r_{1}\times r_{2}\times\cdots\times r_{N}}$, denote the tensor product $\otimes$ as ${\cal A}\otimes{\cal B}\in\mathbb{R}^{m_{1}\times\cdots\times m_{K}\times r_{1}\times\cdots\times r_{N}}$, such that
$({\cal A}\otimes{\cal B})_{i_{1},...,i_{K},j_{1},...,j_{N}}=({\cal A})_{i_{1},...,i_{K}}({\cal B})_{j_{1},...,j_{N}}.$
Let ${\rm{vec}}(\cdot)$ be the vectorization of matrices and tensors. The mode-$k$ unfolding (or matricization) is defined as ${\rm{mat}}_{k}({\cal A})$, which maps a tensor ${\cal A}$ to a matrix ${\rm{mat}}_{k}({\cal A})\in\mathbb{R}^{m_{k}\times m_{-k}}$ where $m_{-k}=\prod_{j\neq k}^{K}m_{j}$. For example, if ${\cal A}\in\mathbb{R}^{m_{1}\times m_{2}\times m_{3}}$, then
$({\rm{mat}}_{1}({\cal A}))_{i,(j+m_{2}(k-1))}=({\rm{mat}}_{2}({\cal A}))_{j,(k+m_{3}(i-1))}=({\rm{mat}}_{3}({\cal A}))_{k,(i+m_{1}(j-1))}={\cal A}_{ijk}.$
For tensor ${\cal A}\in\mathbb{R}^{m_{1}\times m_{2}\times\cdots\times m_{K}}$, the Hilbert Schmidt norm or Frobenius norm is defined as
$\|{\cal A}\|_{{\rm F}}=\sqrt{\sum_{i_{1}=1}^{m_{1}}\cdots\sum_{i_{K}=1}^{m_{K}}({\cal A})_{i_{1},...,i_{K}}^{2}}.$
A.2 Preliminary of matrix and tensor algebra
Fact (Norm inequalities).
  • For any matrix $A,B$, $\|AB\|\geq\|A\|\cdot\sigma_{\min}(B)$.
  • For any matrix $A$, square invertible matrix $B$, $\|A\|\leq\sigma_{\max}(B)\cdot\|AB^{-1}\|$.
  • For any matrix $A$, and square diagonal invertible matrix $B$, $\sigma_{\min}(BA)\geq\sigma_{\min}(B)\cdot\sigma_{\min}(A)$.
Lemma A.1 (Weyl’s inequality).
Let $A,B\in\mathbb{R}^{m\times n}$ where $n\geq k$, then we have for any $i\in[k]$,
$|\sigma_{i}(A)-\sigma_{i}(B)|\leq\|A-B\|$ (S.2)
Lemma A.2 (Davis–Kahan $\sin\Theta$ theorem).
Let $\Sigma,\hat{\Sigma}\in\mathbb{R}^{p\times p}$ be symmetric, with eigenvalues $\lambda_{1}\geq\cdots\geq\lambda_{p}$ and $\hat{\lambda}_{1}\geq\cdots\geq\hat{\lambda}_{p}$ respectively. Fix $1\leq r\leq s\leq p$, let $d=s-r+1$, and let $V=(v_{r},v_{r+1},\ldots,v_{s})\in\mathbb{R}^{p\times d}$ and $\hat{V}=(\hat{v}_{r},\hat{v}_{r+1},\ldots,\hat{v}_{s})\in\mathbb{R}^{p\times d}$ have orthonormal columns satisfying $\Sigma v_{j}=\lambda_{j}v_{j}$ and $\hat{\Sigma}\hat{v}_{j}=\hat{\lambda}_{j}\hat{v}_{j}$ for $j=r,r+1,\ldots,s$. Write $\delta=\inf\{|\hat{\lambda}-\lambda|:\lambda\in[\lambda_{s},\lambda_{r}],\hat{\lambda}\in(-\infty,\hat{\lambda}_{s+1}]\cup[\hat{\lambda}_{r-1},\infty)\}$, where we define $\hat{\lambda}_{0}=-\infty$ and $\hat{\lambda}_{p+1}=\infty$, and assume that $\delta>0$. Then
$\|\sin\Theta(\hat{V},V)\|\leq\frac{\|\hat{\Sigma}-\Sigma\|}{\delta},$ (S.3)
[…]
$\|\sin\Theta(\hat{v}_{j},v_{j})\|_{2}\leq\frac{\|\hat{\Sigma}-\Sigma\|_{2}}{\min(|\hat{\lambda}_{j-1}-\lambda_{j}|,|\hat{\lambda}_{j+1}-\lambda_{j}|)}.$ (S.4)
A.3 Technical lemmas
Lemma A.3 ($\epsilon$-covering of matrix norms, Han et al., (2020)).
Let $d,d_{j},d_{*},r\leq d\wedge d_{j}$ be positive integers, $\epsilon>0$ and $N_{d,\epsilon}=\lfloor(1+2/\epsilon)^{d}\rfloor$.
[…]
$\sup_{M\in\mathbb{R}^{d},\|M\|\leq 1}\|f(M)\|_{*}\leq 2\max_{1\leq j\leq N_{d,1/2}}\|f(M_{j})\|_{*}.$
(ii) Given $\epsilon>0$, there exist $U_{j}\in\mathbb{R}^{d\times r}$ and $V_{j^{\prime}}\in\mathbb{R}^{d^{\prime}\times r}$ with $\|U_{j}\|_{2}\vee\|V_{j^{\prime}}\|_{2}\leq 1$ such that
$\max_{M\in\mathbb{R}^{d\times d^{\prime}},\|M\|_{2}\leq 1,\text{rank}(M)\leq r}\ \min_{j\leq N_{dr,\epsilon/2},j^{\prime}\leq N_{d^{\prime}r,\epsilon/2}}\|M-U_{j}V_{j^{\prime}}^{\top}\|_{2}\leq\epsilon.$
Consequently, for any linear mapping $f$ and norm $\|\cdot\|_{*}$ in the range of $f$,
$\sup_{M,\widetilde{M}\in\mathbb{R}^{d\times d^{\prime}},\|M-\widetilde{M}\|_{2}\leq\epsilon\atop{\|M\|_{2}\vee\|\widetilde{M}\|_{2}\leq 1\atop\text{rank}(M)\vee\text{rank}(\widetilde{M})\leq r}}\frac{\|f(M-\widetilde{M})\|_{*}}{\epsilon 2^{I_{r<d\wedge d^{\prime}}}}\leq\sup_{\|M\|_{2}\leq 1\atop\text{rank}(M)\leq r}\|f(M)\|_{*}\leq 2\max_{1\leq j\leq N_{dr,1/8}\atop 1\leq j^{\prime}\leq N_{d^{\prime}r,1/8}}\|f(U_{j}V_{j^{\prime}}^{\top})\|_{*}.$ (S.5)
(iii) Given $\epsilon>0$, there exist $U_{j,k}\in\mathbb{R}^{d_{k}\times r_{k}}$ and $V_{j^{\prime},k}\in\mathbb{R}^{d^{\prime}_{k}\times r_{k}}$ with $\|U_{j,k}\|_{2}\vee\|V_{j^{\prime},k}\|_{2}\leq 1$ such that
$\max_{M_{k}\in\mathbb{R}^{d_{k}\times d_{k}^{\prime}},\|M_{k}\|_{2}\leq 1\atop\text{rank}(M_{k})\leq r_{k},\forall k\leq K}\ \min_{j_{k}\leq N_{d_{k}r_{k},\epsilon/2}\atop j^{\prime}_{k}\leq N_{d_{k}^{\prime}r_{k},\epsilon/2},\forall k\leq K}\Big{\|}\odot_{k=2}^{K}M_{k}-\odot_{k=2}^{K}(U_{j_{k},k}V_{j_{k}^{\prime},k}^{\top})\Big{\|}_{\rm op}\leq\epsilon(K-1).$
For any linear mapping $f$ and norm $\|\cdot\|_{*}$ in the range of $f$,
$\sup_{M_{k},\widetilde{M}_{k}\in\mathbb{R}^{d_{k}\times d_{k}^{\prime}},\|M_{k}-\widetilde{M}_{k}\|_{2}\leq\epsilon\atop{\text{rank}(M_{k})\vee\text{rank}(\widetilde{M}_{k})\leq r_{k}\atop\|M_{k}\|_{2}\vee\|\widetilde{M}_{k}\|_{2}\leq 1\ \forall k\leq K}}\frac{\|f(\odot_{k=2}^{K}M_{k}-\odot_{k=2}^{K}\widetilde{M}_{k})\|_{*}}{\epsilon(2K-2)}\leq\sup_{M_{k}\in\mathbb{R}^{d_{k}\times d_{k}^{\prime}}\atop{\text{rank}(M_{k})\leq r_{k}\atop\|M_{k}\|_{2}\leq 1,\forall k}}\Big{\|}f\big{(}\odot_{k=2}^{K}M_{k}\big{)}\Big{\|}_{*}$ (S.6)
and
$\sup_{M_{k}\in\mathbb{R}^{d_{k}\times d_{k}^{\prime}},\|M_{k}\|_{2}\leq 1\atop\text{rank}(M_{k})\leq r_{k}\ \forall k\leq K}\Big{\|}f\big{(}\odot_{k=2}^{K}M_{k}\big{)}\Big{\|}_{*}\leq 2\max_{1\leq j_{k}\leq N_{d_{k}r_{k},1/(8K-8)}\atop 1\leq j_{k}^{\prime}\leq N_{d_{k}^{\prime}r_{k},1/(8K-8)}}\Big{\|}f\big{(}\odot_{k=2}^{K}U_{j_{k},k}V_{j_{k}^{\prime},k}^{\top}\big{)}\Big{\|}_{*}.$ (S.7)
Lemma A.4 (Slepian’s inequality, Vershynin, (2018)).
Let $(X_{t})_{t\in T}$ and $(Y_{t})_{t\in T}\in T$ be two mean zero Gaussian Processes. Assume that for all $t,s\in T$, we have
$\mathbb{E}X_{t}^{2}=\mathbb{E}Y_{t}^{2}\quad\text{and}\quad\mathbb{E}[(X_{t}-X_{s})^{2}]\leq\mathbb{E}[(Y_{t}-Y_{s})^{2}].$ (S.8)
Then for every $\tau\in R$, we have
$\mathbb{P}\left\{\sup_{t\in T}X_{t}\geq\tau\right\}\leq\mathbb{P}\left\{\sup_{t\in T}Y_{t}\geq\tau\right\}.$ (S.9)
Consequently,
$\mathbb{E}\sup_{t\in T}X_{t}\leq\mathbb{E}\sup_{t\in T}Y_{t}$ (S.10)
Lemma A.5 (Sudakov-Fernique inequality, Vershynin, (2018)).
Let $(X_{t})_{t\in T}$ and $(Y_{t})_{t\in T}$ be two mean-zero Gaussian processes. Assume that, for all $t,s\in T$, we have
$\mathbb{E}[(X_{t}-X_{s})^{2}]\leq\mathbb{E}[(Y_{t}-Y_{s})^{2}].$ (S.11)
Then,
$\mathbb{E}\left[\sup_{t\in T}\{X_{t}\}\right]\leq\mathbb{E}\left[\sup_{t\in T}\{Y_{t}\}\right].$ (S.12)
A function $f:\mathbb{R}^{d}\to\mathbb{R}$ is $L$-Lipschitz continuous (sometimes called globally Lipschitz continuous) if there exists $L\in\mathbb{R}$ such that
$\|f(x)-f(y)\|\leq L\|x-y\|.$
The smallest constant $L>0$ is denoted $\|f\|_{\rm Lip}$.
Lemma A.6 (Gaussian Concentration of Lipschitz functions, Vershynin, (2018)).
Let $X\sim N(0,I_{d})$ and let $f:\mathbb{R}^{d}\to\mathbb{R}$ be a Lipschitz function. Then,
$\mathbb{P}\left(f(X)-\mathbb{E}f(X)\geq t\right)\leq 2\exp\left(-\frac{t^{2}}{2\|f\|_{\rm Lip}}\right).$
Lemma A.7 (Sigular values of Gaussian random matrices).
Let $X$ be an $n\times p$ matrix with i.i.d. $N(0,1)$ entries. For any $x>0$,
$\mathbb{P}\left(\left\|X^{\top}X-nI_{p}\right\|_{2}\geq p+2\sqrt{pn}+2x(\sqrt{p}+\sqrt{n})+x^{2}\right)\leq 2e^{-x^{2}/2}.$
Proof.
It can be easily derived by Corollary 5.35 in Vershynin, (2010). ∎
A.4 The Determination of Rank
Here, the estimators are constructed with given ranks $r_{1},\ldots,r_{K}$. Determining the number of factors in a data-driven manner has been a significant research topic in the factor model literature. Recently, Han et al., (2022) established a class of rank determination approaches for tensor factor models, based on both the information criterion and the eigen-ratio criterion. These procedures can be adopted for our purposes.
A.5 Tucker Decomposition and iterative TIPUP (iTIPUP) Algorithm
The Higher-Order Orthogonal Iteration (HOOI) algorithm De Lathauwer et al., (2000) is one of the most popular methods for computing the Tucker decomposition, which relies on the orthogonal projection and Singular Value Decomposition (SVD) to update the estimation of loading matrices for each tensor mode in every iteration.
Similar to the idea of HOOI, an iterative refinement version of TIPUP could further improve its performance, just as HOOI enhances the original Tucker decomposition method. The pseudo-code for this iterative refinement is provided in Algorithm S.1.
Algorithm S.1 Iterative TIPUP (iTIPUP) algorithm
1:  Input: $\mathcal{X}_{t}\in\mathbb{R}^{d_{1}\times\ldots\times d_{K}}$ for $t=1,\ldots,n$, rank $r_{k}$ for all $k=1,\ldots,K$, the tolerance parameter $\epsilon>0$, the maximum number of iterations $J$.
2:  Let $j=0$, initiate via TIPUP on ${\cal X}_{1},...,{\cal X}_{n}$ to obtain
$\widehat{A}_{k}^{(0)}=\text{LSVD}_{r_{k}}\left(\frac{1}{n}\sum_{t=1}^{n}\hbox{\rm mat}_{k}({\cal X}_{t})\hbox{\rm mat}_{k}^{\top}({\cal X}_{t})\right),$
[…]
3:  repeat
4:     Let $j=j+1$.
5:     for $k=1,\ldots,K$ do
6:        Given previous estimates $\widehat{A}_{1}^{(j-1)},\ldots,\widehat{A}_{k-1}^{(j-1)},\widehat{A}_{k+1}^{(j-1)},\ldots,\widehat{A}_{K}^{(j-1)}$, sequentially calculate,
$Z_{t,k}^{(j)}={\cal X}_{t}\times_{1}\widehat{A}_{1}^{(j)\top}\times_{2}\ldots\times_{k-1}\widehat{A}_{k-1}^{(j)\top}\times_{k+1}\widehat{A}_{k+1}^{(j-1)\top}\times_{k+2}\ldots\times_{K}\widehat{A}_{K}^{(j-1)\top},$
for $t=1,\ldots,n$.
7:        Perform TIPUP on the new tensor series $(Z_{1,k}^{(j)},\ldots,Z_{n,k}^{(j)})$,
$\widehat{A}_{k}^{(j)}=\text{LSVD}_{r_{k}}\left(\frac{1}{n}\sum_{t=1}^{n}\hbox{\rm mat}_{k}(Z_{t,k}^{(j)})\hbox{\rm mat}_{k}^{\top}(Z_{t,k}^{(j)})\right).$
8:     end for
9:  until $j=J$ or $\max_{1\leq k\leq K}\|\widehat{A}_{k}^{(j)}\widehat{A}_{k}^{(j)\top}-\widehat{A}_{k}^{(j-1)}\widehat{A}_{k}^{(j-1)\top}\|_{2}\leq\epsilon$.
10:  Output:
$\widehat{A}_{k}$ $=\widehat{A}_{k}^{(j)},\quad k=1,\ldots,K,$
$\widehat{F}_{t}$ $=\mathcal{X}_{t}\times_{k=1}^{K}\widehat{A}_{K}^{\top},\quad t=1,\ldots,n.$
A.6 Proof of Proposition 3.1
We focus on the case of $K=2$ as the iTIPUP begins with mode-$k$ matrix unfolding. In particular, we sometimes give explicit expressions only in the case of $k=1$ and $K=2$. For $K=2$, we observe a matrix time series with $X_{t}=A_{1}F_{t}A_{2}^{\top}+{\cal E}_{t}\in\mathbb{R}^{d_{1}\times d_{2}}$. Let $\overline{\mathbb{E}}(\cdot)=\mathbb{E}(\cdot|\{{\cal F}_{1},...,{\cal F}_{n}\})$ and ¯()=(|$\overline{\mathbb{P}}(\cdot)=\mathbb{P}(\cdot|\{{\cal F}_{1},...,{\cal F}_{n}\})$.
Let $L_{k}^{(m)}$ be the loss for $\hat{A}_{k}^{(m)}$ at the $m$-th iteration,
$L_{k}^{(m)}=\|\widehat{P}_{k}^{(m)}-P_{k}\|_{2},\quad L^{(m)}=\max_{1\leq k\leq K}L_{k}^{(m)},$
where $\widehat{P}_{k}^{(m)}=A_{k}^{(m)}(A_{k}^{(m)\top}A_{k}^{(m)})^{-1}A_{k}^{(m)^{\top}}$.
We outline the proof as follows.
(i) Initialization.
Define
${\rm TIPUP}_{1}=\frac{1}{n}\sum_{t=1}^{n}X_{t}X_{t}^{\top}\in\mathbb{R}^{d_{1}\times d_{1}},$
The “noiseles” version of TIPUP1 is
$\Theta_{1}=\frac{1}{n}\sum_{t=1}^{n}A_{1}F_{t}A_{2}^{\top}A_{2}F_{1}^{\top}A_{1}^{\top}\in\mathbb{R}^{d_{1}\times d_{1}},$
which is of rank $r_{1}$. Then, by Davis-Kahan sin $\Theta$ theorem (Lemma A.2),
$L_{1}^{(0)}=\|P_{1}-\widehat{P}_{1}^{(0)}\|_{2}\leq\frac{2\|{\rm TIPUP}_{1}-\Theta_{1}\|_{2}}{\sigma_{r_{1}}(\Theta_{1})}.$ (S.13)
To bound the numerator on the right hand side of (S.13), we write
${\rm TIPUP}_{1}-\Theta_{1}=\frac{1}{n}\sum_{t=1}^{n}A_{1}F_{t}A_{2}^{\top}{\cal E}_{t}^{\top}+\frac{1}{n}\sum_{t=1}^{n}{\cal E}_{t}A_{2}{\cal F}_{t}^{\top}A_{1}^{\top}+\frac{1}{n}\sum_{t=1}^{n}{\cal E}_{t}{\cal E}_{t}^{\top}:=\Delta_{1}+\Delta_{2}+\Delta_{3}.$
For $\Delta_{3}$, by Lemma A.7,
$\mathbb{P}\left(\left\|\sum_{t=1}^{n}\mathcal{E}_{t}\mathcal{E}_{t}^{\top}-n\mathbb{E}\mathcal{E}_{t}\mathcal{E}_{t}^{\top}\right\|_{2}\geq d_{1}+2\sqrt{d_{1}d_{2}n}+2x(\sqrt{d_{1}}+\sqrt{d_{2}n})+x^{2}\right)\leq 2e^{-x^{2}/2}.$
By setting $x\asymp\sqrt{d_{1}}$, with probability at least $1-e^{-d_{1}}$,
$\left\|\Delta_{3}-\mathbb{E}\mathcal{E}_{t}\mathcal{E}_{t}^{\top}\right\|_{2}\lesssim\frac{d_{1}}{n}+\sqrt{\frac{d_{1}d_{2}}{n}}.$
Next consider $\Delta_{1}$ and $\Delta_{2}$. For any $u$ and $v$ in $\mathbb{R}^{d_{1}}$, we have
$\overline{\mathbb{E}}\left(u^{\top}\left(\frac{1}{\sqrt{n}}\sum_{t=1}^{n}A_{1}F_{t}A_{2}^{\top}\mathcal{E}_{t}^{\top}\right)v\right)^{2}\leq\|\Theta_{1}\|_{2}\|u\|_{2}^{2}\|v\|_{2}^{2}.$
Thus for $u_{i}$ and $v_{i}$ with $\|u_{i}\|_{2}=\|v_{i}\|_{2}=1$, $i=1,2$,
$n\|\Theta_{1}\|_{2}^{-1}\overline{\mathbb{E}}\left((u_{1}^{\top}\Delta_{1}v_{1}-u_{2}^{\top}\Delta_{1}v_{2})^{2}\right)$ $\leq\left(\|u_{1}-u_{2}\|_{2}\|v_{1}\|_{2}+\|u_{2}\|_{2}\|v_{1}-v_{2}\|_{2}\right)^{2}$
$\leq 2\left(\|u_{1}-u_{2}\|_{2}^{2}+\|v_{1}-v_{2}\|_{2}^{2}\right)$
$=2\mathbb{E}\left[(u_{1}-u_{2})^{\top}\xi+(v_{1}-v_{2})^{\top}\zeta\right]^{2},$
where $\xi$ and $\zeta$ are i.i.d $N(0,I_{d_{1}})$ vectors. As $\Delta_{1}$ is a $d_{1}\times d_{1}$ Gaussian matrix under $\overline{\mathbb{E}}$, the Sudakov-Fernique inequality (Lemma A.5) yields
$\sqrt{n}\|\Theta_{1}\|_{2}^{-1/2}\overline{\mathbb{E}}\|\Delta_{1}\|_{2}\leq\sqrt{2}\mathbb{E}\sup_{\|u\|_{2}=\|v\|_{2}=1}(u^{\top}\xi+v^{\top}\zeta)=\sqrt{2}\mathbb{E}\left(\|\xi\|_{2}+\|\zeta\|_{2}\right)\leq 2\sqrt{2d_{1}}.$
It follows that
$\overline{\mathbb{E}}\|\Delta_{1}\|_{2}\leq\sqrt{\frac{8d_{1}}{n}}\|\Theta_{1}\|_{2}^{1/2}.$
Elementary calculation shows that $\|\Delta_{1}\|_{2}$ is a $\sqrt{n}\|\Theta_{1}\|_{2}^{1/2}$-Lipschitz function. Then, by Gaussian concentration inequality for Lipschitz functions (Lemma A.6),
¯$\overline{\mathbb{P}}\left(\|\Delta_{1}\|_{2}\geq\sqrt{\frac{8d_{1}}{n}}\|\Theta_{1}\|_{2}^{1/2}+\sqrt{\frac{1}{n}}\|\Theta_{1}\|_{2}^{1/2}x\right)\leq 2e^{-\frac{x^{2}}{2}}.$
With $x\asymp\sqrt{d_{1}}$, with probability at least $1-e^{-d_{1}}$,
$\|\Delta_{2}\|_{2}=\|\Delta_{1}\|_{2}\lesssim\sqrt{\frac{d_{1}}{n}}\sigma_{1}(\Theta_{1}).$
As $\sigma_{1}(\Theta_{1})=\sigma_{r_{1}}(\Theta_{1})\asymp\lambda^{2}$, by (S.13), with probability at least $1-e^{-d_{1}}$,
$L_{1}^{(0)}=\|\widehat{P}_{1}^{(0)}-P_{1}\|_{2}\lesssim\frac{1}{\lambda^{2}}\sqrt{\frac{d}{n}}+\frac{d_{1}}{\lambda^{2}n}+\frac{1}{\lambda}\sqrt{\frac{d_{1}}{n}}.$ (S.14)
(ii) Iterative refinement.
[…]
${\rm TIPUP_{1}}(U_{2})=\frac{1}{n}\sum_{t=1}^{n}X_{t}U_{2}U_{2}^{\top}X_{t}^{\top}\in\mathbb{R}^{d_{2}\times d_{2}},$
for any matrix-valued variable $U_{2}\in\mathbb{R}^{d_{2}\times r_{2}}$. Given $\widehat{A}_{2}^{(m)}$, the $(m+1)$-th iteration produces estimates
$\widehat{A}_{1}^{(m+1)}=\text{LSVD}_{r_{1}}(\text{TIPUP}_{1}(\widehat{A}_{2}^{(m)})),\quad\widehat{P}_{1}^{(m+1)}=\widehat{A}_{1}^{(m+1)}\widehat{A}_{1}^{(m+1)^{\top}}.$ (S.15)
The “noiseless” version of this update is given by
$\Theta_{1}(U_{2})=\frac{1}{n}\sum_{t=1}^{n}A_{1}F_{t}A_{2}^{\top}U_{2}U_{2}^{\top}A_{2}F_{t}^{\top}A_{1}^{\top}=\Theta_{1}$
giving error-free “estimates”,
$\widehat{A}_{1}^{(m+1)}=\text{LSVD}_{r_{1}}(\Theta_{1}(\widehat{A}_{2}^{(m)})),$
where $\Theta_{1}(\widehat{A}_{2}^{(m)})$ is of rank $r_{1}$. Thus, by Davis-Kahan sin $\Theta$ theorem (Lemma A.2)
$L_{1}^{(m+1)}=\|\widehat{P}_{1}^{(m+1)}-P_{1}\|_{2}\leq\frac{2\|\text{TIPUP}_{1}(\widehat{A}_{2}^{(m)})-\Theta_{1}(\widehat{A}_{2}^{(m)})\|_{2}}{\sigma_{r_{1}}[\Theta_{1}(\widehat{A}_{2}^{(m)})]}.$ (S.16)
To bound the numerator on the right-hand side of (S.16), we write
$\text{TIPUP}_{1}(U_{2})-\Theta_{1}(U_{2}):=\Delta_{1}(U_{2}U_{2}^{\top})+\Delta_{2}(U_{2}U_{2}^{\top})+\Delta_{3}(U_{2}U_{2}^{\top}),$
where for any $M_{2}\in\mathbb{R}^{d_{2}\times d_{2}}$,
$\Delta_{1}(M_{2})$ $=\frac{1}{n}\sum_{t=1}^{n}A_{1}F_{t}A_{2}^{\top}M_{2}\mathcal{E}_{t}^{\top},$
$\Delta_{2}(M_{2})$ $=\frac{1}{n}\sum_{t=1}^{n}\mathcal{E}_{t}M_{2}A_{2}F_{t}^{\top}A_{1}^{\top},$
$\Delta_{3}(M_{2})$ $=\frac{1}{n}\sum_{t=1}^{n}\mathcal{E}_{t}M_{2}\mathcal{E}_{t}^{\top}.$
As $\Delta_{i}(M_{2})$ is linear in $M_{2}$, the numerator on the right-hand side of (S.16) can be bounded by
$\|\text{TIPUP}_{1}(\widehat{A}_{2}^{(m)})-\Theta_{1}(\widehat{A}_{2}^{(m)})\|_{2}\leq\|\text{TIPUP}_{1}(A_{2})-\Theta_{1}(A_{2})\|_{2}+L^{(m)}(2K-2)\sum_{i=1}^{3}\|\Delta_{i}(M_{2})\|_{\rm S},$
where
$\|\Delta_{i}(M_{2})\|_{\rm S}=\max_{\|M_{2}\|_{1}\leq 1,\text{rank}(M_{2})\leq r_{2}}\|\Delta_{i}(M_{2})\|_{2}.$
We claim in certain events $\Omega_{i}$, with $\mathbb{P}(\Omega_{i})\geq 1-e^{-d_{1}}$,
$\|\Delta_{i}(M_{2})\|_{\rm S}$ $\lesssim\sqrt{\frac{d_{1}+d_{2}}{n}}\lambda,\quad i=1,2,$ (S.17)
$\|\Delta_{3}(M_{2})\|_{\rm S}$ $\lesssim\frac{d_{1}+d_{2}}{n}+\sqrt{\frac{d_{1}+d_{2}}{n}}.$ (S.18)
We will also prove by induction, the denominator in (S.16) satisfies
$2\sigma_{r_{1}}[\Theta_{1}(\widehat{A}_{2}^{(m)})]\geq\sigma_{r_{1}}[\Theta_{1}(A_{2})].$ (S.19)
Define the ideal version of the ratio in (S.16) as
$L_{k}^{(\text{ideal})}=\frac{\|\text{TIPUP}_{1}(\odot_{j\neq k}A_{j})-\Theta_{k}(\odot_{j\neq k}A_{j})\|_{2}}{\sigma_{r_{1}}[\Theta_{k}(\odot_{j\neq k}A_{j})]},\quad L^{(\text{ideal})}=\max_{1\leq k\leq K}L_{k}^{(\text{ideal})}.$
The proof of (S.17) and (S.18) implies that, in an event with probability at least $1-\frac{k}{2}e^{-d_{k}}$,
$L_{k}^{(\text{ideal})}\lesssim\frac{1}{\lambda^{2}}\sqrt{\frac{d_{k}}{n}}+\frac{d_{k}}{\lambda^{2}n}+\frac{1}{\lambda}\sqrt{\frac{d_{k}}{n}}.$ (S.20)
Putting together (S.13),(S.14),(S.16),(S.17),(S.18) and (S.19), we have
$L_{k}^{(m+1)}\leq L_{k}^{(\text{ideal})}+L^{(m)}\rho,$
with $\rho<1$. By induction,
$L^{(m+1)}\leq(1+\rho+\dots+\rho^{m})L^{(m)}+\rho^{m+1}L^{(0)}.$
We achieve the desired results.
We divided the rest of the proof into 4 steps to prove (S.17) and (S.18)for $i=1,2,3$ and (S.19).
Step 1.
We prove (S.17) for the $\|\Delta_{1}(M_{2})\|_{\rm S}$.
By Lemma A.3 (ii), there exists $\widetilde{M}^{(\ell,\ell^{\prime})}\in\mathbb{R}^{d_{2}\times d_{2}}$ of the forms $W_{\ell}Q_{\ell^{\prime}}^{\top}$ with $W_{\ell}\in\mathbb{R}^{d_{2}\times r_{2}}$, $Q_{\ell^{\prime}}\in\mathbb{R}^{d_{2}\times r_{2}}$, $1\leq\ell,\ell^{\prime}\leq N_{d_{2}r_{2},1/8}=17^{d_{2}r_{2}}$, such that $\|\widetilde{M}^{(\ell,\ell^{\prime})}\|_{2}\leq 1$, $\text{rank}(\widetilde{M}^{(\ell,\ell^{\prime})})\leq r_{2}$ and
$\|\Delta_{1}(M_{2})\|_{\rm S}\leq 2\max_{\ell,\ell^{\prime}\leq N_{d_{2}r_{2},1/8}}\left\|\frac{1}{n}\sum_{t=1}^{n}A_{1}F_{t}A_{2}^{\top}\widetilde{M}^{(\ell,\ell^{\prime})}\mathcal{E}_{t}^{\top}\right\|_{2}.$
Elementary calculation shows that $\|\Delta_{1}(\widetilde{M}^{(\ell,\ell^{\prime})})\|_{2}$ is a $\sqrt{n}\|\Theta_{1}\|_{2}$-Lipschitz function of $\mathcal{E}_{1},\ldots,\mathcal{E}_{n}$. Employing similar argument as the initialization (Sudakov-Fernique inequality, Lemma A.5), we have
$\overline{\mathbb{E}}\left\|\frac{1}{n}\sum_{t=1}^{n}A_{1}F_{t}A_{2}^{\top}\widetilde{M}^{(\ell,\ell^{\prime})}\mathcal{E}_{t}^{\top}\right\|_{2}\leq\sqrt{\frac{8d_{1}}{n}}\|\Theta_{1}\|_{2}^{1/2}.$
Then, by Gaussian concentration inequality for Lipschitz functions (Lemma A.6),
¯$\overline{\mathbb{P}}\left(\left\|\frac{1}{n}\sum_{t=1}^{n}A_{1}F_{t}A_{2}^{\top}\widetilde{M}^{(\ell,\ell^{\prime})}\mathcal{E}_{t}^{\top}\right\|_{2}-\sqrt{\frac{8d_{1}}{n}}\|\Theta_{1}\|_{2}^{1/2}\geq\frac{1}{\sqrt{n}}\|\Theta_{1}\|_{2}^{1/2}x\right)\leq e^{-x^{2}/2}.$
Hence,
¯$\overline{\mathbb{P}}\left(\|\Delta_{1}(M_{2})\|_{\rm S}\geq\sqrt{\frac{8d_{1}}{n}}\|\Theta_{1}\|_{2}^{1/2}+\frac{1}{\sqrt{n}}\|\Theta_{1}\|_{2}^{1/2}x\right)\leq N_{d_{2}r_{2},1/8}^{2}e^{-x^{2}/2}.$
This implies that with $x\asymp\sqrt{d_{2}r_{2}}$ that in an event with probability at least $1-e^{-d_{2}}$,
$\|\Delta_{1}(M_{2})\|_{\rm S}\lesssim\sqrt{\frac{d_{1}+d_{2}r_{2}}{n}}\lambda\lesssim\sqrt{\frac{d_{1}+d_{2}}{n}}\lambda.$
Step 2.
The bound of $\|\Delta_{2}(M_{2})\|_{\rm S}$ follows from the same argument as the above step.
Step 3.
Here we prove (S.18) for $\|\Delta_{3}(M_{2})\|_{\rm S}$. By Lemma A.3(iii), we can find $U_{2}^{(\ell)}\in\mathbb{R}^{d_{2}\times r_{2}}$, $1\leq\ell\leq N_{d_{2}r_{2},1/8}$ such that $\|U_{2}^{(\ell)}\|_{2}\leq 1$ and
$\|\Delta_{3}(M_{2})\|_{\rm S}\leq 2\max_{1\leq\ell\leq N_{d_{2}r_{2},1/8}}\left\|\frac{1}{n}\sum_{t=1}^{n}\mathcal{E}_{t}U_{2}^{(\ell)}U_{2}^{(\ell)\top}\mathcal{E}_{t}^{\top}\right\|_{2}.$
By Lemma A.7,
$\mathbb{P}\left(\left\|\sum_{t=1}^{n}\mathcal{E}_{t}U_{2}^{(\ell)}U_{2}^{(\ell)\top}\mathcal{E}_{t}^{\top}-n\mathbb{E}\mathcal{E}_{t}U_{2}^{(\ell)}U_{2}^{(\ell)\top}\mathcal{E}_{t}^{\top}\right\|_{2}\geq d_{1}+2\sqrt{d_{1}r_{2}n}+2x(\sqrt{d_{1}}+\sqrt{r_{2}n})+x^{2}\right)\leq 2e^{-x^{2}/2}.$
Hence, by setting $x\asymp\sqrt{d_{2}r_{2}}$, we have with probability at least $1-e^{-d_{2}}$,
$\|\Delta_{3}(M_{2})\|_{\rm S}\lesssim\frac{d_{1}}{n}+\sqrt{\frac{d_{1}r_{2}}{n}}+\frac{d_{2}r_{2}}{n}+\sqrt{\frac{d_{2}r_{2}^{2}}{n^{2}}}\lesssim\frac{d_{1}+d_{2}}{n}+\sqrt{\frac{d_{1}+d_{2}}{n}}.$
Step 4.
Next, we prove (S.19). Note that
$\|\Theta_{1}(\widehat{A}_{2}^{(m)})-\Theta_{1}(A_{2})\|_{2}$ $\leq\left\|\frac{1}{n}\sum_{t=1}^{n}A_{1}F_{t}A_{2}^{\top}(\widehat{A}_{2}^{(m)}\widehat{A}_{2}^{(m)\top}-A_{2}A_{2}^{\top})A_{2}F_{t}^{\top}A_{1}^{\top}\right\|_{2}$
$\leq\|\widehat{A}_{2}^{(m)}\widehat{A}_{2}^{(m)\top}-A_{2}A_{2}^{\top}\|_{2}\|\Theta_{1}\|_{2}$
$\leq L^{(m)}\|\Theta_{1}\|_{2}.$
Thus, by Weyl’s inequality (Lemma A.1),
$\sigma_{r_{1}}[\Theta_{1}(\widehat{A}_{2}^{(m)})]\geq\sigma_{r_{1}}[\Theta_{1}(A_{2})]-L^{(m)}\|\Theta_{1}\|_{2}\geq\frac{1}{2}\sigma_{r_{1}}[\Theta_{1}(A_{2})],$
when $\min_{1\leq k\leq K}\frac{\sigma_{r_{k}}[\Theta_{k}(A_{2})]}{\|\Theta_{k}(A_{2})\|_{2}}\geq 2L^{(m)}$. We prove this condition by induction.
Appendix B Supplemental Results to Numerical Studies in Section 4
B.1 Supplemental Notation used in Simulation Studies
(1) For vectors $a_{1},\ldots,a_{K}$ of length $r_{1},\ldots,r_{K}$, the outer product
${\cal A}=a_{1}\circ a_{2}\circ\cdots\circ a_{K}\in\mathbb{R}^{r_{1}\times\cdots\times r_{K}}$
is the $K$-th order tensor, with entries ${\cal A}[i_{1},\ldots,i_{K}]=\prod_{k=1}^{K}a_{k}[i_{k}]$.
(2) Let $M_{1}\in\mathbb{R}^{r_{1}\times R},\ldots M_{K}\in\mathbb{R}^{r_{K}\times R}$ be $K$ matrices with the same column dimension (denoted by $R$), the notation $\llbracket M_{1},\cdots,M_{K}\rrbracket$ represents
$\llbracket M_{1},\ldots,M_{K}\rrbracket=\sum_{r=1}^{R}m_{1r}\circ\cdots\circ m_{Kr},$ (S.21)
where $m_{kr}$ is the $r$th column of $M_{k}$, $k=1,\ldots,K$, $r=1,\ldots,R$.
(3) For two tensor $\mathcal{A}:d_{1}\times\cdots\times d_{K}\times r_{1}\times\cdots\times r_{K}$ and $\mathcal{B}:r_{1}\times\cdots\times r_{K}\times p_{1}\times\cdots\times p_{q}$, the notation $\langle{\cal A},{\cal B}\rangle_{L}$ denote the contracted tensor product (Lock,, 2018)
$\langle\mathcal{A},\mathcal{B}\rangle_{L}\in\mathbb{R}^{d_{1}\times\cdots\times d_{K}\times p_{1}\times\cdots\times p_{q}}$
with the $(i_{1},\ldots,i_{K},j_{1},\ldots,j_{d})$-th element of $\langle\mathcal{A},\mathcal{B}\rangle_{L}$ defined as
$\langle\mathcal{A},\mathcal{B}\rangle_{L}[i_{1},\ldots,i_{K},j_{1},\ldots,j_{d}]=\sum_{l_{1}=1}^{r_{1}}\cdots\sum_{l_{K}=1}^{r_{K}}\mathcal{A}[i_{1},\dots,i_{K},l_{1},\dots,l_{K}]\mathcal{B}[l_{1},\dots,l_{K},j_{1},\dots,j_{d}].$ (S.22)
B.2 Supplemental Results for FAO Data Analysis
Table S.1 illustrates detailed information about the lists of crop types, livestock types, countries, and metrics involved in the two prediction tasks using FAO data:
  • (i)
    (crop $\sim$ crop) using Yield and Production data of 11 different crops for 33 countries in East Asia, North America, and Europe (i.e., ${\cal X}_{t}\in\mathbb{R}^{33\times 2\times 11}$) to predict the Yield and Production quantities of the same crops for 13 countries in South America (i.e., ${\cal Y}_{t}\in\mathbb{R}^{13\times 2\times 11}$);
  • (ii)
    (crop $\sim$ livestock) using four agricultural statistics (Producing Animals, Animals-slaughtered, Milk Animals, Laying) associated with 5 kinds of livestock of 26 selected countries in Europe (i.e., ${\cal X}_{t}\in\mathbb{R}^{26\times 4\times 5}$) to predict three metrics (Area-harvested, Production, and Yield) of 11 crops in the same countries (i.e., ${\cal Y}_{t}\in\mathbb{R}^{26\times 3\times 11}$).
Table S.2 and Table S.3 supplement Figure 3 by providing the numerical values of ground truth and predicted Yield metrics for the plot.
Table S.1: Details of statistics and metrics involved in the FAO data analysis
(a) Prediction Task (i): crop $\sim$ crop
$\mathcal{X}_{t}\in\mathbb{R}^{33\times 2\times 11}$ $\mathcal{Y}_{t}\in\mathbb{R}^{13\times 2\times 11}$
Countries AUT, BGR, CAN, CHN, HRV, CYP, CZE, PRK, DNK, EST, FIN, FRA, DEU, GRC, HUN, IRL, ITA, JPN, LVA, LTU, LUX, MLT, MNG, NLD, POL, PRT, KOR, ROU, SVK, SVN, ESP, SWE, USA (33 countries) ARG, BOL, BRA, CHL, COL, ECU, GUF, GUY, PRY, PER, SUR, URY, VEN (13 countries)
Crop/Livestock selected Cereals, Citrus Fruit, Fibre Crop, Fruit, OliCrops cake equivalent, OliCrops oil equivalent, Pulses, Roots and tubers, Sugar crop, Treenuts  (11 types) the same as in $\mathcal{X}_{t}$
Metrics involved Yield, Production (2 metrics) the same as in $\mathcal{X}_{t}$
(b) Prediction Task (i): crop $\sim$ livestock
$\mathcal{X}_{t}\in\mathbb{R}^{26\times 4\times 5}$ $\mathcal{Y}_{t}\in\mathbb{R}^{26\times 3\times 11}$
Countries AUT, BGR, HRV, CYP, CZE, DNK, EST, FIN, FRA, DEU, GRC, HUN, IRL, ITA, LVA, LTU, LUX, MLT, NLD, POL, PRT, ROU, SVK, SVN, ESP, SWE (26 countries) the same as in $\mathcal{X}_{t}$
Crop/Livestock selected Beef and Buffalo Meat, Eggs, Poultry Meat, Milk, Sheep and Goat Meat (5 types) Cereals, Citrus Fruit, Fibre Crop, Fruit, OliCrops cake equivalent, OliCrops oil equivalent, Pulses, Roots and tubers, Sugar crop, Treenuts  (11 types)
Metrics involved Producing Animals, Animals-slaughtered, Milk Animals, Laying (4 metrics) Yield, Production, Area-harvested (3 metrics)
Note: A full description of the code definitions is available on the United Nations Food and Agriculture Organization Crops and Livestock Products Database website.
Table S.2: Numerical values of ground-truth and predicted Yield of cirtus fruit in 26 European countries using livestock data of the same 26 countries - a supplement to Figure 3
Country Ground Truth FATTNN TCN ($\mathcal{Y}\sim\mathcal{X}$)
Austria 13.47 13.94 15.92
Bulgaria 12.14 15.01 17.34
Croatia 11.98 12.52 2.38
Cyprus 11.52 14.60 14.80
Czechia 13.38 10.68 11.13
Denmark 11.52 9.43 5.97
Estonia 13.29 13.06 17.61
Finland 15.86 16.91 14.45
France 16.19 14.86 11.88
Germany 13.27 15.24 13.31
Greece 12.87 14.67 16.19
Hungary 12.78 14.21 12.57
Ireland 14.15 14.87 14.59
Italy 15.07 14.31 13.43
Latvia 13.72 13.56 8.04
Lithuania 12.68 9.81 4.78
Luxembourg 12.77 9.49 3.20
Malta 9.76 -0.76 -0.49
Netherlands 12.12 12.44 8.33
Poland 15.74 15.40 18.19
Portugal 15.86 16.34 25.16
Romania 14.76 16.14 20.95
Slovakia 11.99 8.17 -0.96
Slovenia 11.36 8.58 0.84
Spain 14.62 9.52 17.50
Sweden 13.63 10.19 16.74
Note: Values are reported on the log-transformed scale.
Table S.3: Numerical values of ground-truth and predicted Area-harvested of cirtus fruit in 26 European countries using livestock data of the same 26 countries - a supplement to Figure 3
Country True Value FATTNN TCN ($\mathcal{Y}\sim\mathcal{X}$)
Austria 10.01 10.74 13.54
Bulgaria 9.36 11.81 13.65
Croatia 9.18 10.03 0.33
Cyprus 8.39 12.07 11.58
Czechia 10.07 9.54 7.65
Denmark 8.55 7.98 1.90
Estonia 9.99 10.53 13.61
Finland 12.14 14.00 11.41
France 12.44 11.48 9.40
Germany 10.07 11.60 10.08
Greece 9.65 12.03 12.86
Hungary 9.14 11.37 8.96
Ireland 10.84 12.24 9.91
Italy 11.82 11.87 9.86
Latvia 10.71 10.94 5.96
Lithuania 9.75 8.53 4.28
Luxembourg 10.05 8.26 3.16
Malta 6.42 -1.43 -0.02
Netherlands 9.65 10.44 7.39
Poland 11.98 12.27 15.52
Portugal 12.61 12.11 22.19
Romania 12.02 14.08 18.88
Slovakia 9.00 6.94 -0.97
Slovenia 8.20 6.61 0.83
Spain 11.17 7.22 14.15
Sweden 10.10 7.71 13.22
Note: Values are reported on the log-transformed scale.
B.3 Supplemental Information for NYC Taxi Data Analysis
Figure S.1 provides a detailed Manhattan district map together with time series plots of hourly traffic trends in Midtown and Downtown Manhattan. The district information is used in segmenting districts into different prediction tasks in NYC Taxi data analysis. We could observe a slightly decreasing trend from the beginning of 2014 to the end of 2015 in both midtown and downtown.
Figure S.1: Left panel: An overview of 69 districts in Manhattan. Districts colored with gold represent Midtown and districts colored with green are categorized as Downtown Manhattan. Right panel: Overall trend of the sum of pick-up and drop-off by hour in Midtown and Downtown Manhattan.

### v2

@@
 Numerical results show that our proposed algorithms achieve substantial increases in prediction accuracy and significant reductions in computational time compared to benchmark methods.
-Keywords: Tensor regression; Tensor-on-tensor regression; Tensor factor model;
-Time series forecasting; Temporal convolutional network.
-1 Introduction
-Motivated by a wide range of industrial and scientific applications, forecasting tasks using one tensor series to predict another has become increasingly prevalent in various fields.
-In finance, tensor models are used to forecast future stock prices by analyzing multi-dimensional stock attributes over time (Tran et al.,, 2017).
-In meteorology, tensor models predict future weather conditions using historical data formatted as multi-dimensional tensors (latitude, longitude, time, etc.) (Bilgin et al.,, 2021).
-In neuroscience, tensor models are employed to process medical imaging like MRI/fMRI scans (Wei et al.,, 2023).
-Though predictive modeling of tensors has emerged rapidly over the past decade, most existing studies primarily focus on scenarios when either covariates or responses are tensors, such as scalar-on-tensor regression (Guo et al.,, 2011; Zhou et al.,, 2013; Wimalawarne et al.,, 2016; Li et al.,, 2018; Ahmed et al.,, 2020), matrix-on-tensor regression (Hoff,, 2015; Kossaifi et al.,, 2020), and tensor response regression (Li and Zhang,, 2017; Sun and Li,, 2017).
+Introduction
+Motivated by a wide range of industrial and scientific applications, forecasting tasks using one tensor series to predict another have become increasingly prevalent in various fields.
+In finance, tensor models are used to forecast future stock prices by analyzing multi-dimensional stock attributes over time (Tran et al.
+2017).
+In meteorology, tensor models predict future weather conditions using historical data formatted as multi-dimensional tensors (latitude, longitude, time, etc.) (Bilgin et al.
+2021).
+In neuroscience, tensor models are employed to process medical imaging like MRI/fMRI scans (Wei et al.
+2023).
+Though predictive modeling of tensors has emerged rapidly over the past decade, most existing studies primarily focus on scenarios when either covariates or responses are tensors, such as scalar-on-tensor regression (Guo, Kotsia, and Patras 2011; Zhou, Li, and Zhu 2013; Wimalawarne, Tomioka, and Sugiyama 2016; Li et al.
+2018; Ahmed, Raja, and Bajwa 2020), matrix-on-tensor regression (Hoff 2015; Kossaifi et al.
+2020), and tensor-on-vector regression (Li and Zhang 2017; Sun and Li 2017).
 Tensor-on-tensor [^1] Following the literature, we distinguish “tensor-on-tensor regression” from “tensor regression”.
-“Tensor regression” refers to regression models with tensor covariates and scalar/vector/matrix responses.
+“Tensor regression” refers to regression models with tensor covariates and scalar responses.
 “Tensor-on-tensor regression” refers to regression models with both tensor covariates and tensor responses.
@@
 prediction, when both covariates and responses are tensors, remains a challenging task due to the inherent complexities of tensor structures.
-One commonly used strategy for tensor-on-tensor modeling is to flatten the tensors into matrics (or even vectors) (Kilmer and Martin,, 2011; Choi and Vishwanathan,, 2014) so that matrix-based analytical tools become applicable.
+One commonly used strategy for tensor-on-tensor modeling is to flatten the tensors into matrices (or even vectors) (Kilmer and Martin 2011; Choi and Vishwanathan 2014) so that matrix-based analytical tools become applicable.
 However, flattening can lead to a loss of spatial or temporal information, especially for datasets that are inherently dependent on data structures.
 For instance, in image processing, the relative positions of pixels carry important information about object shapes, which is lost when the image is flattened.
-Alternatively, there is a handful of recent works on tensor-on-tensor regression models that do not adopt flattening but deal with the tensor structures directly; see Lock, (2018); Liu et al., (2020); Gahrooei et al., (2021); Luo and Zhang, (2022).
+Alternatively, there is a handful of recent works on tensor-on-tensor regression models that do not adopt flattening but deal with the tensor structures directly (Lock 2018; Liu, Liu, and Zhu 2020; Gahrooei et al.
+2021; Luo and Zhang 2024).
 However, these methods posit a linear prediction model between covariates and responses, limiting their capacity to capture potential nonlinear relationships.
-Another emerging trend in tensor-on-tensor prediction involves neural networks, such as recurrent neural networks, convolutional neural networks, convolutional recurrent neural networks, and temporal convolutional networks, to name a few.
+Another emerging trend in tensor-on-tensor prediction involves neural networks, such as recurrent neural networks (RNN), convolutional neural networks (CNN), recurrent convolutional neural networks (RCNN), and temporal convolutional networks (TCN), to name a few.
 Deep neural network is a powerful tool in prediction tasks owing to its ability to learn intricate patterns from complex datasets.
@@
 One key innovation of our method is the integration of tensor factor models into deep neural networks for tensor-on-tensor regression.
-Factor models have been a widely utilized tool in matrix/vector data analysis (Bai and Ng,, 2002; Stock and Watson,, 2002; Pan and Yao,, 2008; Lam and Yao,, 2012) for understanding common dependence among multi-dimensional covariates.
-In recent years, researchers have advanced the methodology of factor models to the context of tensor time series (Chen et al., 2022a, ; Chen et al., 2022b, ; Han et al.,, 2020, 2024, 2022; Han and Zhang,, 2022).
-In our models, we borrow the strengths from tensor factor models, offering a more general approach than traditional vector factor analysis for prediction (Sen et al.,, 2019; Fan et al.,, 2017; Yu et al.,, 2022).
+Factor models have been a widely utilized tool in matrix/vector data analysis (Bai and Ng 2002; Stock and Watson 2002; Pan and Yao 2008; Lam and Yao 2012) for understanding common dependence among multi-dimensional covariates.
+In recent years, researchers have advanced the methodology of factor models to the context of tensor time series (Chen et al.
+2022; Chen, Yang, and Zhang 2022; Han et al.
+2024a, b; Han, Chen, and Zhang 2022; Han and Zhang 2023; Chen, Han, and Yu 2024; Yu et al.
+2024).
+In our models, we borrow the strengths from tensor factor models, offering a more general approach than traditional vector factor analysis for prediction (Sen, Yu, and Dhillon 2019; Fan, Xue, and Yao 2017; Yu, Yao, and Xue 2022).
 Our contributions can be summarized in three folds.
@@
   • FATTNN provides a comprehensive framework for the general supervised learning question of predicting tensor responses using tensor covariates.
-Although initially introduced under the problem settings of tensor-on-tensor time series forecasting, the proposed FATTNN framework is not limited to time series data, but can accommodate a variety of data types.
+Although initially introduced under the problem settings of tensor-on-tensor time series forecasting, the proposed framework is not limited to time series data, but can accommodate a variety of data types.
 With different types of factor models and choices of neural network architectures tailored to specific characteristics of the data, FATTNN can be naturally adapted to tensor-on-tensor regression for various data types.
-This paper is organized as follows.
-Section 2 introduces some useful tensor terminology and sets up the problem of interest.
-Section 3 presents the methodological details of the proposed FATTNN algorithms.
-Section 4 includes applications to three real-world datasets to demonstrate the finite-sample performance of our proposed methods.
-Section 5 concludes the paper with discussions.
-Theoretical proofs, simulation studies, and supplemental numerical results are provided in appendices.
-2 Preliminaries
+Related Work
+Neural networks are powerful tools for processing complex tensor data.
+Early works involve RNN and its variations (e.g., Long Short-Term Memory (LSTM) networks) to accommodate the temporal dependence, and CNN to capture the temporal and spatial relationships underlying the data structures.
+Recently, more advanced architectures have been proposed, integrating the strengths of traditional models to capture the increasingly complex data dependence in a more powerful and efficient manner.
+Notable examples include convolutional LSTM (ConvLSTM) (Shi et al.
+2015), recurrent CNN (RCNN) (Liang and Hu 2015), and temporal convolutional networks (TCN)(Bai, Kolter, and Koltun 2018).
+Another popular thread is to incorporate a tensor decomposition procedure in the model training to reduce computational complexity and potentially enhance interpretability.
+Recent examples include tensorizing neural networks (Novikov et al.
+2015), convolutional tensor-train LSTM (Conv-TT-LSTM) (Su et al.
+2020), tensor regression networks (Kossaifi et al.
+2020) and its efficient variant using tensor dropout (Kolbeinsson et al.
+2021).
+Methods may vary depending on how the tensor decomposition is conducted (Fang et al.
+2022; Wang and Zhe 2022; Tao, Tanaka, and Zhao 2024).
+Preliminaries
 Notation.
@@
 The Tucker rank of an order-$K$ tensor ${\cal X}$, denoted by Tucrank$({\cal X})$, is defined as a $K$-tuple $(r_{1},...,r_{K})$, where $r_{k}=\text{rank}(\hbox{\rm mat}_{k}({\cal X}))$.
-Any Tucker rank-$(r_{1},...,r_{K})$ tensor ${\cal X}$ admits the following Tucker decomposition (Tucker,, 1966): ${\cal X}={\cal S}\times_{1}U_{1}\times_{2}\cdots\times_{K}U_{K}$, where ${\cal S}\in\mathbb{R}^{r_{1}\times\cdots\times r_{K}}$ is the core tensor and $U_{k}=\text{SVD}_{r_{k}}(\hbox{\rm mat}_{k}({\cal X}))$ is the mode-$k$ top $r_{k}$ left singular vectors.
-Here, the mode-$k$ product of ${\cal X}\in\mathbb{R}^{d_{1}\times\cdots\times d_{K}}$ with a matrix $B\in\mathbb{R}^{d_{k}\times r_{k}}$, denoted by ${\cal X}\times_{k}B$, is a $d_{1}\times\cdots\times d_{k-1}\times r_{k}\times d_{k+1}\times\cdots\times d_{K}$-dimensional tensor, and its detailed definition is provided in Section 3 and Appendix A.
+Any Tucker rank-$(r_{1},...,r_{K})$ tensor ${\cal X}$ admits the following Tucker decomposition (Tucker 1966): ${\cal X}={\cal S}\times_{1}U_{1}\times_{2}\cdots\times_{K}U_{K}$, where ${\cal S}\in\mathbb{R}^{r_{1}\times\cdots\times r_{K}}$ is the core tensor and $U_{k}=\text{SVD}_{r_{k}}(\hbox{\rm mat}_{k}({\cal X}))$ is the mode-$k$ top $r_{k}$ left singular vectors.
+Here, the mode-$k$ product of ${\cal X}\in\mathbb{R}^{d_{1}\times\cdots\times d_{K}}$ with a matrix $B\in\mathbb{R}^{d_{k}\times r_{k}}$, denoted by ${\cal X}\times_{k}B$, is a $d_{1}\times\cdots\times d_{k-1}\times r_{k}\times d_{k+1}\times\cdots\times d_{K}$-dimensional tensor, and its detailed definition is provided in Appendix A.
 The following abbreviations are used to denote the tensor-matrix product along multiple modes: ${\cal X}\times_{k=1}^{K}U_{k}={\cal X}\times_{1}U_{1}\times_{2}\cdots\times_{K}U_{K}$ and ${\cal X}\times_{\ell\neq k}U_{\ell}={\cal X}\times_{1}U_{1}\times_{2}\cdots\times_{k-1}U_{k-1}\times_{k+1}U_{k+1}\times_{k+2}\cdots\times_{K}U_{K}$.
-For any two matrices $A\in\mathbb{R}^{m_{1}\times r_{1}},B\in\mathbb{R}^{m_{2}\times r_{2}}$, denote the Kronecker product $\odot$ as $A\odot B\in\mathbb{R}^{m_{1}m_{2}\times r_{1}r_{2}}$.
-For any two tensors ${\cal A}\in\mathbb{R}^{m_{1}\times m_{2}\times\cdots\times m_{K}},{\cal B}\in\mathbb{R}^{r_{1}\times r_{2}\times\cdots\times r_{N}}$, denote the tensor product $\otimes$ as ${\cal A}\otimes{\cal B}\in\mathbb{R}^{m_{1}\times\cdots\times m_{K}\times r_{1}\times\cdots\times r_{N}}$, such that $({\cal A}\otimes{\cal B})_{i_{1},...,i_{K},j_{1},...,j_{N}}=({\cal A})_{i_{1},...,i_{K}}({\cal B})_{j_{1},...,j_{N}}.$
+For any two matrices $A\in\mathbb{R}^{m_{1}\times r_{1}},B\in\mathbb{R}^{m_{2}\times r_{2}}$, we denote the Kronecker product $\odot$ as $A\odot B\in\mathbb{R}^{m_{1}m_{2}\times r_{1}r_{2}}$.
+For any two tensors ${\cal A}\in\mathbb{R}^{m_{1}\times m_{2}\times\cdots\times m_{K}},{\cal B}\in\mathbb{R}^{r_{1}\times r_{2}\times\cdots\times r_{N}}$, we denote the tensor product $\otimes$ as ${\cal A}\otimes{\cal B}\in\mathbb{R}^{m_{1}\times\cdots\times m_{K}\times r_{1}\times\cdots\times r_{N}}$, such that $({\cal A}\otimes{\cal B})_{i_{1},...,i_{K},j_{1},...,j_{N}}=({\cal A})_{i_{1},...,i_{K}}({\cal B})_{j_{1},...,j_{N}}.$
 Problem Setup.
@@
 In subsequent discussions, we denote the series in training time range as $\mathcal{X}^{(tr)}=\{\mathcal{X}_{t}\}_{t=1}^{n}$ and $\mathcal{Y}^{(tr)}=\{\mathcal{Y}_{t}\}_{t=1}^{n}$, and the covariates in forecasting time range as $\mathcal{X}^{(te)}=\{\mathcal{X}_{t}\}_{t=n+1}^{n+m}$.
-3 Methodology: Factor Augmented Tensor-on-Tensor time series regression
+Methodology: Factor Augmented Tensor-on-Tensor Time Series Regression
+Figure 1: A graphical illustration of the proposed FATTNN.
+The input is the observed time series $\{({\cal X}_{t},{\cal Y}_{t})\}_{t=1}^{n}$ and the new covariates $\{{\cal X}_{t}\}_{t=n+1}^{n+m}$.
+The output is the forecasted future tensor responses, denoted by $\{\widehat{\cal Y}_{t}\}_{t=n+1}^{n+m}$
 In this section, we introduce our Factor Augmented Tensor-on-Tensor Neural Network, dubbed as FATTNN.
@@
 First, factor tensor features are extracted by low-rank tensor factorization.
-Second, we develop a Tensor-on-Tensor Neural Network based on Temporal Convolutional Network (TCN) (Bai et al.,, 2018).
-In Section 3.1, we present a tensor factor model.
+Second, we develop a Tensor-on-Tensor Neural Network based on Temporal Convolutional Network (TCN) (Bai, Kolter, and Koltun 2018).
+In the first subsection, we present a tensor factor model.
 This model can capture global patterns among the observed time series covariate tensors, by representing each of the original covariate tensor ${\cal X}_{t}$ as a multilinear combination of $r_{1}\times\cdots\times r_{K}$ basis across each tensor mode $k$ plus noise, where $r_{k}\ll d_{k}$.
-It offers a more general approach than the traditional vector factor analysis for prediction (Sen et al.,, 2019; Fan et al.,, 2017; Luo et al.,, 2022; Yu et al.,, 2022; Fan and Gu,, 2023), and preserves the tensor structures of features.
+It offers a more general approach than the traditional vector factor analysis for prediction (Sen, Yu, and Dhillon 2019; Fan, Xue, and Yao 2017; Luo et al.
+2022; Yu, Yao, and Xue 2022; Fan and Gu 2023), and preserves the tensor structures of features.
 It also provides dimension reduction, which significantly reduces computational complexity.
-In Section 3.2, we describe how the output from the global tensor factor model can be used as an input covariate for a TCN to predict future response tensors.
-Figure 1: A graphical illustration of the proposed FATTNN.
-The input is the observed time series $\{({\cal X}_{t},{\cal Y}_{t})\}_{t=1}^{n}$ and the new covariates $\{{\cal X}_{t}\}_{t=n+1}^{n+m}$.
-The output is the forecasted future tensor responses, denoted by $\{\widehat{\cal Y}_{t}\}_{t=n+1}^{n+m}$
-3.1 Tensor Factor Model (TFM)
+In the second subsection, we describe how the output from the global tensor factor model can be used as an input covariate for a TCN to predict future response tensors.
+Tensor Factor Model (TFM)
 Our representation learning module utilizes low rank tensor factorization for covariate tensor.
@@
 Thus, the covariate tensor ${\cal X}_{t}$ can be thought of to be comprised of a basis tensor features ${\cal F}_{t}$ that capture the global patterns in the whole data-set, and all the original covariates are multilinear combinations of these basis tensor features.
-In the following subsection, we will discuss how a TCN can be utilized to leverage the temporal structure of the training data.
+In the next subsection, we will discuss how a TCN can be utilized to leverage the temporal structure of the training data.
 For predicting future response tensors, given a new covariate tensor ${\cal X}_{t}$, we can also extract low-dimensional feature tensors ${\cal F}_{t}$ using the estimated deterministic loading matrices $A_{k}$ and model (1).
@@
 $\widehat{A}_{k}=\text{LSVD}_{r_{k}}\left(\frac{1}{n}\sum_{t=1}^{n}\hbox{\rm mat}_{k}({\cal X}_{t})\hbox{\rm mat}_{k}^{\top}({\cal X}_{t})\right).$ (2)
-The estimation methods in (2) can be further enhanced through an iterative refinement algorithm, detailed in the appendix. The factor tensors in the training data thus can be estimated by
-^$\widehat{\cal F}_{t}={\cal X}_{t}\times_{k=1}^{K}\widehat{A}_{k}^{\top},\quad\text{for }1\leq t\leq n.$
+The estimation methods in (2) can be further enhanced through an iterative refinement algorithm, detailed in the appendix.
+The factor tensors in the training data thus can be estimated by ^$\widehat{\cal F}_{t}={\cal X}_{t}\times_{k=1}^{K}\widehat{A}_{k}^{\top},\text{for }1\leq t\leq n$.
 In the next proposition, we demonstrate the provable benefits of using a tensor factor model with Tucker low-rank structure (1).
 Specifically, we establish the convergence rates of the linear space of the loading matrices, which in turn ensure the accuracy of the estimated factor tensors.
-Proposition 3.1.
+Proposition 1.
 Assume each elements of the idiosyncratic noise ${\cal E}_{t}$ in (1) are i.i.d.
@@
 Then, the iterative TIPUP estimator satisfies
-$\|\widehat{A}_{k}(\widehat{A}_{k}^{\top}\widehat{A}_{k})^{-1}\widehat{A}_{k}^{\top}-A_{k}(A_{k}^{\top}A_{k})^{-1}A_{k}\|_{2}=O_{\mathbb{P}}\left(\frac{\max_{k}\sqrt{d_{k}}}{\lambda\sqrt{n}}+\frac{\max_{k}\sqrt{d_{k}}}{\lambda^{2}\sqrt{n}}\right),\qquad 1\leq k\leq K.$
-3.2 Tensor-on-Tensor Neural Network based TCN (TTNN)
+$\|\widehat{A}_{k}(\widehat{A}_{k}^{\top}\widehat{A}_{k})^{-1}\widehat{A}_{k}^{\top}-A_{k}(A_{k}^{\top}A_{k})^{-1}A_{k}\|_{2}$
+$=$ $O_{\mathbb{P}}\left(\frac{\max_{k}\sqrt{d_{k}}}{\lambda\sqrt{n}}+\frac{\max_{k}\sqrt{d_{k}}}{\lambda^{2}\sqrt{n}}\right),\qquad 1\leq k\leq K.$
+Tensor-on-Tensor Neural Network based TCN (TTNN)
 If we are equipped with a TCN that identifies the temporal patterns in the training dataset $({\cal X}^{(tr)},{\cal Y}^{(tr)})$, we can use this model to enhance the temporal structures in $({\cal F}^{(tr)},{\cal Y}^{(tr)})$.
@@
 Let $\eta(\cdot)$ represent this network.
-The whole temporal network $\eta(\cdot)$ can be trained using the low-rank factor tensors ^$\widehat{\cal F}_{1},\ldots,\widehat{\cal F}_{n}$ derived from the training dataset.
+The whole temporal network $\eta(\cdot)$ can be trained using the low-rank factor tensors ^$\widehat{\cal F}_{1},\ldots,\widehat{\cal F}_{n}$ derived from training dataset.
+In the literature, factor-augmented linear regression has been well studied recently in statistics, economics and machine learning (Stock and Watson 2002; Bai and Ng 2006; Bair et al.
+2006; Fan, Xue, and Yao 2017; Bing et al.
+2021).
 The trained temporal network $\eta(\cdot)$ can be effectively utilized for multi-step look-ahead prediction in a standard approach.
-The factor tensors for future time steps can be estimated using the following formula:
-^$\widehat{\cal F}_{n+h}={\cal X}_{n+h}\times_{k=1}^{K}\widehat{A}_{k}^{\top},$
-where $h>0$. Using the historical data points of factor tensors ^$\widehat{\cal F}_{t},1\leq t\leq n+h-1$, the prediction for the subsequent time step, ${\cal Y}_{n+h}$, is generated by
-$\widehat{\cal Y}_{n+h}=\eta(\widehat{\cal F}_{1},...,\widehat{\cal F}_{n+h}).$
+The factor tensors for future time steps can be estimated using ^$\widehat{\cal F}_{n+h}={\cal X}_{n+h}\times_{k=1}^{K}\widehat{A}_{k}^{\top}$, where $h>0$.
+Using the historical data points of factor tensors ^$\widehat{\cal F}_{t},1\leq t\leq n+h-1$, the prediction for the subsequent time step, ${\cal Y}_{n+h}$, is generated by $\widehat{\cal Y}_{n+h}=\eta(\widehat{\cal F}_{1},...,\widehat{\cal F}_{n+h})$.
 Our hybrid forecasting model incorporates a TCN that inputs past data points from the original raw response time series and the estimated factor tensors, enhancing prediction accuracy.
 The pseudo-code for our model is outlined in Algorithm 1.
+Due to space limitations, the determination of the rank of the factor process is discussed in Appendix A.
 Algorithm 1 Factor Augmented Tensor-on-Tensor Neural Network (FATTNN)
@@
 2:  Output: Forecasts $\widehat{\mathcal{Y}}_{n+1},\ldots,\widehat{\mathcal{Y}}_{n+m}$.
-3:  Compute the factor tensor ^$\widehat{\cal F}_{t}$ for $t=1,\ldots,n$ by TIPUP (2) or its iterative version in Algorithm S.1.
+3:  Compute the factor tensor ^$\widehat{\cal F}_{t}$ for $t=1,\ldots,n$, by TIPUP (2) or its iterative version in Algorithm S.1.
 4:  Fit a TCN $\eta(\cdot)$ to the input sequence of factor tensors ^$\widehat{\cal F}_{1},\ldots,\widehat{\cal F}_{n}$ and the output sequence of response tensor ${\cal Y}_{1},...,{\cal Y}_{n}$ to capture the temporal dynamics.
@@
 10:  return  $\widehat{\mathcal{Y}}_{n+1},\ldots,\widehat{\mathcal{Y}}_{n+m}$.
-4 Numerical Results
+Numerical Results
 In this section, we investigate the finite-sample performance of our proposed FATTNN methods via simulation studies and real-world applications.
-Specifically, we implement the proposed FATTNN [^2] Our code for implementing FATTNN is available in the supplement.
-with TIPUP algorithm to estimate the factor structures.
-For benchmark methods [^3] Multiway is implemented using the code provided by https://github.com/lockEF/MultiwayRegression, and TCN is implemented using the provided by https://github.com/locuslab/TCN .
-, we consider one traditional statistical model - the multiway tensor-on-tensor regression (Lock,, 2018) (denoted by Multiway), and one conventional deep learning approach – the temporal convolutional network (Bai et al.,, 2018) of $\mathcal{Y}$ regressed on $\mathcal{X}$ (denoted by TCN).
-Due to space limitations, simulation results are presented in Appendix 4.1.
+Specifically, we implement the proposed FATTNN [^2] Our code is available on Github: https://github.com/chenezandie/aaai2025-fattnn.
+, we consider one traditional statistical model - the multiway tensor-on-tensor regression (Lock 2018) (denoted by Multiway), and four state-of-the-art deep learning approaches, including the temporal convolutional network (Bai, Kolter, and Koltun 2018) of $\mathcal{Y}$ regressed on $\mathcal{X}$ (denoted by TCN), the long short-term memory network (Hochreiter and Schmidhuber 1997) (denoted by LSTM) , the convolutional tensor-train LSTM (Su et al.
+2020) (denoted by Conv-TT-LSTM), and tensor regression layer (Kossaifi et al.
+2020) (denoted by TRL).
 Evaluation Metrics.
@@
 In addition, we record the computational time of different methods to evaluate computational efficiency.
-4.1 Simulation Studies
+Simulation Studies
+Data Generating Process.
 We carry out simulation studies under different scenarios with various configurations of sample size, data dimensions, tensor ranks, and relationships between covariates and responses.
 Recall that ${\cal X}_{t}\in\mathbb{R}^{d_{1}\times\cdots\times d_{K}}$, ${\cal F}_{t}\in\mathbb{R}^{r_{1}\times\cdots\times r_{K}}$, ${\cal Y}_{t}\in\mathbb{R}^{p_{1}\times\cdots\times p_{q}}$, for $1\leq t\leq n$.
-Data Generating Process.
 We begin with generating the low-rank core tensor ${\cal F}_{t}$ from a tensor autoregressive model ${\rm vec}({\cal F}_{t})=\Phi\cdot{\rm vec}({\cal F}_{t-1})+{\rm vec}({\cal W}_{t})$ for $t=1,\ldots,n$.
-Here, ${\rm vec}({\cal F}_{t})$ denotes the vectorization [^4] We would like to clarify that our proposed FATTNN approach does not involve tensor vectorization in the model fitting.
+Here, ${\rm vec}({\cal F}_{t})$ denotes the vectorization [^3] We’d like to clarify that our proposed FATTNN approach does not involve tensor vectorization in the model fitting.
 This vectorization is solely for the convenience of data generation in simulation.
@@
 […]
-We consider the following three experimental configurations:
-  1. (1)
-    $(d_{1},d_{2},d_{3})=(25,25,12)$,$(r_{1},r_{2},r_{3})=(6,8,6)$, $(p_{1},p_{2},p_{3})=(3,3,2)$
-    $n=500$, $s(z)=cos(z)$, $\sigma_{u}^{2}=1$.
-  2. (2)
-    $(d_{1},d_{2},d_{3})=(30,6,12)$, $(r_{1},r_{2},r_{3})=(6,3,2)$, $(p_{1},p_{2},p_{3})=(8,6,4)$,
-    $n=400$, $s(z)=log(|z|)$, $\sigma_{u}^{2}=1$.
-  3. (3)
-    $(d_{1},d_{2},d_{3})=(12,3,12)$, $(r_{1},r_{2},r_{3})=(4,3,4)$, $(p_{1},p_{2},p_{3})=(3,3,3)$,
-    $n=100$, $s(z)=log(e^{(z)}+1)$, $\sigma_{u}^{2}=0.5$.
+We consider the following experimental configurations:
+(1) $(d_{1},d_{2},d_{3})=(25,25,12)$,$(r_{1},r_{2},r_{3})=(3,3,2)$, $(p_{1},p_{2},p_{3})=(6,8,6)$, $n=500$, $s(z)=cos(z)$, $\sigma_{u}^{2}=1$.
+(2) $(d_{1},d_{2},d_{3})=(30,6,12)$, $(r_{1},r_{2},r_{3})=(6,3,2)$, $(p_{1},p_{2},p_{3})=(8,6,4)$, $n=400$, $s(z)=log(|z|)$, $\sigma_{u}^{2}=1$.
+(3) $(d_{1},d_{2},d_{3})=(12,3,12)$, $(r_{1},r_{2},r_{3})=(4,3,4)$, $(p_{1},p_{2},p_{3})=(3,3,3)$, $n=100$, $s(z)=log(e^{(z)}+1)$, $\sigma_{u}^{2}=0.5$.
 Simulation Evaluations.
 In each setting, we use 70% for training the models and 30% for model evaluation.
-Comparisons of MSE, and computational time are summarized in Table 1 and Table 2.
+Comparisons of MSE, and computational time are summarized in Table 1.
 All reported metrics are averaged over 20 replications.
 Experiments were run on the research computing cluster provided by the University Center for Research Computing.
-From Table 1, we can see that FATTNN improves over the two benchmark methods by a large margin.
+From Table 1, we can see that FATTNN consistently improves over the rest benchmark methods by a large margin.
 The improvements can be attributed to two facets: (a) the utilization of tensor factors, summarizing useful predictive information effectively and reducing the number of noise covariates, and (b) the utilization of neural networks, possessing extraordinary ability in capturing complex relationships in the data and therefore exhibiting exceptional predictive power.
 On a separate note, Table 1 reveals the evident computational advantage of FATTNN compared to other neural network models such as TCN, owing to the effective dimension reduction through tensor factor models.
-Table 1: Comparisons of prediction accuracy of different methods in simulation studies
-(measured by MSE over the testing data)
-Prediction Task TCN FATTNN MultiWay
-Simulation 1 3.642 (1.91) 2.451 (1.57) 3.987 (2.00)
-Simulation 2 670.5 (25.8) 511.9 (22.5) 805.1 (28.4)
-Simulation 3 1074 (32.6) 696.7 (26.4) 1670 (40.3)
-Note: The results are averaged over 20 replications.
-Numbers in the brackets are standard errors.
-Table 2: Comparisons of computational time (in the format of hh:mm:ss)
-of different methods in simulation studies
-Prediction Task TCN FATTNN Muiltway
-Simulation 1 01:22:02 00:16:18 00:02:15
-Simulation 2 01:26:30 00:17:54 00:02:30
-Simulation 3 00:11:16 00:03:41 00:01:03
-4.2 Real Data Examples
+Table 1: Comparisons of prediction accuracy (measured by MSE over the testing data) and computational time (in the format of hh:mm:ss) in simulation studies
+Prediction Task Simulation 1 Simulation 2 Simulation 3
+Test MSE
+TCN 9.812 (3.13) 670.5 (25.8) 1074 (32.6)
+FATTNN 6.353 (2.52) 511.9 (22.5) 696.7 (26.4)
+MultiWay 12.31 (3.51) 805.1 (28.4) 1670 (40.3)
+LSTM 10.48 (3.23) 878.3 (29.6) 807.8 (28.3)
+TRL 12.29 (3.50) 852.6 (29.0) 2003 (38.4)
+Conv-TT-LSTM 10.32 (3.15) 1643 (39.5) 2812 (52.8)
+Computational Time
+TCN 01:22:02 01:26:30 00:11:16
+FATTNN 00:16:18 00:17:54 00:03:41
+MultiWay 00:20:15 00:12:30 00:01:03
+LSTM 00:03:36 00:03:01 00:02:09
+TRL 00:02:25 00:01:30 00:00:50
+Conv-TT-LSTM 03:02:15 05:31:25 00:35:37
+Note: The results are averaged over 20 replications. Numbers in the brackets are standard errors.
+Real Data Examples
 We evaluate the performance of different methods using five prediction tasks on three real-world datasets.
@@
 For the prediction tasks using the FAO dataset, we use 80% data for training and 20% for testing due to its relatively small sample size.
-We compare the prediction accuracy and computation time of three methods: FATTNN, TCN, and Multiway.
-The numerical prediction results along with the corresponding computation time are summarized in Table 4.
+We compare the prediction accuracy and computation time of six methods illustrated in the simulation section.
+The numerical prediction results along with the corresponding computational time are summarized in Table 3.
 Table 2: Data information of the real-world applications
@@
 FAO crop $\sim$ livestock 62 (26,4,5) (26,3,11) (6,4,2)
-NYC taxi – midtown 504 (2,12,12,8) (2,12,12,8) (2,4,4,2)
-NYC taxi – downtown 504 (2,19,19,8) (2,19,19,8) (2,4,4,2)
+NYC taxi – midtown 504 (2,12,12,8) (12,12,8) (2,4,4,2)
+NYC taxi – downtown 504 (2,19,19,8) (19,19,8) (2,4,4,2)
 FMRI data 1452 (25,64,64) (1,64,64) (8,23,23)
-Table 4: Comparisons of prediction accuracy (measured by MSE over the testing data) and computational time (in the format of hh:mm:ss) of different methods in real-world applications
-MSE Time
-Prediction Task TCN FATTNN Multiway TCN FATTNN Multiway
-FAO crop $\sim$ crop 14.11 7.437 12.29 00:55:33 00:15:28 00:03:18
-FAO $\sim$ livestock 32.53 21.65 48.04 00:24:34 00:12:57 00:03:21
-NYC taxi – midtown 12.87 7.399 22.84 01:57:15 00:34:34 00:04:52
-NYC taxi – downtown 14.58 9.493 28.10 03:25:55 00:32:46 00:25:35
-FMRI data 0.1236 0.06634 0.1397 03:54:50 00:20:58 10:01:20
+Table 3: Comparisons of prediction accuracy (measured by MSE over the testing data) and computational time
+(in the format of hh:mm:ss) of different methods in real-world applications
+Prediction Task TCN FATTNN Multiway TRL LSTM Conv-TT-LSTM
+Test MSE
+FAO crop $\sim$ crop 14.11 [10.83, 17.58] 7.437 [6.18, 11.44] 12.29 [11.86, 32.45] 17.84 [15.89, 22.74] 15.76 [13.16, 18.34] 53.39 [52.41, 63.95]
+FAO $\sim$ livestock 32.53 [24.87, 36.10] 21.65 [19.56, 25.10] 48.04 [45.56, 165.80] 66.25 [56.85, 76.13] 34.23 [30.47, 38.48] 69.61 [69.06, 78.44]
+NYC taxi – midtown 12.87 [12.06, 15.52] 7.399 [6.53, 8.03] 22.84 [20.86, 44.79] 8.565 [7.89, 10.12] 8.404 [7.95, 9.63] 25.80 [21.55, 28.62]
+NYC taxi – downtown 14.58 [13.36, 17.42] 9.493 [8.46, 11.03] 28.10 [23.99, 52.12] 12.43 [10.82, 14.50] 12.45 [11.09, 14.38] 42.18 [36.27, 47.66]
+FMRI data 0.1236 [0.0997, 0.148] 0.06634 [0.053, 0.075] 0.1397 [0.135, 0.158] 0.07902 [0.072, 0.084] 0.1964 [0.163, 0.224] 0.03871 [0.0337, 0.0914]
+Computational Time
+FAO crop $\sim$ crop 00:55:33 00:15:28 00:03:18 00:09:41 00:02:08 00:20:07
+FAO $\sim$ livestock 00:24:34 00:12:57 00:03:21 00:07:53 00:01:55 00:16:18
+NYC taxi – midtown 01:57:15 00:34:34 00:04:52 00:19:26 00:05:01 01:54:28
+NYC taxi – downtown 03:25:55 00:32:46 00:25:35 00:22:09 00:05:34 03:07:05
+FMRI data 03:54:50 00:20:58 10:01:20 00:43:51 00:06:27 17:27:05
+Note: Numbers in the square brackets are bootstrap confidence intervals (CIs) of test MSE over 100 bootstrapping replications.
 (1) The United Nations Food and Agriculture Organization (FAO) Crops and Livestock Products Data.
 The database provides agricultural statistics (including crop, livestock, and forestry sub-sectors) collected from countries and territories since 1961.
-It is publicly available at https://www.fao.org/faostat/en/#data/QCL.
-Our analyses focus on 11 crops and 5 livestock products of 46 countries in East Asia, North America, South America, and Europe from 1961 to 2022.
+It is publicly available at https://www.fao.org.
+Our analyses focus on 11 crops and 5 livestock products from 46 countries in East Asia, North America, South America, and Europe from 1961 to 2022.
 Detailed information about the types of crops, livestock, and countries are presented in Table S.1 of Appendix B.
-We study two prediction tasks using the FAO dataset: (i) using Yield and Production data of 11 different crops for 33 countries in East Asia, North America, and Europe (i.e., ${\cal X}_{t}\in\mathbb{R}^{33\times 2\times 11}$) to predict the Yield and Production quantities of the same crops for 13 selected countries in South America (i.e., ${\cal Y}_{t}\in\mathbb{R}^{13\times 2\times 11}$); and (ii) using four agricultural statistics (Producing Animals, Animals-slaughtered, Milk Animals, Laying) associated with 5 kinds of livestock of 26 selected countries in Europe (i.e., ${\cal X}_{t}\in\mathbb{R}^{26\times 4\times 5}$) to predict three metrics (Area-harvested, Production, and Yield) of 11 crops in the same countries (i.e., ${\cal Y}_{t}\in\mathbb{R}^{26\times 3\times 11}$).
+We study two prediction tasks: (i) using Yield and Production data of 11 different crops for 33 countries in East Asia, North America, and Europe (i.e., ${\cal X}_{t}\in\mathbb{R}^{33\times 2\times 11}$) to predict the Yield and Production quantities of the same crops for 13 selected countries in South America (i.e., ${\cal Y}_{t}\in\mathbb{R}^{13\times 2\times 11}$); and (ii) using four agricultural statistics (Producing Animals, Animals-slaughtered, Milk Animals, Laying) associated with 5 kinds of livestock of 26 selected countries in Europe (i.e., ${\cal X}_{t}\in\mathbb{R}^{26\times 4\times 5}$) to predict three metrics (Area-harvested, Production, and Yield) of 11 crops in the same countries (i.e., ${\cal Y}_{t}\in\mathbb{R}^{26\times 3\times 11}$).
 Observing high variability in the raw data, we adopt a log transformation on the raw series and fit our models using log-transformed data.
-Table 4 shows that FATTNN brings in substantial reductions in MSE compared to TCN and Multiway.
-In task (i), FATTNN yields 46.92% and 39.49% reductions in MSE compared to TCN and Multiway, respectively, and in task (ii), the improvements are 33.45% and 54.93% reductions in MSE, respectively.
-As reflected by the table, FATTNN computes much faster than TCN, about 3.6 times faster in task (i) and 2 times faster in tasks (ii).
-We also provide graphical comparisons between ground truth and predictions from different methods; see Figures 2 and 3.
+Figure 2: Left Panel: Comparisons between ground-truth values and predicted Production of Treenuts (top row) and Yield of Oilcrops (bottom row) in South America using the crop data of 33 countries in East Asia, North America, and Europe; Right Panel: Comparisons between ground-truth values and predicted Area-harvested (top row) and Yield (bottom row) of citrus fruit in 26 selected countries in Europe using livestock data of the same 26 countries.
+Values are plotted on the log-transformed scale.
+For readability, numerical values from the figures are tabulated in Appendix B.
+In each panel, from left to right: Ground truth, FATTNN, and TCN.
+Table 3 shows that FATTNN brings in substantial reductions in MSE compared to other methods.
+In task (i), FATTNN yields 47.3%, 39.5%, 58.3%, 52.8%, and 86.1% reductions in MSE compared to TCN, Multiway, TRL, LSTM, and Conv-TT-LSTM, respectively, and in task (ii), the improvements are 33.4%, 54.9%, 67.3%, 36.7%, and 68.9% reductions in MSE, respectively.
+The table also shows that the upper bound of the bootstrap CI for FATTNN is smaller than the lower bounds of the CIs for almost all the benchmarks, indicating FATTNN possesses a statistically significant increase in prediction accuracy.
+What’s more, FATTNN produces the shortest CI, implying the lowest variability among the methods compared.
+As reflected by the table, FATTNN computes much faster than TCN, about 3.6 times faster in task (i) and 2 times faster in task (ii).
+We also provide graphical comparisons between ground truth and predictions from different methods; see Figure 2.
 The plots confirm that FATTNN yields results that most closely align with the ground-truth patterns.
-Figure 2: Comparisons between ground-truth values and predicted Production of Treenuts (top row) and Oilcrops (bottom row) in South America using the crop data of 33 countries in East Asia, North America, and Europe.
-Values are reported on the log-transformed scale.
-From left to right: Ground truth, FATTNN, and TCN.
-Figure 3: Comparisons between ground-truth values and predicted Yield (top row) and Area-harvested (bottom row) of citrus fruit in 26 selected countries in Europe using livestock data of the same 26 countries.
-Values are plotted on the log-transformed scale.
-For readability, numerical numbers are reported in Tables S.2S.3 instead of being superimposed on the plots.
-From left to right: Ground truth, FATTNN, and TCN.
 (2) New York City (NYC) Taxi Trip Data.
 The data contains 24-hour taxi pick-up and drop-off information of 69 areas in New York City for all the business days in 2014 and 2015.
-It is publicly available at https://www.nyc.gov/site/tlc/about/tlc-trip-record-data.page.
+It is publicly available at https://www.nyc.gov.
 Considering the autoregressive nature of this taxi data, we include the lag-1 response as an additional covariate when fitting the models.
@@
 A detailed Manhattan district map is provided in Figure S.3 of Appendix B.
-The Midtown prediction task yields a tensor covariate and response $({\cal X}_{t},{\cal Y}_{t})\in\mathbb{R}^{2\times 12\times 12\times 8}\times\mathbb{R}^{2\times 12\times 12\times 8}$ for each business day, in which the 1st dimension represents two different taxi companies, the 2nd and 3rd dimensions encode the pick-up and drop-off districts, and the 4th dimension represents the hour of day.
-Similarly, the Downtown prediction task involves $({\cal X}_{t},{\cal Y}_{t})\in\mathbb{R}^{2\times 19\times 19\times 8}\times\mathbb{R}^{2\times 19\times 19\times 8}$ for each business day.
-From Table 4, we observe that FATTNN significantly outperforms the other approaches in terms of prediction accuracy, with a 42.5% reduction in MSE compared to TCN and a 67.6% reduction in MSE compared to Multiway in Midtown traffic prediction.
-In Downtown traffic prediction, FATTNN shows a 34.9% reduction in MSE compared to TCN and a 66.2% reduction in MSE compared to Multiway.
-As shown in the table, FATTNN is much more computationally efficient than TCN, about 6.4 times faster than TCN in Downtown prediction and 3.4 times faster in Midtown prediction.
+The Midtown prediction task yields a tensor covariate and response $({\cal X}_{t},{\cal Y}_{t})\in\mathbb{R}^{2\times 12\times 12\times 8}\times\mathbb{R}^{12\times 12\times 8}$ for each business day, in which the 1st dimension stands for the observed information at time $t$, concatenated with lag 1 response, the 2nd and 3rd dimensions encode the pick-up and drop-off districts, and the 4th dimension represents the hour of day.
+Similarly, the Downtown prediction task involves $({\cal X}_{t},{\cal Y}_{t})\in\mathbb{R}^{2\times 19\times 19\times 8}\times\mathbb{R}^{19\times 19\times 8}$for each business day.
+From Table 3, we observe that FATTNN significantly outperforms all other approaches in terms of prediction accuracy, with 42.5%, 67.6%, 13.6%, 12.0%, and 71.3% reductions in MSE compared to TCN, Multiway, TRL, LSTM, and Conv-TT-LSTM, respectively, in Midtown traffic prediction.
+In Downtown traffic prediction, FATTNN shows 34.9%, 66.2%, 23.6%, 23.7%, and 77.5% reductions in MSE compared to TCN, Multiway, TRL, LSTM, and Conv-TT-LSTM, respectively.
+As shown in the table, FATTNN is much more computationally efficient than TCN which directly feeds tensor observations into a convolutional network, about 6.4 times faster than TCN in Downtown prediction and 3.4 times faster in Midtown prediction.
 Additionally, Figure 3 visualizes the comparisons using a 5-day moving average of ground-truth values and predicted pick-up and drop-off volumes by day (where we sum the number of pick-up and drop-off across 8-hour testing periods).
 The FATTNN-predicted lines appear to be closer to the ground-truth lines than TCN-predicted lines in most plots.
-(3) Functional Magnetic Resonance Imaging (FMRI) Data.
-We consider the Haxby dataset (Haxby et al.,, 2001), a well-known public dataset for research in brain imaging and cognitive neuroscience, which can be retrieved from the “NiLearn” Python library.
-The data contains slices of $64\times 64$ FMRI brain images collected from 1452 samples.
-We use the first 25 slices of each sample to predict the 26th slice of that sample, i.e., $(\mathcal{X}_{t},\mathcal{Y}_{t})\in\mathbb{R}^{25\times 64\times 64}\times\mathbb{R}^{1\times 64\times 64}$.
-Similar to the previous two datasets, FATTNN significantly outperforms the other methods in both prediction accuracy and computational efficiency.
-Compared to TCN, FATTNNyields a 46.3% reduction in MSE and computes 11 times faster.
-To further illustrate the algorithm performances, we visualize the predicted images in Figure 5.
-It is evident that FATTNN produces images that more accurately resemble the ground truth.
-In summary, our proposed methods, especially the FATTNN, achieve substantial increases in prediction accuracy compared to benchmark methods – with $\sim 40\%$ reductions in MSE compared to TCN, and even larger reductions compared to Multiway Regression.
-At the same time, the computational costs of FATTNN methods are much lower compared to TCN which directly feeds tensor observations into a convolutional network, ranging from 2 to 11 times faster than TCN.
 Figure 3: Comparisons between ground-truth values and predicted pick-up and drop-off volumes using various methods.
@@
 The district numbering is assigned according to the Manhattan district map shown in Figure S.3.
-Left panel: Midtown Manhattan; Right panel: Downtown Manhattan.
-Figure 5: Comparison between ground-truth values and predicted outputs for selected samples in FMRI dataset. From left to right: Ground truth, FATTNN, and TCN.
-5 Summary & Discussion
+An enlarged figure and more detailed descriptions are in Appendix B.
+Left: Midtown Manhattan; Right: Downtown Manhattan.
+(3) Functional Magnetic Resonance Imaging (FMRI) Data.
+We consider the Haxby dataset (Haxby et al.
+2001), a well-known public dataset for research in brain imaging and cognitive neuroscience, which can be retrieved from the “NiLearn” Python library.
+The data contains slices of $64\times 64$ FMRI brain images collected from 1452 samples.
+We use the first 25 slices of each sample to predict the 26th slice of that sample, i.e., $(\mathcal{X}_{t},\mathcal{Y}_{t})\in\mathbb{R}^{25\times 64\times 64}\times\mathbb{R}^{1\times 64\times 64}$.
+Conv-TT-LSTM yields the lowest MSE in this task.
+This is not surprising as Conv-TT-LSTM is specifically designed for image prediction while FATTNN is more suitable for time-series tensor prediction.
+Other than Conv-TT-LSTM, FATTNN still outperforms other methods.
+Additionally, though Conv-TT-LSTM has the best prediction accuracy, it is quite computationally heavy - about 50 times slower than FATTNN.
+The FATTNN is also 2 times faster than the TRL method which has the 3rd lowest MSE in this task.
+In summary, our proposed FATTNN performs the best in the four prediction tasks on FAO and NYC datasets.
+In the FMRI dataset, the FATTNN yields higher MSE than Conv-TT-LSTM, a method specialized in image prediction.
+That being said, FATTNN has a huge computational advantage over Conv-TT-LSTM.
+We also observe that Conv-TT-LSTM underperforms other methods in the FAO and NYC datasets by a large margin, when the prediction tasks are not focusing on image data.
+To summarize, the FATTNN demonstrates the overall best performance on all the prediction tasks, and strikes an ideal balance between prediction accuracy and computational cost.
+The FATTNN achieves substantial increases in prediction accuracy compared to benchmark methods.
+At the same time, the computational costs of FATTNN methods are much lower compared to TCN and Conv-TT-LSTM, ranging from 2 to 50 times faster.
+Summary & Discussion
 In this paper, we propose FATTNN, a hybrid method that integrates a tensor factor model into deep neural networks, for tensor-on-tensor time series forecasting.
@@
 This approach unifies the strengths of both factor models and neural networks.
-Numerical studies confirm that our proposed methods successfully achieve substantial increases in prediction accuracy and significant reductions in computational time, compared to traditional linear tensor-on-tensor regression models (which often fail to capture the intricate nonlinearity in covariate-response relationships) and conventional deep learning approaches (that often operate in a black-box manner without intention to making use of the inherent structures among observations).
-Limitations.
-In this work, we focus on tensor-on-tensor time series forecasting.
-The integration of TIPUP and TCN, both designed specifically for handling temporal data, makes our proposed approach a powerful tool for forecasting tensor series over time.
-When the data are not time series, the TIPUP-TCN-based FATTNN is still applicable for tensor-on-tensor prediction tasks, but the strengths of TIPUP and TCN may not be fully utilized for modeling non-temporal data.
-This limitation can be easily rectified by adopting proper types of tensor factor models and neural networks in the implementation of FATTNN, depending on the nature of the prediction tasks and data types involved.
-Please see the paragraph on generalizations of FATTNN.
-Generalizations.
+Numerical studies confirm that our proposed methods successfully achieve substantial increases in prediction accuracy and significant reductions in computational time, compared to traditional linear tensor-on-tensor regression models (which often fail to capture the intricate nonlinearity in covariate-response relationships), conventional deep learning methods (that often operate in a black-box manner without making use of the inherent structures among observations), and state-of-the-art tensor-decomposition-based learning methods.
+Limitation and Extension.
 Though we presented our models for the purpose of tensor-on-tensor time series forecasting, our proposed FATTNN framework is not confined to temporal data but can accommodate a variety of data types.
 The main idea of FATTNN is to use a low-rank tensor factor model to capture the intrinsic patterns among the observed covariates, and then proceed with a neural network to model the intricate relationships between covariates and responses.
-With appropriate choices of factor models and neural network architectures, the proposed FATTNN can be naturally adapted to any simple (e.g., i.i.d observations) or complex tensor-type data (e.g., image, graph, network, text, video).
 When handling time series tensors, we adopt TIPUP factorization and TCN to fully capitalize on their temporal nature.
+For non-temporal data, the TIPUP-TCN-based FATTNN is still applicable for tensor-on-tensor prediction tasks, but the strengths of TIPUP and TCN may not be fully utilized.
+For higher-order tensors (e.g.
+$K>4$ ), Tensor-Train (Oseledets 2011) and Hierarchical Tucker decompositions (Lubich et al.
+2013) may be better suited.
+This limitation can be easily rectified: with appropriate choices of factor models and neural network architectures, the proposed FATTNN can be naturally adapted to any simple (e.g., i.i.d observations) or complex tensor-type data (e.g., image, network, text, video).
 To extend beyond time series data, different types of factor models and neural network architectures should be deployed, depending on the nature of the data.
@@
 For text data, Transformer models could substitute for the TCN in our FATTNN framework.
-Our proposed FATTNN is flexible in accommodating different types of tensor factor models and deep learning architecture for tensor-on-tensor prediction using diverse data types.
-Appendices
-Appendix A Tensor Factor Models
-A.1 Notation
+An extended discussion on the generality of the FATTNN framework is presented in Appendix C.
+Our proposed FATTNN is flexible in accommodating different types of tensor factor models and deep learning architectures for tensor-on-tensor prediction using diverse data types.
+Acknowledgments
+We thank the anonymous reviewers for helpful suggestions.
+The work of Yuefeng Han is supported in part by National Science Foundation Grants DMS-2412578.
+The work of Xiufan Yu is supported in part by National Institutes of Health grant R01GM152812.
+Appendix A Appendix A: Tensor Factor Models
+Notation
 For a vector $x=(x_{1},...,x_{p})^{\top}$, define $\|x\|_{q}=(x_{1}^{q}+...+x_{p}^{q})^{1/q}$, $q\geq 1$.
@@
 $\|\widehat{U}\widehat{U}^{\top}-UU^{\top}\|_{2}=\|\sin\Theta(U,\widehat{U})\|_{2}=\sqrt{1-\sigma_{r}^{2}},$ (S.1)
-which equals to the sine of the largest principle angle between the column spaces of $U$ and $\widehat{U}$.
+which equals to the sine of the largest principal angle between the column spaces of $U$ and $\widehat{U}$.
 For any two matrices $A\in\mathbb{R}^{m_{1}\times r_{1}},B\in\mathbb{R}^{m_{2}\times r_{2}}$, denote the Kronecker product $\odot$ as $A\odot B\in\mathbb{R}^{m_{1}m_{2}\times r_{1}r_{2}}$.
@@
 For example, if ${\cal A}\in\mathbb{R}^{m_{1}\times m_{2}\times m_{3}}$, then
-$({\rm{mat}}_{1}({\cal A}))_{i,(j+m_{2}(k-1))}=({\rm{mat}}_{2}({\cal A}))_{j,(k+m_{3}(i-1))}=({\rm{mat}}_{3}({\cal A}))_{k,(i+m_{1}(j-1))}={\cal A}_{ijk}.$
+$({\rm{mat}}_{1}({\cal A}))_{i,(j+m_{2}(k-1))}=({\rm{mat}}_{2}({\cal A}))_{j,(k+m_{3}(i-1))}$
+$=({\rm{mat}}_{3}({\cal A}))_{k,(i+m_{1}(j-1))}={\cal A}_{ijk}.$
 For tensor ${\cal A}\in\mathbb{R}^{m_{1}\times m_{2}\times\cdots\times m_{K}}$, the Hilbert Schmidt norm or Frobenius norm is defined as
 $\|{\cal A}\|_{{\rm F}}=\sqrt{\sum_{i_{1}=1}^{m_{1}}\cdots\sum_{i_{K}=1}^{m_{K}}({\cal A})_{i_{1},...,i_{K}}^{2}}.$
-A.2 Preliminary of matrix and tensor algebra
+For vectors $a_{1},\ldots,a_{K}$ of length $r_{1},\ldots,r_{K}$, the outer product
+${\cal A}=a_{1}\circ a_{2}\circ\cdots\circ a_{K}\in\mathbb{R}^{r_{1}\times\cdots\times r_{K}}$
+is the $K$-th order tensor, with entries ${\cal A}[i_{1},\ldots,i_{K}]=\prod_{k=1}^{K}a_{k}[i_{k}]$.
+Let $M_{1}\in\mathbb{R}^{r_{1}\times R},\ldots M_{K}\in\mathbb{R}^{r_{K}\times R}$ be $K$ matrices with the same column dimension (denoted by $R$), the notation $\llbracket M_{1},\cdots,M_{K}\rrbracket$ represents
+$\llbracket M_{1},\ldots,M_{K}\rrbracket=\sum_{r=1}^{R}m_{1r}\circ\cdots\circ m_{Kr},$ (S.2)
+where $m_{kr}$ is the $r$th column of $M_{k}$, $k=1,\ldots,K$, $r=1,\ldots,R$.
+For two tensor $\mathcal{A}:d_{1}\times\cdots\times d_{K}\times r_{1}\times\cdots\times r_{K}$ and $\mathcal{B}:r_{1}\times\cdots\times r_{K}\times p_{1}\times\cdots\times p_{q}$, the notation $\langle{\cal A},{\cal B}\rangle_{L}$ denote the contracted tensor product (Lock 2018)
+$\langle\mathcal{A},\mathcal{B}\rangle_{L}\in\mathbb{R}^{d_{1}\times\cdots\times d_{K}\times p_{1}\times\cdots\times p_{q}}$
+with the $(i_{1},\ldots,i_{K},j_{1},\ldots,j_{d})$-th element of $\langle\mathcal{A},\mathcal{B}\rangle_{L}$ defined as
+[…]
+Preliminary of matrix and tensor algebra
 Fact (Norm inequalities).
@@
   • For any matrix $A$, and square diagonal invertible matrix $B$, $\sigma_{\min}(BA)\geq\sigma_{\min}(B)\cdot\sigma_{\min}(A)$.
-Lemma A.1 (Weyl’s inequality).
+Lemma 1 (Weyl’s inequality).
 Let $A,B\in\mathbb{R}^{m\times n}$ where $n\geq k$, then we have for any $i\in[k]$,
-$|\sigma_{i}(A)-\sigma_{i}(B)|\leq\|A-B\|$ (S.2)
-Lemma A.2 (Davis–Kahan $\sin\Theta$ theorem).
+$|\sigma_{i}(A)-\sigma_{i}(B)|\leq\|A-B\|.$ (S.4)
+Lemma 2 (Davis–Kahan $\sin\Theta$ theorem).
 Let $\Sigma,\hat{\Sigma}\in\mathbb{R}^{p\times p}$ be symmetric, with eigenvalues $\lambda_{1}\geq\cdots\geq\lambda_{p}$ and $\hat{\lambda}_{1}\geq\cdots\geq\hat{\lambda}_{p}$ respectively.
@@
 $\|\sin\Theta(\hat{v}_{j},v_{j})\|_{2}\leq\frac{\|\hat{\Sigma}-\Sigma\|_{2}}{\min(|\hat{\lambda}_{j-1}-\lambda_{j}|,|\hat{\lambda}_{j+1}-\lambda_{j}|)}.$ (S.6)
-A.3 Technical lemmas
-Lemma A.3 ($\epsilon$-covering of matrix norms, Han et al., (2020)).
+Technical lemmas
+Lemma 3 ($\epsilon$-covering of matrix norms, Han et al. (2024a)).
 Let $d,d_{j},d_{*},r\leq d\wedge d_{j}$ be positive integers, $\epsilon>0$ and $N_{d,\epsilon}=\lfloor(1+2/\epsilon)^{d}\rfloor$.
@@
 (ii) Given $\epsilon>0$, there exist $U_{j}\in\mathbb{R}^{d\times r}$ and $V_{j^{\prime}}\in\mathbb{R}^{d^{\prime}\times r}$ with $\|U_{j}\|_{2}\vee\|V_{j^{\prime}}\|_{2}\leq 1$ such that
-$\max_{M\in\mathbb{R}^{d\times d^{\prime}},\|M\|_{2}\leq 1,\text{rank}(M)\leq r}\ \min_{j\leq N_{dr,\epsilon/2},j^{\prime}\leq N_{d^{\prime}r,\epsilon/2}}\|M-U_{j}V_{j^{\prime}}^{\top}\|_{2}\leq\epsilon.$
+$\max_{M\in\mathbb{R}^{d\times d^{\prime}},\|M\|_{2}\leq 1,\text{rank}(M)\leq r}$
+$\min_{j\leq N_{dr,\epsilon/2},j^{\prime}\leq N_{d^{\prime}r,\epsilon/2}}\|M-U_{j}V_{j^{\prime}}^{\top}\|_{2}\leq\epsilon.$
 Consequently, for any linear mapping $f$ and norm $\|\cdot\|_{*}$ in the range of $f$,
-$\sup_{M,\widetilde{M}\in\mathbb{R}^{d\times d^{\prime}},\|M-\widetilde{M}\|_{2}\leq\epsilon\atop{\|M\|_{2}\vee\|\widetilde{M}\|_{2}\leq 1\atop\text{rank}(M)\vee\text{rank}(\widetilde{M})\leq r}}\frac{\|f(M-\widetilde{M})\|_{*}}{\epsilon 2^{I_{r<d\wedge d^{\prime}}}}\leq\sup_{\|M\|_{2}\leq 1\atop\text{rank}(M)\leq r}\|f(M)\|_{*}\leq 2\max_{1\leq j\leq N_{dr,1/8}\atop 1\leq j^{\prime}\leq N_{d^{\prime}r,1/8}}\|f(U_{j}V_{j^{\prime}}^{\top})\|_{*}.$ (S.5)
-(iii) Given $\epsilon>0$, there exist $U_{j,k}\in\mathbb{R}^{d_{k}\times r_{k}}$ and $V_{j^{\prime},k}\in\mathbb{R}^{d^{\prime}_{k}\times r_{k}}$ with $\|U_{j,k}\|_{2}\vee\|V_{j^{\prime},k}\|_{2}\leq 1$ such that
-$\max_{M_{k}\in\mathbb{R}^{d_{k}\times d_{k}^{\prime}},\|M_{k}\|_{2}\leq 1\atop\text{rank}(M_{k})\leq r_{k},\forall k\leq K}\ \min_{j_{k}\leq N_{d_{k}r_{k},\epsilon/2}\atop j^{\prime}_{k}\leq N_{d_{k}^{\prime}r_{k},\epsilon/2},\forall k\leq K}\Big{\|}\odot_{k=2}^{K}M_{k}-\odot_{k=2}^{K}(U_{j_{k},k}V_{j_{k}^{\prime},k}^{\top})\Big{\|}_{\rm op}\leq\epsilon(K-1).$
+$&\sup_{M,\widetilde{M}\in\mathbb{R}^{d\times d^{\prime}},\|M-\widetilde{M}\|_{2}\leq\epsilon\atop{\|M\|_{2}\vee\|\widetilde{M}\|_{2}\leq 1\atop\text{rank}(M)\vee\text{rank}(\widetilde{M})\leq r}}\frac{\|f(M-\widetilde{M})\|_{*}}{\epsilon 2^{I_{r<d\wedge d^{\prime}}}}\\ \leq&\sup_{\|M\|_{2}\leq 1\atop\text{rank}(M)\leq r}\|f(M)\|_{*}\leq 2\max_{1\leq j\leq N_{dr,1/8}\atop 1\leq j^{\prime}\leq N_{d^{\prime}r,1/8}}\|f(U_{j}V_{j^{\prime}}^{\top})\|_{*}.\end{aligned}$ (S.7)
+(iii) Given $\epsilon>0$, there exist $U_{j,k}\in\mathbb{R}^{d_{k}\times r_{k}}$ and $V_{j^{\prime},k}\in\mathbb{R}^{d^{\prime}_{k}\times r_{k}}$ with $\|U_{j,k}\|_{2}\vee\|V_{j^{\prime},k}\|_{2}\leq 1$ such that $\forall k\leq K$
+$\max_{M_{k}\in\mathbb{R}^{d_{k}\times d_{k}^{\prime}},\atop\|M_{k}\|_{2}\leq 1,\text{rank}(M_{k})\leq r_{k}}\min_{j_{k}\leq N_{d_{k}r_{k},\epsilon/2}\atop j^{\prime}_{k}\leq N_{d_{k}^{\prime}r_{k},\epsilon/2}}\Big{\|}\odot_{k=2}^{K}M_{k}-\odot_{k=2}^{K}(U_{j_{k},k}V_{j_{k}^{\prime},k}^{\top})\Big{\|}_{\rm op}$
+$\leq\epsilon(K-1).$
 For any linear mapping $f$ and norm $\|\cdot\|_{*}$ in the range of $f$,
-$\sup_{M_{k},\widetilde{M}_{k}\in\mathbb{R}^{d_{k}\times d_{k}^{\prime}},\|M_{k}-\widetilde{M}_{k}\|_{2}\leq\epsilon\atop{\text{rank}(M_{k})\vee\text{rank}(\widetilde{M}_{k})\leq r_{k}\atop\|M_{k}\|_{2}\vee\|\widetilde{M}_{k}\|_{2}\leq 1\ \forall k\leq K}}\frac{\|f(\odot_{k=2}^{K}M_{k}-\odot_{k=2}^{K}\widetilde{M}_{k})\|_{*}}{\epsilon(2K-2)}\leq\sup_{M_{k}\in\mathbb{R}^{d_{k}\times d_{k}^{\prime}}\atop{\text{rank}(M_{k})\leq r_{k}\atop\|M_{k}\|_{2}\leq 1,\forall k}}\Big{\|}f\big{(}\odot_{k=2}^{K}M_{k}\big{)}\Big{\|}_{*}$ (S.6)
+$&\sup_{M_{k},\widetilde{M}_{k}\in\mathbb{R}^{d_{k}\times d_{k}^{\prime}},\|M_{k}-\widetilde{M}_{k}\|_{2}\leq\epsilon\atop{\text{rank}(M_{k})\vee\text{rank}(\widetilde{M}_{k})\leq r_{k}\atop\|M_{k}\|_{2}\vee\|\widetilde{M}_{k}\|_{2}\leq 1\ \forall k\leq K}}\frac{\|f(\odot_{k=2}^{K}M_{k}-\odot_{k=2}^{K}\widetilde{M}_{k})\|_{*}}{\epsilon(2K-2)}\\ &\leq\sup_{M_{k}\in\mathbb{R}^{d_{k}\times d_{k}^{\prime}}\atop{\text{rank}(M_{k})\leq r_{k}\atop\|M_{k}\|_{2}\leq 1,\forall k}}\Big{\|}f\big{(}\odot_{k=2}^{K}M_{k}\big{)}\Big{\|}_{*}\end{aligned}$ (S.8)
 and
-$\sup_{M_{k}\in\mathbb{R}^{d_{k}\times d_{k}^{\prime}},\|M_{k}\|_{2}\leq 1\atop\text{rank}(M_{k})\leq r_{k}\ \forall k\leq K}\Big{\|}f\big{(}\odot_{k=2}^{K}M_{k}\big{)}\Big{\|}_{*}\leq 2\max_{1\leq j_{k}\leq N_{d_{k}r_{k},1/(8K-8)}\atop 1\leq j_{k}^{\prime}\leq N_{d_{k}^{\prime}r_{k},1/(8K-8)}}\Big{\|}f\big{(}\odot_{k=2}^{K}U_{j_{k},k}V_{j_{k}^{\prime},k}^{\top}\big{)}\Big{\|}_{*}.$ (S.7)
-Lemma A.4 (Slepian’s inequality, Vershynin, (2018)).
+$&\sup_{M_{k}\in\mathbb{R}^{d_{k}\times d_{k}^{\prime}},\|M_{k}\|_{2}\leq 1\atop\text{rank}(M_{k})\leq r_{k}\ \forall k\leq K}\Big{\|}f\big{(}\odot_{k=2}^{K}M_{k}\big{)}\Big{\|}_{*}\\ \leq&2\max_{1\leq j_{k}\leq N_{d_{k}r_{k},1/(8K-8)}\atop 1\leq j_{k}^{\prime}\leq N_{d_{k}^{\prime}r_{k},1/(8K-8)}}\Big{\|}f\big{(}\odot_{k=2}^{K}U_{j_{k},k}V_{j_{k}^{\prime},k}^{\top}\big{)}\Big{\|}_{*}.\end{aligned}$ (S.9)
+Lemma 4 (Slepian’s inequality, Vershynin (2018)).
 Let $(X_{t})_{t\in T}$ and $(Y_{t})_{t\in T}\in T$ be two mean zero Gaussian Processes. Assume that for all $t,s\in T$, we have
@@
 Consequently,
-$\mathbb{E}\sup_{t\in T}X_{t}\leq\mathbb{E}\sup_{t\in T}Y_{t}$ (S.10)
-Lemma A.5 (Sudakov-Fernique inequality, Vershynin, (2018)).
+$\mathbb{E}\sup_{t\in T}X_{t}\leq\mathbb{E}\sup_{t\in T}Y_{t}.$ (S.12)
+Lemma 5 (Sudakov-Fernique inequality, Vershynin (2018)).
 Let $(X_{t})_{t\in T}$ and $(Y_{t})_{t\in T}$ be two mean-zero Gaussian processes. Assume that, for all $t,s\in T$, we have
@@
 The smallest constant $L>0$ is denoted $\|f\|_{\rm Lip}$.
-Lemma A.6 (Gaussian Concentration of Lipschitz functions, Vershynin, (2018)).
+Lemma 6 (Gaussian Concentration of Lipschitz functions, Vershynin (2018)).
 Let $X\sim N(0,I_{d})$ and let $f:\mathbb{R}^{d}\to\mathbb{R}$ be a Lipschitz function. Then,
 $\mathbb{P}\left(f(X)-\mathbb{E}f(X)\geq t\right)\leq 2\exp\left(-\frac{t^{2}}{2\|f\|_{\rm Lip}}\right).$
-Lemma A.7 (Sigular values of Gaussian random matrices).
-Let $X$ be an $n\times p$ matrix with i.i.d. $N(0,1)$ entries. For any $x>0$,
-$\mathbb{P}\left(\left\|X^{\top}X-nI_{p}\right\|_{2}\geq p+2\sqrt{pn}+2x(\sqrt{p}+\sqrt{n})+x^{2}\right)\leq 2e^{-x^{2}/2}.$
+Lemma 7 (Sigular values of Gaussian random matrices).
+Let $X$ be an $n\times p$ matrix with i.i.d.
+$N(0,1)$ entries.
+For any $x>0$, $\mathbb{P}\left(\left\|X^{\top}X-nI_{p}\right\|_{2}\geq p+2\sqrt{pn}+2x(\sqrt{p}+\sqrt{n})+x^{2}\right)\leq 2e^{-x^{2}/2}.$
 Proof.
-It can be easily derived by Corollary 5.35 in Vershynin, (2010). ∎
-A.4 The Determination of Rank
+It can be easily derived by Corollary 5.35 in Vershynin (2010). ∎
+The Determination of Rank
 Here, the estimators are constructed with given ranks $r_{1},\ldots,r_{K}$.
 Determining the number of factors in a data-driven manner has been a significant research topic in the factor model literature.
-Recently, Han et al., (2022) established a class of rank determination approaches for tensor factor models, based on both the information criterion and the eigen-ratio criterion.
+Recently, (Han, Chen, and Zhang 2022) established a class of rank determination approaches for tensor factor models, based on both the information criterion and the eigen-ratio criterion.
 These procedures can be adopted for our purposes.
-A.5 Tucker Decomposition and iterative TIPUP (iTIPUP) Algorithm
-The Higher-Order Orthogonal Iteration (HOOI) algorithm De Lathauwer et al., (2000) is one of the most popular methods for computing the Tucker decomposition, which relies on the orthogonal projection and Singular Value Decomposition (SVD) to update the estimation of loading matrices for each tensor mode in every iteration.
+Tucker Decomposition and iterative TIPUP (iTIPUP) Algorithm
+The Higher-Order Orthogonal Iteration (HOOI) algorithm (De Lathauwer, De Moor, and Vandewalle 2000) is one of the most popular methods for computing the Tucker decomposition, which relies on the orthogonal projection and Singular Value Decomposition (SVD) to update the estimation of loading matrices for each tensor mode in every iteration.
 Similar to the idea of HOOI, an iterative refinement version of TIPUP could further improve its performance, just as HOOI enhances the original Tucker decomposition method.
 The pseudo-code for this iterative refinement is provided in Algorithm S.1.
+Proof of Proposition 1
+Proof.
+We focus on the case of $K=2$ as the iTIPUP begins with mode-$k$ matrix unfolding.
+In particular, we sometimes give explicit expressions only in the case of $k=1$ and $K=2$.
+For $K=2$, we observe a matrix time series with $X_{t}=A_{1}F_{t}A_{2}^{\top}+{\cal E}_{t}\in\mathbb{R}^{d_{1}\times d_{2}}$.
+Let $\overline{\mathbb{E}}(\cdot)=\mathbb{E}(\cdot|\{{\cal F}_{1},...,{\cal F}_{n}\})$ and ¯()=(|$\overline{\mathbb{P}}(\cdot)=\mathbb{P}(\cdot|\{{\cal F}_{1},...,{\cal F}_{n}\})$.
+Let $L_{k}^{(m)}$ be the loss for $\hat{A}_{k}^{(m)}$ at the $m$-th iteration,
+$L_{k}^{(m)}=\|\widehat{P}_{k}^{(m)}-P_{k}\|_{2},\quad L^{(m)}=\max_{1\leq k\leq K}L_{k}^{(m)},$
+where $\widehat{P}_{k}^{(m)}=A_{k}^{(m)}(A_{k}^{(m)\top}A_{k}^{(m)})^{-1}A_{k}^{(m)^{\top}}$.
+We outline the proof as follows.
+(i) Initialization.
+Define
+${\rm TIPUP}_{1}=\frac{1}{n}\sum_{t=1}^{n}X_{t}X_{t}^{\top}\in\mathbb{R}^{d_{1}\times d_{1}},$
+The “noiseles” version of TIPUP1 is
+$\Theta_{1}=\frac{1}{n}\sum_{t=1}^{n}A_{1}F_{t}A_{2}^{\top}A_{2}F_{1}^{\top}A_{1}^{\top}\in\mathbb{R}^{d_{1}\times d_{1}},$
+which is of rank $r_{1}$. Then, by Davis-Kahan sin $\Theta$ theorem (Lemma 2),
+$L_{1}^{(0)}=\|P_{1}-\widehat{P}_{1}^{(0)}\|_{2}\leq\frac{2\|{\rm TIPUP}_{1}-\Theta_{1}\|_{2}}{\sigma_{r_{1}}(\Theta_{1})}.$ (S.15)
+To bound the numerator on the right-hand side of (S.15), we write
+${\rm TIPUP}_{1}-\Theta_{1}$
+$=$ $\frac{1}{n}\sum_{t=1}^{n}A_{1}F_{t}A_{2}^{\top}{\cal E}_{t}^{\top}+\frac{1}{n}\sum_{t=1}^{n}{\cal E}_{t}A_{2}{\cal F}_{t}^{\top}A_{1}^{\top}+\frac{1}{n}\sum_{t=1}^{n}{\cal E}_{t}{\cal E}_{t}^{\top}$
+::=:= $\Delta_{1}+\Delta_{2}+\Delta_{3}.$
+For $\Delta_{3}$, by Lemma 7,
+$\mathbb{P}\left(\left\|\sum_{t=1}^{n}\mathcal{E}_{t}\mathcal{E}_{t}^{\top}-n\mathbb{E}\mathcal{E}_{t}\mathcal{E}_{t}^{\top}\right\|_{2}\right.$
+$\qquad\geq\left.d_{1}+2\sqrt{d_{1}d_{2}n}+2x(\sqrt{d_{1}}+\sqrt{d_{2}n})+x^{2}\right)$
+$\leq$ $2e^{-x^{2}/2}.$
+By setting $x\asymp\sqrt{d_{1}}$, with probability at least $1-e^{-d_{1}}$,
+$\left\|\Delta_{3}-\mathbb{E}\mathcal{E}_{t}\mathcal{E}_{t}^{\top}\right\|_{2}\lesssim\frac{d_{1}}{n}+\sqrt{\frac{d_{1}d_{2}}{n}}.$
+Next consider $\Delta_{1}$ and $\Delta_{2}$. For any $u$ and $v$ in $\mathbb{R}^{d_{1}}$, we have
+$\overline{\mathbb{E}}\left(u^{\top}\left(\frac{1}{\sqrt{n}}\sum_{t=1}^{n}A_{1}F_{t}A_{2}^{\top}\mathcal{E}_{t}^{\top}\right)v\right)^{2}\leq\|\Theta_{1}\|_{2}\|u\|_{2}^{2}\|v\|_{2}^{2}.$
+Thus for $u_{i}$ and $v_{i}$ with $\|u_{i}\|_{2}=\|v_{i}\|_{2}=1$, $i=1,2$,
+$n\|\Theta_{1}\|_{2}^{-1}\overline{\mathbb{E}}\left((u_{1}^{\top}\Delta_{1}v_{1}-u_{2}^{\top}\Delta_{1}v_{2})^{2}\right)$
+$\leq\left(\|u_{1}-u_{2}\|_{2}\|v_{1}\|_{2}+\|u_{2}\|_{2}\|v_{1}-v_{2}\|_{2}\right)^{2}$
+$\leq 2\left(\|u_{1}-u_{2}\|_{2}^{2}+\|v_{1}-v_{2}\|_{2}^{2}\right)$
+$=2\mathbb{E}\left[(u_{1}-u_{2})^{\top}\xi+(v_{1}-v_{2})^{\top}\zeta\right]^{2},$
+where $\xi$ and $\zeta$ are i.i.d $N(0,I_{d_{1}})$ vectors. As $\Delta_{1}$ is a $d_{1}\times d_{1}$ Gaussian matrix under $\overline{\mathbb{E}}$, the Sudakov-Fernique inequality (Lemma 5) yields
+$\sqrt{n}\|\Theta_{1}\|_{2}^{-1/2}\overline{\mathbb{E}}\|\Delta_{1}\|_{2}\leq\sqrt{2}\mathbb{E}\sup_{\|u\|_{2}=\|v\|_{2}=1}(u^{\top}\xi+v^{\top}\zeta)$
+$=$ $\sqrt{2}\mathbb{E}\left(\|\xi\|_{2}+\|\zeta\|_{2}\right)\leq 2\sqrt{2d_{1}}.$
+It follows that
+$\overline{\mathbb{E}}\|\Delta_{1}\|_{2}\leq\sqrt{\frac{8d_{1}}{n}}\|\Theta_{1}\|_{2}^{1/2}.$
+Elementary calculation shows that $\|\Delta_{1}\|_{2}$ is a $\sqrt{n}\|\Theta_{1}\|_{2}^{1/2}$-Lipschitz function. Then, by Gaussian concentration inequality for Lipschitz functions (Lemma 6),
+¯$\overline{\mathbb{P}}\left(\|\Delta_{1}\|_{2}\geq\sqrt{\frac{8d_{1}}{n}}\|\Theta_{1}\|_{2}^{1/2}+\sqrt{\frac{1}{n}}\|\Theta_{1}\|_{2}^{1/2}x\right)\leq 2e^{-\frac{x^{2}}{2}}.$
+With $x\asymp\sqrt{d_{1}}$, with probability at least $1-e^{-d_{1}}$,
+$\|\Delta_{2}\|_{2}=\|\Delta_{1}\|_{2}\lesssim\sqrt{\frac{d_{1}}{n}}\sigma_{1}(\Theta_{1}).$
+As $\sigma_{1}(\Theta_{1})=\sigma_{r_{1}}(\Theta_{1})\asymp\lambda^{2}$, by (S.15), with probability at least $1-e^{-d_{1}}$,
+$L_{1}^{(0)}=\|\widehat{P}_{1}^{(0)}-P_{1}\|_{2}\lesssim\frac{1}{\lambda^{2}}\sqrt{\frac{d}{n}}+\frac{d_{1}}{\lambda^{2}n}+\frac{1}{\lambda}\sqrt{\frac{d_{1}}{n}}.$ (S.16)
+(ii) Iterative refinement.
+[…]
+${\rm TIPUP_{1}}(U_{2})=\frac{1}{n}\sum_{t=1}^{n}X_{t}U_{2}U_{2}^{\top}X_{t}^{\top}\in\mathbb{R}^{d_{2}\times d_{2}},$
+for any matrix-valued variable $U_{2}\in\mathbb{R}^{d_{2}\times r_{2}}$. Given $\widehat{A}_{2}^{(m)}$, the $(m+1)$-th iteration produces estimates
+$\widehat{A}_{1}^{(m+1)}$ $=\text{LSVD}_{r_{1}}(\text{TIPUP}_{1}(\widehat{A}_{2}^{(m)})),$
+$\widehat{P}_{1}^{(m+1)}$ $=\widehat{A}_{1}^{(m+1)}\widehat{A}_{1}^{(m+1)^{\top}}.$
+The “noiseless” version of this update is given by
+$\Theta_{1}(U_{2})=\frac{1}{n}\sum_{t=1}^{n}A_{1}F_{t}A_{2}^{\top}U_{2}U_{2}^{\top}A_{2}F_{t}^{\top}A_{1}^{\top}=\Theta_{1}$
+giving error-free “estimates”,
+$\widehat{A}_{1}^{(m+1)}=\text{LSVD}_{r_{1}}(\Theta_{1}(\widehat{A}_{2}^{(m)})),$
+where $\Theta_{1}(\widehat{A}_{2}^{(m)})$ is of rank $r_{1}$. Thus, by Davis-Kahan sin $\Theta$ theorem (Lemma 2)
+$L_{1}^{(m+1)}=\|\widehat{P}_{1}^{(m+1)}-P_{1}\|_{2}\leq\frac{2\|\text{TIPUP}_{1}(\widehat{A}_{2}^{(m)})-\Theta_{1}(\widehat{A}_{2}^{(m)})\|_{2}}{\sigma_{r_{1}}[\Theta_{1}(\widehat{A}_{2}^{(m)})]}.$ (S.17)
+To bound the numerator on the right-hand side of (S.17), we write
+$\text{TIPUP}_{1}(U_{2})-\Theta_{1}(U_{2}):=\Delta_{1}(U_{2}U_{2}^{\top})+\Delta_{2}(U_{2}U_{2}^{\top})+\Delta_{3}(U_{2}U_{2}^{\top}),$
+where for any $M_{2}\in\mathbb{R}^{d_{2}\times d_{2}}$,
+$\Delta_{1}(M_{2})$ $=\frac{1}{n}\sum_{t=1}^{n}A_{1}F_{t}A_{2}^{\top}M_{2}\mathcal{E}_{t}^{\top},$
+$\Delta_{2}(M_{2})$ $=\frac{1}{n}\sum_{t=1}^{n}\mathcal{E}_{t}M_{2}A_{2}F_{t}^{\top}A_{1}^{\top},$
+$\Delta_{3}(M_{2})$ $=\frac{1}{n}\sum_{t=1}^{n}\mathcal{E}_{t}M_{2}\mathcal{E}_{t}^{\top}.$
+As $\Delta_{i}(M_{2})$ is linear in $M_{2}$, the numerator on the right-hand side of (S.17) can be bounded by
+$\|\text{TIPUP}_{1}(\widehat{A}_{2}^{(m)})-\Theta_{1}(\widehat{A}_{2}^{(m)})\|_{2}$
+$\leq$ $\|\text{TIPUP}_{1}(A_{2})-\Theta_{1}(A_{2})\|_{2}+L^{(m)}(2K-2)\sum_{i=1}^{3}\|\Delta_{i}(M_{2})\|_{\rm S},$
+where
+$\|\Delta_{i}(M_{2})\|_{\rm S}=\max_{\|M_{2}\|_{1}\leq 1,\text{rank}(M_{2})\leq r_{2}}\|\Delta_{i}(M_{2})\|_{2}.$
+We claim in certain events $\Omega_{i}$, with $\mathbb{P}(\Omega_{i})\geq 1-e^{-d_{1}}$,
+$\|\Delta_{i}(M_{2})\|_{\rm S}$ $\lesssim\sqrt{\frac{d_{1}+d_{2}}{n}}\lambda,\quad i=1,2,$ (S.18)
+$\|\Delta_{3}(M_{2})\|_{\rm S}$ $\lesssim\frac{d_{1}+d_{2}}{n}+\sqrt{\frac{d_{1}+d_{2}}{n}}.$ (S.19)
+We will also prove by induction, the denominator in (S.17) satisfies
+$2\sigma_{r_{1}}[\Theta_{1}(\widehat{A}_{2}^{(m)})]\geq\sigma_{r_{1}}[\Theta_{1}(A_{2})].$ (S.20)
+Define the ideal version of the ratio in (S.17) as
+$L_{k}^{(\text{ideal})}$ $=\frac{\|\text{TIPUP}_{1}(\odot_{j\neq k}A_{j})-\Theta_{k}(\odot_{j\neq k}A_{j})\|_{2}}{\sigma_{r_{1}}[\Theta_{k}(\odot_{j\neq k}A_{j})]},$
+$L^{(\text{ideal})}$ $=\max_{1\leq k\leq K}L_{k}^{(\text{ideal})}.$
+The proof of (S.18) and (S.19) implies that, in an event with probability at least $1-\frac{k}{2}e^{-d_{k}}$,
+$L_{k}^{(\text{ideal})}\lesssim\frac{1}{\lambda^{2}}\sqrt{\frac{d_{k}}{n}}+\frac{d_{k}}{\lambda^{2}n}+\frac{1}{\lambda}\sqrt{\frac{d_{k}}{n}}.$ (S.21)
+Putting together (S.15), (S.16), (S.17), (S.18), (S.19) and (S.20), we have
+$L_{k}^{(m+1)}\leq L_{k}^{(\text{ideal})}+L^{(m)}\rho,$
+with $\rho<1$. By induction,
+$L^{(m+1)}\leq(1+\rho+\dots+\rho^{m})L^{(m)}+\rho^{m+1}L^{(0)}.$
+We achieve the desired results.
+We divided the rest of the proof into 4 steps to prove (S.18) and (S.19) for $i=1,2,3$ and (S.20).
+Step 1.
+We prove (S.18) for the $\|\Delta_{1}(M_{2})\|_{\rm S}$.
+By Lemma 3 (ii), there exists $\widetilde{M}^{(\ell,\ell^{\prime})}\in\mathbb{R}^{d_{2}\times d_{2}}$ of the forms $W_{\ell}Q_{\ell^{\prime}}^{\top}$ with $W_{\ell}\in\mathbb{R}^{d_{2}\times r_{2}}$, $Q_{\ell^{\prime}}\in\mathbb{R}^{d_{2}\times r_{2}}$, $1\leq\ell,\ell^{\prime}\leq N_{d_{2}r_{2},1/8}=17^{d_{2}r_{2}}$, such that $\|\widetilde{M}^{(\ell,\ell^{\prime})}\|_{2}\leq 1$, $\text{rank}(\widetilde{M}^{(\ell,\ell^{\prime})})\leq r_{2}$ and
+$\|\Delta_{1}(M_{2})\|_{\rm S}\leq 2\max_{\ell,\ell^{\prime}\leq N_{d_{2}r_{2},1/8}}\left\|\frac{1}{n}\sum_{t=1}^{n}A_{1}F_{t}A_{2}^{\top}\widetilde{M}^{(\ell,\ell^{\prime})}\mathcal{E}_{t}^{\top}\right\|_{2}.$
+Elementary calculation shows that $\|\Delta_{1}(\widetilde{M}^{(\ell,\ell^{\prime})})\|_{2}$ is a $\sqrt{n}\|\Theta_{1}\|_{2}$-Lipschitz function of $\mathcal{E}_{1},\ldots,\mathcal{E}_{n}$.
+Employing similar argument as the initialization (Sudakov-Fernique inequality, Lemma 5), we have
+$\overline{\mathbb{E}}\left\|\frac{1}{n}\sum_{t=1}^{n}A_{1}F_{t}A_{2}^{\top}\widetilde{M}^{(\ell,\ell^{\prime})}\mathcal{E}_{t}^{\top}\right\|_{2}\leq\sqrt{\frac{8d_{1}}{n}}\|\Theta_{1}\|_{2}^{1/2}.$
+Then, by Gaussian concentration inequality for Lipschitz functions (Lemma 6),
+¯$\overline{\mathbb{P}}\left(\left\|\frac{1}{n}\sum_{t=1}^{n}A_{1}F_{t}A_{2}^{\top}\widetilde{M}^{(\ell,\ell^{\prime})}\mathcal{E}_{t}^{\top}\right\|_{2}-\sqrt{\frac{8d_{1}}{n}}\|\Theta_{1}\|_{2}^{1/2}\right.$
+$\qquad\geq\left.\frac{1}{\sqrt{n}}\|\Theta_{1}\|_{2}^{1/2}x\right)$
+$\leq$ $e^{-x^{2}/2}.$
+Hence,
+¯$\overline{\mathbb{P}}\left(\|\Delta_{1}(M_{2})\|_{\rm S}\geq\sqrt{\frac{8d_{1}}{n}}\|\Theta_{1}\|_{2}^{1/2}+\frac{1}{\sqrt{n}}\|\Theta_{1}\|_{2}^{1/2}x\right)$
+$\leq$ $N_{d_{2}r_{2},1/8}^{2}e^{-x^{2}/2}.$
+This implies that with $x\asymp\sqrt{d_{2}r_{2}}$ that in an event with probability at least $1-e^{-d_{2}}$,
+$\|\Delta_{1}(M_{2})\|_{\rm S}\lesssim\sqrt{\frac{d_{1}+d_{2}r_{2}}{n}}\lambda\lesssim\sqrt{\frac{d_{1}+d_{2}}{n}}\lambda.$
+Step 2.
+The bound of $\|\Delta_{2}(M_{2})\|_{\rm S}$ follows from the same argument as the above step.
+Step 3.
+Here we prove (S.19) for $\|\Delta_{3}(M_{2})\|_{\rm S}$.
+By Lemma 3(iii), we can find $U_{2}^{(\ell)}\in\mathbb{R}^{d_{2}\times r_{2}}$, $1\leq\ell\leq N_{d_{2}r_{2},1/8}$ such that $\|U_{2}^{(\ell)}\|_{2}\leq 1$ and
+$\|\Delta_{3}(M_{2})\|_{\rm S}\leq 2\max_{1\leq\ell\leq N_{d_{2}r_{2},1/8}}\left\|\frac{1}{n}\sum_{t=1}^{n}\mathcal{E}_{t}U_{2}^{(\ell)}U_{2}^{(\ell)\top}\mathcal{E}_{t}^{\top}\right\|_{2}.$
+By Lemma 7,
+$\mathbb{P}\left(\left\|\sum_{t=1}^{n}\mathcal{E}_{t}U_{2}^{(\ell)}U_{2}^{(\ell)\top}\mathcal{E}_{t}^{\top}-n\mathbb{E}\mathcal{E}_{t}U_{2}^{(\ell)}U_{2}^{(\ell)\top}\mathcal{E}_{t}^{\top}\right\|_{2}\right.$
+$\qquad\geq\left.d_{1}+2\sqrt{d_{1}r_{2}n}+2x(\sqrt{d_{1}}+\sqrt{r_{2}n})+x^{2}\right)$
+$\leq$ $2e^{-x^{2}/2}.$
+Hence, by setting $x\asymp\sqrt{d_{2}r_{2}}$, we have with probability at least $1-e^{-d_{2}}$,
+$\|\Delta_{3}(M_{2})\|_{\rm S}$ $\lesssim\frac{d_{1}}{n}+\sqrt{\frac{d_{1}r_{2}}{n}}+\frac{d_{2}r_{2}}{n}+\sqrt{\frac{d_{2}r_{2}^{2}}{n^{2}}}$
+$\lesssim\frac{d_{1}+d_{2}}{n}+\sqrt{\frac{d_{1}+d_{2}}{n}}.$
+Step 4.
+Next, we prove (S.20). Note that
+$\|\Theta_{1}(\widehat{A}_{2}^{(m)})-\Theta_{1}(A_{2})\|_{2}$
+$\leq\left\|\frac{1}{n}\sum_{t=1}^{n}A_{1}F_{t}A_{2}^{\top}(\widehat{A}_{2}^{(m)}\widehat{A}_{2}^{(m)\top}-A_{2}A_{2}^{\top})A_{2}F_{t}^{\top}A_{1}^{\top}\right\|_{2}$
+$\leq\|\widehat{A}_{2}^{(m)}\widehat{A}_{2}^{(m)\top}-A_{2}A_{2}^{\top}\|_{2}\|\Theta_{1}\|_{2}$
+$\leq L^{(m)}\|\Theta_{1}\|_{2}.$
+Thus, by Weyl’s inequality (Lemma 1),
+$\sigma_{r_{1}}[\Theta_{1}(\widehat{A}_{2}^{(m)})]\geq\sigma_{r_{1}}[\Theta_{1}(A_{2})]-L^{(m)}\|\Theta_{1}\|_{2}\geq\frac{1}{2}\sigma_{r_{1}}[\Theta_{1}(A_{2})],$
+when $\min_{1\leq k\leq K}\frac{\sigma_{r_{k}}[\Theta_{k}(A_{2})]}{\|\Theta_{k}(A_{2})\|_{2}}\geq 2L^{(m)}$. We prove this condition by induction. ∎
+Appendix B Appendix B: Supplemental Numerical Results
+Supplement to FAO Data Analysis
+Studies on FAO data can reveal temporal causal links that help inform decisions, policies, and investments related to food and agriculture.
+The dataset has also been examined in other studies (e.g., Chepeliev (2020); Lee and Wang (2023); Chevuru et al.
+(2023)) to address similar issues discussed in this article.
+Predicting yield and production in South America based on data from East Asia, North America and Europe can aid in analyzing global climate effects.
+For example, El Ni no causes droughts in Asia and heavy rains in South America.
+Studying El Ni no’s impact on crop yields in East Asia and North America helps anticipate its effects on South American agriculture (Anderson et al.
+2017).
+Such predictions also enhance our understanding of global market dynamics.
+South America and the US are top soybean producers, and changes in production on one continent influence prices and planting decisions on the other.
+A US drought reducing soybean yields could drive up prices, prompting South American farmers to increase cultivation, affecting future yields (Robinson et al.
+2015).
+Table S.1 illustrates detailed information about the lists of crop types, livestock types, countries, and metrics involved in the two prediction tasks using FAO data:
+  • (i)
+    (crop $\sim$ crop) using Yield and Production data of 11 different crops for 33 countries in East Asia, North America, and Europe (i.e., ${\cal X}_{t}\in\mathbb{R}^{33\times 2\times 11}$) to predict the Yield and Production quantities of the same crops for 13 countries in South America (i.e., ${\cal Y}_{t}\in\mathbb{R}^{13\times 2\times 11}$);
+  • (ii)
+    (crop $\sim$ livestock) using four agricultural statistics (Producing Animals, Animals-slaughtered, Milk Animals, Laying) associated with 5 kinds of livestock of 26 selected countries in Europe (i.e., ${\cal X}_{t}\in\mathbb{R}^{26\times 4\times 5}$) to predict three metrics (Area-harvested, Production, and Yield) of 11 crops in the same countries (i.e., ${\cal Y}_{t}\in\mathbb{R}^{26\times 3\times 11}$).
+Figures 2, S.1 and S.2, visualize the comparisons between ground-truth values and predicted metrics in the two prediction tasks for FAO data.
+The predicted and actual quantities are represented by the color intensity of each country.
+A closer examination reveals that our proposed FATTNN model produces predictions that more closely align with the ground truth compared to the benchmark methods (right panel).
+The color gradients in the FATTNN predictions more accurately reflect the actual data, indicating superior spatial prediction performance.
+We have also included numerical values tabularly reported in Table S.2 and Table S.3, providing a quantitative comparison that reflects the advantages of our method.
+Supplement to NYC Taxi Data Analysis
+Factorized temporal tensor acts like time-varying principal components, summarizing the key information of data in reduced dimensionality.
+Specific interpretation of core tensor varies across datasets.
+For example, in taxi data, the dimension reduction from (19,19) (pick-up, drop-off districts) in raw data to (4,4) in core tensor reflects clusters of geographical regions that share similar traffic patterns, and the reduction from 8 (hours) in raw data to 2 in core tensor summarizes traffic patterns into rush hour and off-peak hour.
+Figure S.3 provides a detailed Manhattan district map together with time series plots of hourly traffic trends in Midtown and Downtown Manhattan.
+The district information is used in segmenting districts into different prediction tasks in NYC Taxi data analysis.
+We could observe a slightly decreasing trend from the beginning of 2014 to the end of 2015 in both midtown and downtown.
+Figure 3 (with an enlarged version shown in Figure S.4) presents a comparison between ground-truth values and predicted pick-up and drop-off volumes using various methods.
+In this figure, the ground truth is depicted by the blue solid line, while our FATTNN predictions are shown as a dark red line with smaller dots.
+The FATTNN line closely follows the ground truth, both in shape and location, outperforming the benchmark method represented by the red line.
+This demonstrates our model’s enhanced ability to capture temporal patterns in the New York taxi data.
+Supplement to FMRI Data Analysis
+Figure S.5 presents a comparison between ground-truth values and predicted outputs for selected samples in FMRI dataset.
+The figure shows that the FMRI image predicted by FATTNN closely resembles the actual FMRI image shown on the left.
+In contrast, the images generated by TCN, LSTM, and TRL without considering tensor factor structures, shows a more blur pattern and noticeable discrepancies.
+The Conv-TT-LSTM also produces a plot that resembles the actual FMRI image well.
+The figure is also consistent with the test MSE results summarized in Table 3, that Conv-TT-LSTM yields the lowest MSE in the FMRI prediction task, followed by the FATTNN.
+This is not surprising as Conv-TT-LSTM is specifically designed for image prediction.
+We also observe a noticeable difference in computation time.
+Though Conv-TT-LSTM has the best prediction accuracy, it is quite computationally heavy - about 50 times slower than FATTNN.
+Appendix C Appendix C: An Extended Discussion
+on the Generalizations of FATTNN
+In this article, we focus on tensor-on-tensor time series forecasting.
+The integration of TIPUP and TCN, both designed specifically for handling temporal data, makes the proposed approach a powerful tool for forecasting tensor series over time.
+For non-temporal data, the TIPUP-TCN-based FATTNN is still applicable for tensor-on-tensor prediction tasks, but the strengths of TIPUP and TCN may not be fully utilized.
+This limitation can be easily rectified by adopting proper types of tensor factor models and neural networks in the implementation of FATTNN, depending on the nature of the prediction tasks and data types involved.
+Our proposed FATTNN framework is not confined to temporal data but can accommodate a variety of data types, though we presented our models for the purpose of tensor-on-tensor time series forecasting.
+The main idea of FATTNN is to use a low-rank tensor factor model to capture the intrinsic patterns among the observed covariates, and then proceed with a neural network to model the intricate relationships between covariates and responses.
+With appropriate choices of factor models and neural network architectures, the proposed FATTNN can be naturally adapted to any simple (e.g., i.i.d observations) or complex tensor-type data (e.g., image, graph, network, text, video).
+When handling time series tensors, we adopt TIPUP factorization and TCN to fully capitalize on their temporal nature.
+For higher-order tensors (e.g.
+$K>4$ ), Tensor-Train (Oseledets 2011) and Hierarchical Tucker decompositions (Lubich et al.
+2013) may be better suited.
+To extend beyond time series data, different types of factor models and neural network architectures should be deployed, depending on the nature of the data.
+For example, for graph data, Graph Neural Network is a more suitable alternative to TCN.
+For text data, Transformer models could substitute for the TCN in our FATTNN framework.
+Our proposed FATTNN is flexible in accommodating different types of tensor factor models and deep learning architectures for tensor-on-tensor prediction using diverse data types.
 Algorithm S.1 Iterative TIPUP (iTIPUP) algorithm
@@
 $\widehat{F}_{t}$ $=\mathcal{X}_{t}\times_{k=1}^{K}\widehat{A}_{K}^{\top},\quad t=1,\ldots,n.$
-A.6 Proof of Proposition 3.1
-We focus on the case of $K=2$ as the iTIPUP begins with mode-$k$ matrix unfolding.
-In particular, we sometimes give explicit expressions only in the case of $k=1$ and $K=2$.
-For $K=2$, we observe a matrix time series with $X_{t}=A_{1}F_{t}A_{2}^{\top}+{\cal E}_{t}\in\mathbb{R}^{d_{1}\times d_{2}}$.
-Let $\overline{\mathbb{E}}(\cdot)=\mathbb{E}(\cdot|\{{\cal F}_{1},...,{\cal F}_{n}\})$ and ¯()=(|$\overline{\mathbb{P}}(\cdot)=\mathbb{P}(\cdot|\{{\cal F}_{1},...,{\cal F}_{n}\})$.
-Let $L_{k}^{(m)}$ be the loss for $\hat{A}_{k}^{(m)}$ at the $m$-th iteration,
-$L_{k}^{(m)}=\|\widehat{P}_{k}^{(m)}-P_{k}\|_{2},\quad L^{(m)}=\max_{1\leq k\leq K}L_{k}^{(m)},$
-where $\widehat{P}_{k}^{(m)}=A_{k}^{(m)}(A_{k}^{(m)\top}A_{k}^{(m)})^{-1}A_{k}^{(m)^{\top}}$.
-We outline the proof as follows.
-(i) Initialization.
-Define
-${\rm TIPUP}_{1}=\frac{1}{n}\sum_{t=1}^{n}X_{t}X_{t}^{\top}\in\mathbb{R}^{d_{1}\times d_{1}},$
-The “noiseles” version of TIPUP1 is
-$\Theta_{1}=\frac{1}{n}\sum_{t=1}^{n}A_{1}F_{t}A_{2}^{\top}A_{2}F_{1}^{\top}A_{1}^{\top}\in\mathbb{R}^{d_{1}\times d_{1}},$
-which is of rank $r_{1}$. Then, by Davis-Kahan sin $\Theta$ theorem (Lemma A.2),
-$L_{1}^{(0)}=\|P_{1}-\widehat{P}_{1}^{(0)}\|_{2}\leq\frac{2\|{\rm TIPUP}_{1}-\Theta_{1}\|_{2}}{\sigma_{r_{1}}(\Theta_{1})}.$ (S.13)
-To bound the numerator on the right hand side of (S.13), we write
-${\rm TIPUP}_{1}-\Theta_{1}=\frac{1}{n}\sum_{t=1}^{n}A_{1}F_{t}A_{2}^{\top}{\cal E}_{t}^{\top}+\frac{1}{n}\sum_{t=1}^{n}{\cal E}_{t}A_{2}{\cal F}_{t}^{\top}A_{1}^{\top}+\frac{1}{n}\sum_{t=1}^{n}{\cal E}_{t}{\cal E}_{t}^{\top}:=\Delta_{1}+\Delta_{2}+\Delta_{3}.$
-For $\Delta_{3}$, by Lemma A.7,
-$\mathbb{P}\left(\left\|\sum_{t=1}^{n}\mathcal{E}_{t}\mathcal{E}_{t}^{\top}-n\mathbb{E}\mathcal{E}_{t}\mathcal{E}_{t}^{\top}\right\|_{2}\geq d_{1}+2\sqrt{d_{1}d_{2}n}+2x(\sqrt{d_{1}}+\sqrt{d_{2}n})+x^{2}\right)\leq 2e^{-x^{2}/2}.$
-By setting $x\asymp\sqrt{d_{1}}$, with probability at least $1-e^{-d_{1}}$,
-$\left\|\Delta_{3}-\mathbb{E}\mathcal{E}_{t}\mathcal{E}_{t}^{\top}\right\|_{2}\lesssim\frac{d_{1}}{n}+\sqrt{\frac{d_{1}d_{2}}{n}}.$
-Next consider $\Delta_{1}$ and $\Delta_{2}$. For any $u$ and $v$ in $\mathbb{R}^{d_{1}}$, we have
-$\overline{\mathbb{E}}\left(u^{\top}\left(\frac{1}{\sqrt{n}}\sum_{t=1}^{n}A_{1}F_{t}A_{2}^{\top}\mathcal{E}_{t}^{\top}\right)v\right)^{2}\leq\|\Theta_{1}\|_{2}\|u\|_{2}^{2}\|v\|_{2}^{2}.$
-Thus for $u_{i}$ and $v_{i}$ with $\|u_{i}\|_{2}=\|v_{i}\|_{2}=1$, $i=1,2$,
-$n\|\Theta_{1}\|_{2}^{-1}\overline{\mathbb{E}}\left((u_{1}^{\top}\Delta_{1}v_{1}-u_{2}^{\top}\Delta_{1}v_{2})^{2}\right)$ $\leq\left(\|u_{1}-u_{2}\|_{2}\|v_{1}\|_{2}+\|u_{2}\|_{2}\|v_{1}-v_{2}\|_{2}\right)^{2}$
-$\leq 2\left(\|u_{1}-u_{2}\|_{2}^{2}+\|v_{1}-v_{2}\|_{2}^{2}\right)$
-$=2\mathbb{E}\left[(u_{1}-u_{2})^{\top}\xi+(v_{1}-v_{2})^{\top}\zeta\right]^{2},$
-where $\xi$ and $\zeta$ are i.i.d $N(0,I_{d_{1}})$ vectors. As $\Delta_{1}$ is a $d_{1}\times d_{1}$ Gaussian matrix under $\overline{\mathbb{E}}$, the Sudakov-Fernique inequality (Lemma A.5) yields
-$\sqrt{n}\|\Theta_{1}\|_{2}^{-1/2}\overline{\mathbb{E}}\|\Delta_{1}\|_{2}\leq\sqrt{2}\mathbb{E}\sup_{\|u\|_{2}=\|v\|_{2}=1}(u^{\top}\xi+v^{\top}\zeta)=\sqrt{2}\mathbb{E}\left(\|\xi\|_{2}+\|\zeta\|_{2}\right)\leq 2\sqrt{2d_{1}}.$
-It follows that
-$\overline{\mathbb{E}}\|\Delta_{1}\|_{2}\leq\sqrt{\frac{8d_{1}}{n}}\|\Theta_{1}\|_{2}^{1/2}.$
-Elementary calculation shows that $\|\Delta_{1}\|_{2}$ is a $\sqrt{n}\|\Theta_{1}\|_{2}^{1/2}$-Lipschitz function. Then, by Gaussian concentration inequality for Lipschitz functions (Lemma A.6),
-¯$\overline{\mathbb{P}}\left(\|\Delta_{1}\|_{2}\geq\sqrt{\frac{8d_{1}}{n}}\|\Theta_{1}\|_{2}^{1/2}+\sqrt{\frac{1}{n}}\|\Theta_{1}\|_{2}^{1/2}x\right)\leq 2e^{-\frac{x^{2}}{2}}.$
-With $x\asymp\sqrt{d_{1}}$, with probability at least $1-e^{-d_{1}}$,
-$\|\Delta_{2}\|_{2}=\|\Delta_{1}\|_{2}\lesssim\sqrt{\frac{d_{1}}{n}}\sigma_{1}(\Theta_{1}).$
-As $\sigma_{1}(\Theta_{1})=\sigma_{r_{1}}(\Theta_{1})\asymp\lambda^{2}$, by (S.13), with probability at least $1-e^{-d_{1}}$,
-$L_{1}^{(0)}=\|\widehat{P}_{1}^{(0)}-P_{1}\|_{2}\lesssim\frac{1}{\lambda^{2}}\sqrt{\frac{d}{n}}+\frac{d_{1}}{\lambda^{2}n}+\frac{1}{\lambda}\sqrt{\frac{d_{1}}{n}}.$ (S.14)
-(ii) Iterative refinement.
-[…]
-${\rm TIPUP_{1}}(U_{2})=\frac{1}{n}\sum_{t=1}^{n}X_{t}U_{2}U_{2}^{\top}X_{t}^{\top}\in\mathbb{R}^{d_{2}\times d_{2}},$
-for any matrix-valued variable $U_{2}\in\mathbb{R}^{d_{2}\times r_{2}}$. Given $\widehat{A}_{2}^{(m)}$, the $(m+1)$-th iteration produces estimates
-$\widehat{A}_{1}^{(m+1)}=\text{LSVD}_{r_{1}}(\text{TIPUP}_{1}(\widehat{A}_{2}^{(m)})),\quad\widehat{P}_{1}^{(m+1)}=\widehat{A}_{1}^{(m+1)}\widehat{A}_{1}^{(m+1)^{\top}}.$ (S.15)
-The “noiseless” version of this update is given by
-$\Theta_{1}(U_{2})=\frac{1}{n}\sum_{t=1}^{n}A_{1}F_{t}A_{2}^{\top}U_{2}U_{2}^{\top}A_{2}F_{t}^{\top}A_{1}^{\top}=\Theta_{1}$
-giving error-free “estimates”,
-$\widehat{A}_{1}^{(m+1)}=\text{LSVD}_{r_{1}}(\Theta_{1}(\widehat{A}_{2}^{(m)})),$
-where $\Theta_{1}(\widehat{A}_{2}^{(m)})$ is of rank $r_{1}$. Thus, by Davis-Kahan sin $\Theta$ theorem (Lemma A.2)
-$L_{1}^{(m+1)}=\|\widehat{P}_{1}^{(m+1)}-P_{1}\|_{2}\leq\frac{2\|\text{TIPUP}_{1}(\widehat{A}_{2}^{(m)})-\Theta_{1}(\widehat{A}_{2}^{(m)})\|_{2}}{\sigma_{r_{1}}[\Theta_{1}(\widehat{A}_{2}^{(m)})]}.$ (S.16)
-To bound the numerator on the right-hand side of (S.16), we write
-$\text{TIPUP}_{1}(U_{2})-\Theta_{1}(U_{2}):=\Delta_{1}(U_{2}U_{2}^{\top})+\Delta_{2}(U_{2}U_{2}^{\top})+\Delta_{3}(U_{2}U_{2}^{\top}),$
-where for any $M_{2}\in\mathbb{R}^{d_{2}\times d_{2}}$,
-$\Delta_{1}(M_{2})$ $=\frac{1}{n}\sum_{t=1}^{n}A_{1}F_{t}A_{2}^{\top}M_{2}\mathcal{E}_{t}^{\top},$
-$\Delta_{2}(M_{2})$ $=\frac{1}{n}\sum_{t=1}^{n}\mathcal{E}_{t}M_{2}A_{2}F_{t}^{\top}A_{1}^{\top},$
-$\Delta_{3}(M_{2})$ $=\frac{1}{n}\sum_{t=1}^{n}\mathcal{E}_{t}M_{2}\mathcal{E}_{t}^{\top}.$
-As $\Delta_{i}(M_{2})$ is linear in $M_{2}$, the numerator on the right-hand side of (S.16) can be bounded by
-$\|\text{TIPUP}_{1}(\widehat{A}_{2}^{(m)})-\Theta_{1}(\widehat{A}_{2}^{(m)})\|_{2}\leq\|\text{TIPUP}_{1}(A_{2})-\Theta_{1}(A_{2})\|_{2}+L^{(m)}(2K-2)\sum_{i=1}^{3}\|\Delta_{i}(M_{2})\|_{\rm S},$
-where
-$\|\Delta_{i}(M_{2})\|_{\rm S}=\max_{\|M_{2}\|_{1}\leq 1,\text{rank}(M_{2})\leq r_{2}}\|\Delta_{i}(M_{2})\|_{2}.$
-We claim in certain events $\Omega_{i}$, with $\mathbb{P}(\Omega_{i})\geq 1-e^{-d_{1}}$,
-$\|\Delta_{i}(M_{2})\|_{\rm S}$ $\lesssim\sqrt{\frac{d_{1}+d_{2}}{n}}\lambda,\quad i=1,2,$ (S.17)
-$\|\Delta_{3}(M_{2})\|_{\rm S}$ $\lesssim\frac{d_{1}+d_{2}}{n}+\sqrt{\frac{d_{1}+d_{2}}{n}}.$ (S.18)
-We will also prove by induction, the denominator in (S.16) satisfies
-$2\sigma_{r_{1}}[\Theta_{1}(\widehat{A}_{2}^{(m)})]\geq\sigma_{r_{1}}[\Theta_{1}(A_{2})].$ (S.19)
-Define the ideal version of the ratio in (S.16) as
-$L_{k}^{(\text{ideal})}=\frac{\|\text{TIPUP}_{1}(\odot_{j\neq k}A_{j})-\Theta_{k}(\odot_{j\neq k}A_{j})\|_{2}}{\sigma_{r_{1}}[\Theta_{k}(\odot_{j\neq k}A_{j})]},\quad L^{(\text{ideal})}=\max_{1\leq k\leq K}L_{k}^{(\text{ideal})}.$
-The proof of (S.17) and (S.18) implies that, in an event with probability at least $1-\frac{k}{2}e^{-d_{k}}$,
-$L_{k}^{(\text{ideal})}\lesssim\frac{1}{\lambda^{2}}\sqrt{\frac{d_{k}}{n}}+\frac{d_{k}}{\lambda^{2}n}+\frac{1}{\lambda}\sqrt{\frac{d_{k}}{n}}.$ (S.20)
-Putting together (S.13),(S.14),(S.16),(S.17),(S.18) and (S.19), we have
-$L_{k}^{(m+1)}\leq L_{k}^{(\text{ideal})}+L^{(m)}\rho,$
-with $\rho<1$. By induction,
-$L^{(m+1)}\leq(1+\rho+\dots+\rho^{m})L^{(m)}+\rho^{m+1}L^{(0)}.$
-We achieve the desired results.
-We divided the rest of the proof into 4 steps to prove (S.17) and (S.18)for $i=1,2,3$ and (S.19).
-Step 1.
-We prove (S.17) for the $\|\Delta_{1}(M_{2})\|_{\rm S}$.
-By Lemma A.3 (ii), there exists $\widetilde{M}^{(\ell,\ell^{\prime})}\in\mathbb{R}^{d_{2}\times d_{2}}$ of the forms $W_{\ell}Q_{\ell^{\prime}}^{\top}$ with $W_{\ell}\in\mathbb{R}^{d_{2}\times r_{2}}$, $Q_{\ell^{\prime}}\in\mathbb{R}^{d_{2}\times r_{2}}$, $1\leq\ell,\ell^{\prime}\leq N_{d_{2}r_{2},1/8}=17^{d_{2}r_{2}}$, such that $\|\widetilde{M}^{(\ell,\ell^{\prime})}\|_{2}\leq 1$, $\text{rank}(\widetilde{M}^{(\ell,\ell^{\prime})})\leq r_{2}$ and
-$\|\Delta_{1}(M_{2})\|_{\rm S}\leq 2\max_{\ell,\ell^{\prime}\leq N_{d_{2}r_{2},1/8}}\left\|\frac{1}{n}\sum_{t=1}^{n}A_{1}F_{t}A_{2}^{\top}\widetilde{M}^{(\ell,\ell^{\prime})}\mathcal{E}_{t}^{\top}\right\|_{2}.$
-Elementary calculation shows that $\|\Delta_{1}(\widetilde{M}^{(\ell,\ell^{\prime})})\|_{2}$ is a $\sqrt{n}\|\Theta_{1}\|_{2}$-Lipschitz function of $\mathcal{E}_{1},\ldots,\mathcal{E}_{n}$.
-Employing similar argument as the initialization (Sudakov-Fernique inequality, Lemma A.5), we have
-$\overline{\mathbb{E}}\left\|\frac{1}{n}\sum_{t=1}^{n}A_{1}F_{t}A_{2}^{\top}\widetilde{M}^{(\ell,\ell^{\prime})}\mathcal{E}_{t}^{\top}\right\|_{2}\leq\sqrt{\frac{8d_{1}}{n}}\|\Theta_{1}\|_{2}^{1/2}.$
-Then, by Gaussian concentration inequality for Lipschitz functions (Lemma A.6),
-¯$\overline{\mathbb{P}}\left(\left\|\frac{1}{n}\sum_{t=1}^{n}A_{1}F_{t}A_{2}^{\top}\widetilde{M}^{(\ell,\ell^{\prime})}\mathcal{E}_{t}^{\top}\right\|_{2}-\sqrt{\frac{8d_{1}}{n}}\|\Theta_{1}\|_{2}^{1/2}\geq\frac{1}{\sqrt{n}}\|\Theta_{1}\|_{2}^{1/2}x\right)\leq e^{-x^{2}/2}.$
-Hence,
-¯$\overline{\mathbb{P}}\left(\|\Delta_{1}(M_{2})\|_{\rm S}\geq\sqrt{\frac{8d_{1}}{n}}\|\Theta_{1}\|_{2}^{1/2}+\frac{1}{\sqrt{n}}\|\Theta_{1}\|_{2}^{1/2}x\right)\leq N_{d_{2}r_{2},1/8}^{2}e^{-x^{2}/2}.$
-This implies that with $x\asymp\sqrt{d_{2}r_{2}}$ that in an event with probability at least $1-e^{-d_{2}}$,
-$\|\Delta_{1}(M_{2})\|_{\rm S}\lesssim\sqrt{\frac{d_{1}+d_{2}r_{2}}{n}}\lambda\lesssim\sqrt{\frac{d_{1}+d_{2}}{n}}\lambda.$
-Step 2.
-The bound of $\|\Delta_{2}(M_{2})\|_{\rm S}$ follows from the same argument as the above step.
-Step 3.
-Here we prove (S.18) for $\|\Delta_{3}(M_{2})\|_{\rm S}$.
-By Lemma A.3(iii), we can find $U_{2}^{(\ell)}\in\mathbb{R}^{d_{2}\times r_{2}}$, $1\leq\ell\leq N_{d_{2}r_{2},1/8}$ such that $\|U_{2}^{(\ell)}\|_{2}\leq 1$ and
-$\|\Delta_{3}(M_{2})\|_{\rm S}\leq 2\max_{1\leq\ell\leq N_{d_{2}r_{2},1/8}}\left\|\frac{1}{n}\sum_{t=1}^{n}\mathcal{E}_{t}U_{2}^{(\ell)}U_{2}^{(\ell)\top}\mathcal{E}_{t}^{\top}\right\|_{2}.$
-By Lemma A.7,
-$\mathbb{P}\left(\left\|\sum_{t=1}^{n}\mathcal{E}_{t}U_{2}^{(\ell)}U_{2}^{(\ell)\top}\mathcal{E}_{t}^{\top}-n\mathbb{E}\mathcal{E}_{t}U_{2}^{(\ell)}U_{2}^{(\ell)\top}\mathcal{E}_{t}^{\top}\right\|_{2}\geq d_{1}+2\sqrt{d_{1}r_{2}n}+2x(\sqrt{d_{1}}+\sqrt{r_{2}n})+x^{2}\right)\leq 2e^{-x^{2}/2}.$
-Hence, by setting $x\asymp\sqrt{d_{2}r_{2}}$, we have with probability at least $1-e^{-d_{2}}$,
-$\|\Delta_{3}(M_{2})\|_{\rm S}\lesssim\frac{d_{1}}{n}+\sqrt{\frac{d_{1}r_{2}}{n}}+\frac{d_{2}r_{2}}{n}+\sqrt{\frac{d_{2}r_{2}^{2}}{n^{2}}}\lesssim\frac{d_{1}+d_{2}}{n}+\sqrt{\frac{d_{1}+d_{2}}{n}}.$
-Step 4.
-Next, we prove (S.19). Note that
-$\|\Theta_{1}(\widehat{A}_{2}^{(m)})-\Theta_{1}(A_{2})\|_{2}$ $\leq\left\|\frac{1}{n}\sum_{t=1}^{n}A_{1}F_{t}A_{2}^{\top}(\widehat{A}_{2}^{(m)}\widehat{A}_{2}^{(m)\top}-A_{2}A_{2}^{\top})A_{2}F_{t}^{\top}A_{1}^{\top}\right\|_{2}$
-$\leq\|\widehat{A}_{2}^{(m)}\widehat{A}_{2}^{(m)\top}-A_{2}A_{2}^{\top}\|_{2}\|\Theta_{1}\|_{2}$
-$\leq L^{(m)}\|\Theta_{1}\|_{2}.$
-Thus, by Weyl’s inequality (Lemma A.1),
-$\sigma_{r_{1}}[\Theta_{1}(\widehat{A}_{2}^{(m)})]\geq\sigma_{r_{1}}[\Theta_{1}(A_{2})]-L^{(m)}\|\Theta_{1}\|_{2}\geq\frac{1}{2}\sigma_{r_{1}}[\Theta_{1}(A_{2})],$
-when $\min_{1\leq k\leq K}\frac{\sigma_{r_{k}}[\Theta_{k}(A_{2})]}{\|\Theta_{k}(A_{2})\|_{2}}\geq 2L^{(m)}$. We prove this condition by induction.
-Appendix B Supplemental Results to Numerical Studies in Section 4
-B.1 Supplemental Notation used in Simulation Studies
-(1) For vectors $a_{1},\ldots,a_{K}$ of length $r_{1},\ldots,r_{K}$, the outer product
-${\cal A}=a_{1}\circ a_{2}\circ\cdots\circ a_{K}\in\mathbb{R}^{r_{1}\times\cdots\times r_{K}}$
-is the $K$-th order tensor, with entries ${\cal A}[i_{1},\ldots,i_{K}]=\prod_{k=1}^{K}a_{k}[i_{k}]$.
-(2) Let $M_{1}\in\mathbb{R}^{r_{1}\times R},\ldots M_{K}\in\mathbb{R}^{r_{K}\times R}$ be $K$ matrices with the same column dimension (denoted by $R$), the notation $\llbracket M_{1},\cdots,M_{K}\rrbracket$ represents
-$\llbracket M_{1},\ldots,M_{K}\rrbracket=\sum_{r=1}^{R}m_{1r}\circ\cdots\circ m_{Kr},$ (S.21)
-where $m_{kr}$ is the $r$th column of $M_{k}$, $k=1,\ldots,K$, $r=1,\ldots,R$.
-(3) For two tensor $\mathcal{A}:d_{1}\times\cdots\times d_{K}\times r_{1}\times\cdots\times r_{K}$ and $\mathcal{B}:r_{1}\times\cdots\times r_{K}\times p_{1}\times\cdots\times p_{q}$, the notation $\langle{\cal A},{\cal B}\rangle_{L}$ denote the contracted tensor product (Lock,, 2018)
-$\langle\mathcal{A},\mathcal{B}\rangle_{L}\in\mathbb{R}^{d_{1}\times\cdots\times d_{K}\times p_{1}\times\cdots\times p_{q}}$
-with the $(i_{1},\ldots,i_{K},j_{1},\ldots,j_{d})$-th element of $\langle\mathcal{A},\mathcal{B}\rangle_{L}$ defined as
-$\langle\mathcal{A},\mathcal{B}\rangle_{L}[i_{1},\ldots,i_{K},j_{1},\ldots,j_{d}]=\sum_{l_{1}=1}^{r_{1}}\cdots\sum_{l_{K}=1}^{r_{K}}\mathcal{A}[i_{1},\dots,i_{K},l_{1},\dots,l_{K}]\mathcal{B}[l_{1},\dots,l_{K},j_{1},\dots,j_{d}].$ (S.22)
-B.2 Supplemental Results for FAO Data Analysis
-Table S.1 illustrates detailed information about the lists of crop types, livestock types, countries, and metrics involved in the two prediction tasks using FAO data:
-  • (i)
-    (crop $\sim$ crop) using Yield and Production data of 11 different crops for 33 countries in East Asia, North America, and Europe (i.e., ${\cal X}_{t}\in\mathbb{R}^{33\times 2\times 11}$) to predict the Yield and Production quantities of the same crops for 13 countries in South America (i.e., ${\cal Y}_{t}\in\mathbb{R}^{13\times 2\times 11}$);
-  • (ii)
-    (crop $\sim$ livestock) using four agricultural statistics (Producing Animals, Animals-slaughtered, Milk Animals, Laying) associated with 5 kinds of livestock of 26 selected countries in Europe (i.e., ${\cal X}_{t}\in\mathbb{R}^{26\times 4\times 5}$) to predict three metrics (Area-harvested, Production, and Yield) of 11 crops in the same countries (i.e., ${\cal Y}_{t}\in\mathbb{R}^{26\times 3\times 11}$).
-Table S.2 and Table S.3 supplement Figure 3 by providing the numerical values of ground truth and predicted Yield metrics for the plot.
 Table S.1: Details of statistics and metrics involved in the FAO data analysis
@@
 Note: A full description of the code definitions is available on the United Nations Food and Agriculture Organization Crops and Livestock Products Database website.
-Table S.2: Numerical values of ground-truth and predicted Yield of cirtus fruit in 26 European countries using livestock data of the same 26 countries - a supplement to Figure 3
+Figure S.1: Comparisons between ground-truth values and predicted Production of Treenuts (top row) and Yield of Oilcrops (bottom row) in South America using the crop data of 33 countries in East Asia, North America, and Europe.
+Values are plotted on the log-transformed scale.
+From left to right: Ground truth, FATTNN, TCN, LSTM, TRL, and Conv-TT-LSTM.
+Figure S.2: Comparisons between ground-truth values and predicted Area-harvested (top row) and Yield (bottom row) of citrus fruit in 26 selected countries in Europe using the livestock data of the same 26 countries.
+Values are plotted on the log-transformed scale.
+From left to right: Ground truth, FATTNN, TCN, LSTM, TRL, and Conv-TT-LSTM.
+Table S.2: Numerical values of ground-truth and predicted (a) Production of Treenuts and (b) yield Oilcrops in 13 South America countries using crop data of 33 countries in East Asia, North America, and Europe - a supplement to Figure 2.
+(a) Predicted Production of Treenuts
+Country Ground Truth FATTNN TCN ($\mathcal{Y}\sim\mathcal{X}$)
+Argentina 11.77 14.77 4.30
+Bolivia 10.84 11.15 6.48
+Brazil 13.55 13.73 9.56
+Chile 9.91 8.88 7.81
+Colombia 11.35 10.66 6.51
+Ecuador 10.51 11.32 7.67
+French Guiana -9.21 -0.71 2.80
+Guyana 7.30 4.34 3.98
+Paraguay 9.32 9.68 6.11
+Peru 11.20 10.40 7.46
+Suriname 7.42 7.03 6.02
+Uruguay 9.61 5.87 5.78
+Venezuela 10.69 15.72 11.30
+(b) Predicted Yield of Oilcrops
+Country Ground Truth FATTNN TCN ($\mathcal{Y}\sim\mathcal{X}$)
+Argentina 9.69 13.53 4.31
+Bolivia 9.91 9.49 7.11
+Brazil 7.99 9.91 6.48
+Chile 10.18 10.30 8.30
+Colombia 10.31 9.67 1.51
+Ecuador -9.21 -5.90 -6.36
+French Guiana -9.21 -9.99 -10.10
+Guyana -9.21 -8.37 -8.57
+Paraguay -9.21 -6.97 -7.43
+Peru 11.24 7.06 8.53
+Suriname -9.21 -8.32 -5.64
+Uruguay -9.21 -10.50 -7.90
+Venezuela -9.21 -11.01 -5.29
+Table S.3: Numerical values of ground-truth and (a) predicted Area-Harvested of citrus fruit (b) predicted Yield of citrus fruit, in 26 European countries using livestock data from the same 26 countries - a supplement to Figure 2.
+(a) Predicted Area-Harvested of Citrus Fruit
+Country Ground Truth FATTNN TCN ($\mathcal{Y}\sim\mathcal{X}$)
+Austria 10.01 10.74 13.54
+Bulgaria 9.36 11.81 13.65
+Croatia 9.18 10.03 0.33
+Cyprus 8.39 12.07 11.58
+Czechia 10.07 9.54 7.65
+Denmark 8.55 7.98 1.90
+Estonia 9.99 10.53 13.61
+Finland 12.14 14.00 11.41
+France 12.44 11.48 9.40
+Germany 10.07 11.60 10.08
+Greece 9.65 12.03 12.86
+Hungary 9.14 11.37 8.96
+Ireland 10.84 12.24 9.91
+Italy 11.82 11.87 9.86
+Latvia 10.71 10.94 5.96
+Lithuania 9.75 8.53 4.28
+Luxembourg 10.05 8.26 3.16
+Malta 6.42 -1.43 -0.02
+Netherlands 9.65 10.44 7.39
+Poland 11.98 12.27 15.52
+Portugal 12.61 12.11 22.19
+Romania 12.02 14.08 18.88
+Slovakia 9.00 6.94 -0.97
+Slovenia 8.20 6.61 0.83
+Spain 11.17 7.22 14.15
+Sweden 10.10 7.71 13.22
+(b) Predicted Yield of Citrus Fruit
 Country Ground Truth FATTNN TCN ($\mathcal{Y}\sim\mathcal{X}$)
@@
 Sweden 13.63 10.19 16.74
-Note: Values are reported on the log-transformed scale.
-Table S.3: Numerical values of ground-truth and predicted Area-harvested of cirtus fruit in 26 European countries using livestock data of the same 26 countries - a supplement to Figure 3
-Country True Value FATTNN TCN ($\mathcal{Y}\sim\mathcal{X}$)
-Austria 10.01 10.74 13.54
-Bulgaria 9.36 11.81 13.65
-Croatia 9.18 10.03 0.33
-Cyprus 8.39 12.07 11.58
-Czechia 10.07 9.54 7.65
-Denmark 8.55 7.98 1.90
-Estonia 9.99 10.53 13.61
-Finland 12.14 14.00 11.41
-France 12.44 11.48 9.40
-Germany 10.07 11.60 10.08
-Greece 9.65 12.03 12.86
-Hungary 9.14 11.37 8.96
-Ireland 10.84 12.24 9.91
-Italy 11.82 11.87 9.86
-Latvia 10.71 10.94 5.96
-Lithuania 9.75 8.53 4.28
-Luxembourg 10.05 8.26 3.16
-Malta 6.42 -1.43 -0.02
-Netherlands 9.65 10.44 7.39
-Poland 11.98 12.27 15.52
-Portugal 12.61 12.11 22.19
-Romania 12.02 14.08 18.88
-Slovakia 9.00 6.94 -0.97
-Slovenia 8.20 6.61 0.83
-Spain 11.17 7.22 14.15
-Sweden 10.10 7.71 13.22
-Note: Values are reported on the log-transformed scale.
-B.3 Supplemental Information for NYC Taxi Data Analysis
-Figure S.1 provides a detailed Manhattan district map together with time series plots of hourly traffic trends in Midtown and Downtown Manhattan.
-The district information is used in segmenting districts into different prediction tasks in NYC Taxi data analysis.
-We could observe a slightly decreasing trend from the beginning of 2014 to the end of 2015 in both midtown and downtown.
 Figure S.3: Left panel: An overview of 69 districts in Manhattan.
@@
 Right panel: Overall trend of the sum of pick-up and drop-off by hour in Midtown and Downtown Manhattan.
+Figure S.4: Comparisons between ground-truth values and predicted pick-up and drop-off volumes using various methods.
+This plot is an enlarged version of Figure 3 of the main context.
+“District A to B” denotes the traffic volume that passengers were picked up in District A and dropped off in District B.
+The district numbering is assigned according to the Manhattan district map shown in Figure S.3.
+Left: Midtown Manhattan; Right: Downtown Manhattan.
+Figure S.5: Comparisons between ground-truth values and predicted outputs for selected samples in FMRI dataset. From left to right: Ground truth, FATTNN, TCN, LSTM, LSTM, TRL, and Conv-TT-LSTM,.